### v1

# 3cm3cm3cm3cm

Abstract.
We study the joint action of the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ for the general linear Lie algebra ${\mathfrak{gl}}_{d}$ with respect to ${\bf w}\in{\mathbb{C}}^{d}$ and the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ for the general linear Lie superalgebra ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$ on the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators. We establish a duality, called the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$, which is an equivalence between the actions of the algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on the Fock space. As an application, we show that the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on each weight space of the evaluation module $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})$, where $\underline{M\mkern-4.0mu}\mkern 4.0mu$ is a $d$-fold tensor product of certain classes of infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules, is a cyclic ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module, and that ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ is diagonalizable with a simple spectrum on the weight space for generic ${\bf w}$ and ${\bf z}$.
1. Introduction
Let $\mathcal{F}:={\mathbb{C}}[{\bf x},{\bf y},{\boldsymbol{\eta}},{\boldsymbol{\zeta}}]$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$ and $\zeta^{a}_{s}$, for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$ and $a=1,\ldots,d$. Here $x^{a}_{i}$ and $y^{a}_{r}$ are even variables while $\eta^{a}_{j}$ and $\zeta^{a}_{s}$ are odd. The superalgebra $\mathcal{F}$ can be realized as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators. Let $\mathcal{D}:={\mathbb{D}}[{\bf x},{\bf y},{\boldsymbol{\eta}},{\boldsymbol{\zeta}}]$ be the corresponding Weyl superalgebra, i.e., $\mathcal{D}$ is the superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$ and $\zeta^{a}_{s}$ as well as their derivatives $\frac{\partial}{\partial x_{i}^{a}}$, $\frac{\partial}{\partial\eta_{j}^{a}}$, $\frac{\partial}{\partial y_{r}^{a}}$ and $\frac{\partial}{\partial\zeta_{r}^{a}}$. There is a natural action of $\mathcal{D}$ on $\mathcal{F}$, and hence $\mathcal{D}$ is a subalgebra of the endomorphism algebra ${\rm End}(\mathcal{F})$ of $\mathcal{F}$.
Moreover, there exist superalgebra homomorphisms $\phi:U({\mathfrak{gl}}_{d})\longrightarrow\mathcal{D}$ and $\overline{\phi}:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$ from the universal enveloping algebras of ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$ to $\mathcal{D}$ defined by (4.2) and (4.3), respectively. These maps induce a ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-action on $\mathcal{F}$. It is well-known that $\mathcal{F}$ decomposes into a direct sum of unitarizable ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-modules. This is called the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ ([CLZ]). Such unitarizable modules are, in general, infinite-dimensional. Taking $p=q=0$, the maps $\phi$ and $\overline{\phi}$ specialize to the superalgebra homomorphisms $U({\mathfrak{gl}}_{d})\longrightarrow{\mathbb{D}}[{\bf x},{\boldsymbol{\eta}}]$ and $U({\mathfrak{gl}}_{m|n})\longrightarrow{\mathbb{D}}[{\bf x},{\boldsymbol{\eta}}]$, respectively, and the duality reduces to the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$, which yields a decomposition of ${\mathbb{C}}[{\bf x},{\boldsymbol{\eta}}]$ into a direct sum of finite-dimensional ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{m|n}$-modules ([CW, Se1, Se2]). Assuming further that $n=0$, it gives the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ ([H]).
Fix ${\bf w}\in{\mathbb{C}}^{\ell}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$. There are natural extensions of $\phi$ and $\overline{\phi}$ to the universal enveloping algebras of current algebras, and we have superalgebra homomorphisms $\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])\longrightarrow\mathcal{D}$ and $\overline{\phi}_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])\longrightarrow\mathcal{D}$; see (4.4) and (4.5). It is natural to ask if dualities exist in the context of current algebras. When $p=q=n=0$, a duality phenomenon is noticed in [TL] (see also [TV]). In [MTV3], a precise statement of the duality between $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{m}[t])$ is made and proved. So far, various generalizations have emerged (see [HM, TU, VY]). All of these are related to the Gaudin models for general linear Lie (super)algebras.
We will establish a duality between $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$ in this paper. We consider the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ for ${\mathfrak{gl}}_{d}$ with respect to ${\bf w}$ and the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ for ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}$; see Section 3.3 and Section 3.4. As the algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ are respectively subalgebras of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$, the maps $\phi_{\bf z}$ and $\overline{\phi}_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ and an action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on the polynomial superalgebra $\mathcal{F}$, respectively. We show that these two actions are equivalent. More precisely, we have:
Theorem 1.1 (Theorem 4.7).
$\phi_{\bf z}\big{(}\mathcal{B}_{d}^{\bf w}\big{)}=\overline{\phi}_{\bf w}\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}$.
We name Theorem 1.1 the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$. It recovers several dualities established earlier. After specializing to $p=q=0$, it recovers the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$ due to Huang and Mukhin [HM]. If we assume further that $n=0$ (resp., $m=0$), Theorem 1.1 gives the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ obtained by Mukhin, Tarasov, and Varchenko [MTV3] (resp., a variant of the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{n})$ obtained by Tarasov and Uvarov [TU]). Furthermore, we will show in [ChL3] that there is another duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for Gaudin models with irregular singularities, which specializes to a variant of Theorem 1.1.
The Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ enables us to study the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on $\mathcal{F}$ via the action of $\mathcal{B}_{d}^{\bf w}$ on $\mathcal{F}$, and vice versa. Using the properties of the action of $\mathcal{B}_{d}^{\bf w}$ on finite-dimensional ${\mathfrak{gl}}_{d}$-modules, we will better understand the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules. Let $\underline{M\mkern-4.0mu}\mkern 4.0mu:=M_{1}\otimes\cdots\otimes M_{d}$, where each $M_{i}$ is an infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module with highest weight corresponding to a generalized partition of depth 1 (see Section 5.2 for a detailed description). We will see in Theorem 5.1 that each weight space of the evaluation ${\mathfrak{gl}}_{p+m|q+n}[t]$-module $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})$ is a cyclic ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module, and that ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ is diagonalizable with a simple spectrum on the weight space for generic ${\bf w}$ and ${\bf z}$.
Let us explain a little how we prove Theorem 1.1. The Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ are determined by the column determinant ${\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})$ and the Berezinian ${\rm Ber}^{\bf s}\,({\mathcal{L}}_{{p+m|q+n}}^{\bf z})$ of type ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})$, respectively; see (3.6) and (3.9). In the special case where $p=q=0$, the specializations of the series $\phi_{\bf z}({\rm cdet}\,({\mathcal{L}}_{d}^{\bf w}))$ and $\overline{\phi}_{\bf w}({\rm Ber}^{\bf s}\,({\mathcal{L}}_{{p+m|q+n}}^{\bf z}))$ admit Capelli-type identities [HM, Proposition 5.4 and Proposition 5.6], and the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$ is obtained by matching the identities. However, the situation is somewhat complicated in our case as no such identities are known to exist. Based on the properties of Berezinians, we can relate $\overline{\phi}_{\bf w}({\rm Ber}^{\bf s}\,({\mathcal{L}}_{{p+m|q+n}}^{\bf z}))$ to $\phi_{\bf z}({\rm cdet}\,({\mathcal{L}}_{d}^{\bf w}))$ (for general $p,q,m,n$) by examining an anti-involution on the Weyl superalgebra $\mathcal{D}$ and an anti-involution on the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ (see Proposition 4.5). Theorem 1.1 will then follow.
We organize the paper as follows. In Section 2, we set up our notations. We recall the notions of the universal affine vertex algebra at the critical level, the Feigin–Frenkel center and the Bethe algebra. In Section 3, we discuss the Bethe algebras for general linear Lie (super)algebras and their basic properties. In Section 4, we introduce the joint action of the Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on the polynomial superalgebra $\mathcal{F}$ for any sequences ${\bf w}\in{\mathbb{C}}^{\ell}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$. We prove Theorem 1.1, i.e., the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$. In Section 5, we apply the Bethe duality to the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules and obtain Theorem 5.1.
Notations. Throughout the paper, the symbol ${\mathbb{Z}}$ (resp., ${\mathbb{N}}$ and ${\mathbb{Z}}_{+}$) stands for the set of all (resp., positive and non-negative) integers, the symbol ${\mathbb{C}}$ for the field of complex numbers, and the symbol ${\mathbb{Z}}_{2}:=\{\bar{0},\bar{1}\}$ for the field of integers modulo 2. All vector spaces, algebras, tensor products, etc., are over ${\mathbb{C}}$. We fix $d\in{\mathbb{N}}$ and $p,q,m,n\in{\mathbb{Z}}_{+}$.
2. Preliminaries
In this section, we set up our notations for the general linear Lie (super)algebras and their related concepts. We review several notions needed in this paper, including current algebras, the Feigin–Frenkel centers and the Bethe algebras.
2.1. The general linear Lie (super)algebras
For $p,q,m,n\in{\mathbb{Z}}_{+}$ that are not all zero, let
[…]
For $i\in\mathbb{I}$, define $|i|\in{\mathbb{Z}}_{2}$ by
$|i|=\begin{cases}\,\bar{0}&\,\,\mbox{if}\ \ i\in\{1,\ldots,p\}\cup\{p+q+1,\ldots,p+q+m\},\\ \,\bar{1}&\,\,\mbox{otherwise}.\end{cases}$
[…]
For any $i,j\in\mathbb{I}$, let $E^{i}_{j}$ denote the ${\mathbb{C}}$-linear endomorphism on ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ defined by
$E^{i}_{j}(e_{k})=\delta_{j,k}e_{i}\quad\mbox{ for $k\in\mathbb{I}$,}$
where $\delta$ is the Kronecker delta. The parity of $E^{i}_{j}$ is given by $|E^{i}_{j}|=|i|+|j|$. The superspace of ${\mathbb{C}}$-linear endomorphisms on ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ has a natural structure of a Lie superalgebra, called a general linear Lie (super)algebra and denoted by ${\mathfrak{gl}}_{p+m|q+n}$. The commutation relations of ${\mathfrak{gl}}_{p+m|q+n}$ are given by
$[E^{i}_{j},E^{k}_{\ell}]=\delta_{j,k}E^{i}_{\ell}-(-1)^{(|i|+|j|)(|k|+|\ell|)}\delta_{i,\ell}E^{k}_{j}\qquad\mbox{for $i,j,k,\ell\in\mathbb{I}$.}$
Note that $\{E^{i}_{j}\,|\,i,j\in\mathbb{I}\}$ is a homogeneous basis for ${\mathfrak{gl}}_{p+m|q+n}$.
Let $\mathfrak{b}_{{p+m|q+n}}=\bigoplus_{\begin{subarray}{c}i,j\in\mathbb{I},i\leq j\end{subarray}}{\mathbb{C}}E^{i}_{j}$ be a Borel subalgebra of ${\mathfrak{gl}}_{p+m|q+n}$. The corresponding Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ has a basis $\{E^{i}_{i}\,|\,i\in\mathbb{I}\}$, and the dual basis in ${\mathfrak{h}}_{{p+m|q+n}}^{*}$ is denoted by $\{\epsilon_{i}\,|\,i\in\mathbb{I}\}$, where the parity of $\epsilon_{i}$ is given by $|\epsilon_{i}|=|i|$. Note that the Borel subalgebra $\mathfrak{b}_{{p+m|q+n}}$ is not the standard one unless some of $p,q,m,n$ are set to 0.
For $m=d$ and $p=q=n=0$, we write
${\mathfrak{gl}}_{d}={\mathfrak{gl}}_{d|0},\quad{\mathfrak{b}}_{d}={\mathfrak{b}}_{d|0},\quad{\mathfrak{h}}_{d}={\mathfrak{h}}_{d|0},$
and
$\hskip 28.45274pte_{ij}=E^{i}_{j}\qquad\mbox{for $i,j=1,\ldots,d$.}$
2.2. Current algebras and evaluation modules
For any Lie (super)algebra ${\mathfrak{g}}$ and an even variable $t$, let ${\mathfrak{g}}[t,t^{-1}]:={\mathfrak{g}}\otimes{\mathbb{C}}[t,t^{-1}]$ denote the loop algebra of ${\mathfrak{g}}$, which is a Lie superalgebra with commutation relations
$\hskip 28.45274pt\big{[}A\otimes t^{r},B\otimes t^{s}\big{]}=[A,B]\otimes t^{r+s}\qquad\mbox{for $A,B\in{\mathfrak{g}}$ and $r,s\in{\mathbb{Z}}$.}$
Here $[A,B]$ is the supercommutator of $A$ and $B$. We will use the notation
$\hskip 28.45274ptA[r]:=A\otimes t^{r}\qquad\mbox{for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{Z}}$.}$
The current algebra ${\mathfrak{g}}[t]:={\mathfrak{g}}\otimes{\mathbb{C}}[t]$ is a subalgebra of ${\mathfrak{g}}[t,t^{-1}]$ defined in an obvious way. We identify the Lie (super)algebra ${\mathfrak{g}}$ with the subalgebra ${\mathfrak{g}}\otimes 1$ of constant polynomials in ${\mathfrak{g}}[t]$ and hence $U({\mathfrak{g}})\subseteq U({\mathfrak{g}}[t])$. Hereafter $U(\mathcal{G})$ denotes the universal enveloping algebra of $\mathcal{G}$ for any Lie (super)algebra $\mathcal{G}$.
For $\alpha\in{\mathbb{C}}$, we have an evaluation homomorphism ${\rm ev}_{\alpha}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}})$ given by
${\rm ev}_{\alpha}(A[r])=\alpha^{r}A\qquad\mbox{for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{Z}}_{+}$}.$
Any ${\mathfrak{g}}$-module $M$ is a ${\mathfrak{g}}[t]$-module via ${\rm ev}_{\alpha}$, which is called the evaluation module and denoted by $M(\alpha)$. For any $A\in{\mathfrak{g}}$ and any even variable $z$, we set
$A(z)=\sum_{r=0}^{\infty}A[r]z^{-r-1}.$
It is a formal power series in $z^{-1}$ with coefficients in ${\mathfrak{g}}[t]$ and acts on $M(\alpha)$ by
$A(z)v=\frac{Av}{z-\alpha}\qquad\mbox{for $v\in M(\alpha)$.}$
Here every rational function in $z$ represents its power series expression at $\infty$.
Let $\ell\in{\mathbb{N}}$ and ${\bf z}=(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}$. Let $\Delta^{(\ell-1)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}}[t])^{\otimes\ell}$ be the $(\ell-1)$-fold coproduct on $U({\mathfrak{g}}[t])$. We have the homomorphism
$\underline{{\rm ev}}_{\bf z}:=({\rm ev}_{z_{1}}\otimes\ldots\otimes{\rm ev}_{z_{\ell}})\circ\Delta^{(\ell-1)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}})^{\otimes{\ell}},$
given by
$\underline{{\rm ev}}_{\bf z}(A[r])=\sum_{i=1}^{\ell}z_{i}^{r}A^{(i)}\qquad\mbox{for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{Z}}_{+}$}.$
Here $A^{(i)}:=\underbrace{1\otimes\cdots\otimes 1\otimes\stackrel{{\scriptstyle i}}{{A}}\otimes 1\otimes\cdots\otimes 1}_{\ell}$ for $i=1,\ldots,\ell$. We call $\underline{{\rm ev}}_{\bf z}$ the evaluation map at ${\bf z}$.
For any ${\mathfrak{g}}$-modules $M_{1},\ldots,M_{\ell}$, let
$\underline{M\mkern-4.0mu}\mkern 4.0mu=M_{1}\otimes\cdots\otimes M_{\ell}.$ (2.1)
The tensor product $\underline{M\mkern-4.0mu}\mkern 4.0mu$ is a ${\mathfrak{g}}[t]$-module via $\underline{{\rm ev}}_{\bf z}$, which coincides with the tensor product of evaluation ${\mathfrak{g}}[t]$-modules
$\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z}):=M_{1}(z_{1})\otimes\cdots\otimes M_{\ell}(z_{\ell}).$ (2.2)
As ${\mathfrak{g}}$-modules, $\underline{M\mkern-4.0mu}\mkern 4.0mu$ and $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ are isomorphic via the identity map. For any $A\in{\mathfrak{g}}$, the series $A(z)$ acts on $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ by
$\hskip 42.67912ptA(z)v=\sum_{i=1}^{\ell}\frac{A^{(i)}v}{z-z_{i}}\qquad\mbox{for $v\in\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$}.$
2.3. The Feigin–Frenkel center
The Lie algebra ${\mathfrak{gl}}_{d}$ admits a nondegenerate invariant symmetric bilinear form
$(A,B)={\rm tr}(AB)-\frac{1}{d}{\rm tr}(A){\rm tr}(B),\quad\mbox{for $A,B\in{\mathfrak{gl}}_{d}$},$
where ${\rm tr}$ denotes the trace. The affine Lie superalgebra
$\widehat{{\mathfrak{gl}}_{d}}:={\mathfrak{gl}}_{d}[t,t^{-1}]\oplus{\mathbb{C}}K,$
where $K$ is even and central, is a Lie superalgebra with commutation relations
$\big{[}A[r],B[s]\big{]}=[A,B][r+s]+r\delta_{r+s,0}(A,B)K$
for $A,B\in{\mathfrak{gl}}_{d}$ and $r,s\in{\mathbb{Z}}$.
The vacuum module $V_{{\rm crit}}({\mathfrak{gl}}_{d})$ at the critical level is
$V_{{\rm crit}}({\mathfrak{gl}}_{d}):=U(\widehat{{\mathfrak{gl}}_{d}})/I,$
where $I$ is the left ideal of $U(\widehat{{\mathfrak{gl}}_{d}})$ generated by ${\mathfrak{gl}}_{d}[t]$ and $K+d$. It possesses a unique vertex algebra structure, called the universal affine vertex algebra associated to ${\mathfrak{gl}}_{d}$ at the critical level; see [FBZ, Theorem 4.4.1]. The center of the vertex algebra $V_{{\rm crit}}({\mathfrak{gl}}_{d})$ is given by
${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d}):=\left\{\,\!v\in V_{{\rm crit}}({\mathfrak{gl}}_{d})\;\middle|\;{\mathfrak{gl}}_{d}[t]v=0\!\,\right\},$
called the Feigin–Frenkel center. Any element of ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$ is called a Segal–Sugawara vector.
We also have the universal affine vertex algebra $V_{{\rm crit}}({\mathfrak{gl}}_{p+m|q+n})$ at the critical level and the Feigin–Frenkel center ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$ for ${\mathfrak{gl}}_{p+m|q+n}$. We skip details here; see [MR, Section 1.3] (and also [ChL2, Section 3.1]).
Let ${\mathfrak{g}}$ be a general linear Lie (super)algebra. The algebra $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ is equipped with the (even) derivation ${\rm T}:=-d/dt$ defined by
${\rm T}(1)=0\qquad\text{and}\qquad{\rm T}(A[-r])=rA[-r-1]$
for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{N}}$. Note that $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ is isomorphic to $V_{{\rm crit}}({\mathfrak{g}})$ as superspaces by the Poincaré–Birkhoff–Witt theorem, and that the derivation ${\rm T}$ corresponds to the translator operator on $V_{{\rm crit}}({\mathfrak{g}})$. It is known that ${\mathfrak{z}}(\widehat{{\mathfrak{g}}})$ can be viewed as a commutative subalgebra of $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ and is ${\rm T}$-invariant. We refer the reader to [Fr, FBZ, Mo, MR] for further details.
2.4. Bethe algebras
For $\mu\in{\mathfrak{gl}}_{d}^{*}$, there is an algebra homomorphism
$\Psi^{\mu}:U(t^{-1}{\mathfrak{gl}}_{d}[t^{-1}])\longrightarrow U({\mathfrak{gl}}_{d}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$
given by
$\hskip 28.45274pt\Psi^{\mu}(A[-r])=A\otimes(t-z)^{-r}+\delta_{r,1}\mu(A),\quad\mbox{for $A\in{\mathfrak{gl}}_{d}\,$ and $\,r\in{\mathbb{N}}$.}$ (2.3)
The Bethe algebra $\mathcal{B}_{d}^{\mu}$ of ${\mathfrak{gl}}_{d}$ with respect to $\mu$ is defined to be the subalgebra of $U({\mathfrak{gl}}_{d}[t])$ generated by the coefficients of $\Psi^{\mu}(S)$ for $S\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$. Since ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$ is commutative, so is $\mathcal{B}_{d}^{\mu}$ ([FFTL, Ry]).
For any $\mathcal{B}_{d}^{\mu}$-module $V$, let $(\mathcal{B}_{d}^{\mu})_{V}$ denote the image of the Bethe algebra $\mathcal{B}_{d}^{\mu}$ in the endomorphism algebra ${\rm End}(V)$ of $V$. We call $(\mathcal{B}_{d}^{\mu})_{V}$ the Bethe algebra of $V$ with respect to $\mu$. For $\alpha\in{\mathbb{C}}$, any ${\mathfrak{gl}}_{d}$-module $M$ is the evaluation module $M(\alpha)$ over ${\mathfrak{gl}}_{d}[t]$, and so $\mathcal{B}_{d}^{\mu}$ acts on $M(\alpha)$. If the centralizer ${\mathfrak{gl}}_{d}^{\mu}$ of $\mu$ in ${\mathfrak{gl}}_{d}$ contains the Cartan subalgebra ${\mathfrak{h}}_{d}$, then $\mathcal{B}_{d}^{\mu}$ commutes with ${\mathfrak{h}}_{d}$ and hence acts on any $\gamma$-weight space $M(\alpha)_{\gamma}$ of $M(\alpha)$, where $\gamma\in{\mathfrak{h}}_{d}^{*}$ ([Ry, Proposition 4]).
Let
${\boldsymbol{X}}_{\!\ell}=\left\{\,\!(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}\;\middle|\;z_{i}\not=z_{j}\,\,\mbox{for any $i\not=j$}\!\,\right\}$
be the configuration space of $\ell$ distinct points on ${\mathbb{C}}^{\ell}$, and let ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$. Let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible ${\mathfrak{gl}}_{d}$-modules. The Bethe algebra $\mathcal{B}_{d}^{\mu}\subseteq U({\mathfrak{gl}}_{d}[t])$ acts on $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ (for the notation, see (2.2)) via the evaluation map $\underline{{\rm ev}}_{\bf z}$.
An element $\mu\in{\mathfrak{gl}}_{d}^{*}$ is called regular if the dimension of ${\mathfrak{gl}}_{d}^{\mu}$ is $d$. The following remarkable results for $\mathcal{B}_{d}^{\mu}$, due to Feigin, Frenkel, and Rybnikov, will be used in this paper.
Theorem 2.1 (cf. [FFRy, Corollary 5]).
For any regular $\mu\in{\mathfrak{gl}}_{d}^{*}$ and ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ is a cyclic $\mathcal{B}_{d}^{\mu}$-module.
Theorem 2.2 (cf. [FFRy, Corollary 6]).
For generic $\mu\in{\mathfrak{gl}}_{d}^{*}$ and generic ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, the Bethe algebra $(\mathcal{B}_{d}^{\mu})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is diagonalizable with a simple spectrum.
3. Bethe algebras for general linear Lie (super)algebras
In this section, we recall column determinants, Berezinians and pseudo-differential operators. We discuss the Bethe algebras of the general linear Lie (super)algebras and their properties.
3.1. Column determinants and Berezinians
Let $\mathcal{A}$ be an associative unital superalgebra over ${\mathbb{C}}$. The parity of a homogeneous element $a\in\mathcal{A}$ is denoted by $|a|$, which lies in ${\mathbb{Z}}_{2}$. Fix $k\in{\mathbb{N}}$. For any $k\times k$ matrix $A=\big{[}a_{i,j}\big{]}_{i,j=1,\ldots,k}$ over $\mathcal{A}$, the column determinant of $A$ is defined to be
${\rm cdet}\,A=\sum_{\sigma\in\mathfrak{S}_{k}}(-1)^{l(\sigma)}\,a_{\sigma(1),1}\ldots a_{\sigma(k),k}.$
Here $\mathfrak{S}_{k}$ denotes the symmetric group on $\{1,\ldots,k\}$ and $l(\sigma)$ denotes the length of $\sigma$. For any nonempty subset $P=\{i_{1}<\ldots<i_{\ell}\}$ of $\{1,\ldots,k\}$, the matrix $A_{P}:=\big{[}a_{i,j}\big{]}_{i,j\in P}$ is called a standard submatrix of $A$.
Assume that $A$ has a two-sided inverse $A^{-1}=\big{[}\widetilde{a}_{i,j}\big{]}$. For all $i,j=1,\ldots,k$, the $(i,j)$th quasideterminant of $A$ is defined to be $|A|_{ij}:=\widetilde{a}_{j,i}^{-1}$ provided that $\widetilde{a}_{j,i}$ has an inverse in $\mathcal{A}$. According to [GGRW, Proposition 1.2.6], if $k\geq 2$, then
$|A|_{ij}=a_{i,j}-r_{i}^{j}(A^{ij})^{-1}c_{j}^{i},$
where $A^{ij}$ is the $(k-1)\times(k-1)$ submatrix of $A$ obtained by deleting the $i$th row and the $j$th column of $A$ and is assumed to be invertible, $r_{i}^{j}$ is the $i$th row of $A$ with $a_{i,j}$ removed, and $c_{j}^{i}$ is the $j$th column of $A$ with $a_{i,j}$ removed. Following the notation of [GGRW], it is convenient to write
$|A|_{ij}=\begin{vmatrix}a_{1,1}&\ldots&a_{1,j}&\ldots&a_{1,k}\\ \ldots&\ldots&\ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,j}$}&\ldots&a_{i,k}\\ \ldots&\ldots&\ldots&\ldots&\ldots\\ a_{k,1}&\ldots&a_{k,j}&\ldots&a_{k,k}\end{vmatrix}.$
For $i=1,\ldots,k$, we define
$d_{i}(A)=\begin{vmatrix}a_{1,1}&\ldots&a_{1,i}\\ \ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,i}$}\end{vmatrix},$
which are called the principal quasiminors of $A$.
A $k\times k$ matrix $A$ over $\mathcal{A}$ is called sufficiently invertible if every principal quasiminor of $A$ is well defined, and $A$ is called amply invertible if each of its standard submatrices is sufficiently invertible.
Let ${{\bf s}}=(s_{1},\ldots,s_{m+n})$ be a sequence of 0’s and 1’s such that exactly $m$ of the $s_{i}$’s are 0 and the others are 1. We call such a sequence a $0^{m}1^{n}$-sequence. Every $0^{m}1^{n}$-sequence can be written in the form $(0^{m_{1}},1^{n_{1}},\ldots,0^{m_{r}},1^{n_{r}}),$ where the sequence begins with $m_{1}$ copies of 00’s, followed by $n_{1}$ copies of $1$’s, and so on. The set of all $0^{m}1^{n}$-sequences is denoted by $\mathcal{S}_{{m|n}}$.
Let ${\bf s}\in\mathcal{S}_{{m|n}}$. For any $\sigma\in\mathfrak{S}_{m+n}$ and any $(m+n)\times(m+n)$ matrix $A=[a_{i,j}]$ over $\mathcal{A}$, we define ${\bf s}^{\sigma}=\left(s_{\sigma^{-1}(1)},s_{\sigma^{-1}(2)},\ldots,s_{\sigma^{-1}(m+n)}\right)$ and $A^{\sigma}=\big{[}a_{\sigma^{-1}(i),\sigma^{-1}(j)}\big{]}$. We say that $A$ is of type ${\bf s}$ if $a_{i,j}$ is a homogeneous element of parity $|a_{i,j}|=\bar{s}_{i}+\bar{s}_{j}$ for any $i,j=1,\ldots,m+n$.
For any $(m+n)\times(m+n)$ sufficiently invertible matrix $A$ of type ${\bf s}$ over $\mathcal{A}$, the Berezinian of type ${\bf s}$ of $A$ is defined to be
${\rm Ber}^{\bf s}\,A=d_{1}(A)^{\hat{s}_{1}}\ldots d_{m+n}(A)^{\hat{s}_{m+n}},$
where $\hat{s}_{i}:=(-1)^{s_{i}}$ for $1\leq i\leq m+n$ (see [HM, (3.3)]). We refer the reader to [Ber, Na, MR] for earlier definitions of Berezinians. We have the following proposition.
Proposition 3.1 (cf. [HM, Proposition 3.5]).
Let $A$ be an $(m+n)\times(m+n)$ amply invertible matrix of type ${\bf s}$ over $\mathcal{A}$. Fix $k\in\{1,\ldots,m+n-1\}$. We write
$A=\begin{bmatrix}W&X\\ Y&Z\end{bmatrix},$ (3.1)
where $W,X,Y,Z$ are respectively $k\times k$, $k\times(m+n-k)$, $(m+n-k)\times k$, and $(m+n-k)\times(m+n-k)$ matrices. Then $W$ and $Z-YW^{-1}X$ are sufficiently invertible matrices of types ${\bf s}^{\prime}:=(s_{1},\ldots,s_{k})$ and ${\bf s}^{\prime\prime}:=(s_{k+1},\ldots,s_{m+n})$, respectively. Moreover,
${\rm Ber}^{\bf s}\,A={\rm Ber}^{{\bf s}^{\prime}}W\cdot{\rm Ber}^{{\bf s}^{\prime\prime}}\big{(}Z-YW^{-1}X\big{)}.$
An $(m+n)\times(m+n)$ matrix $A=\big{[}a_{i,j}\big{]}$ over $\mathcal{A}$ is called a Manin matrix of type ${\bf s}$ if $A$ is a matrix of type ${\bf s}$ satisfying the following relations
$[a_{i,j},a_{k,l}]=(-1)^{s_{i}s_{j}+s_{i}s_{k}+s_{j}s_{k}}[a_{k,j},a_{i,l}]\qquad\text{for all}\quad i,j,k,l=1,\ldots,m+n,$
where $[a,b]:=ab-(-1)^{|a||b|}ba$ for any homogeneous elements $a,b\in\mathcal{A}$. We recall some basic facts about Manin matrices.
Proposition 3.2 (cf. [HM, Section 3]).
Let $A$ be an $(m+n)\times(m+n)$ Manin matrix of type ${\bf s}$ over $\mathcal{A}$. Then
  1. (i)
    $A^{-1}$ is an $(m+n)\times(m+n)$ Manin matrix of type ${\bf s}$ over $\mathcal{A}$.
  2. (ii)
    If $P=\{i_{1}<\ldots<i_{\ell}\}$ is a nonempty subset of $\{1,\ldots,m+n\}$, then the standard submatrix $A_{P}$ of $A$ is a Manin matrix of type ${{\bf s}}_{P}:=(s_{i_{1}},\ldots s_{i_{\ell}})$.
  3. (iii)
    For any $\sigma\in\mathfrak{S}_{m+n}$, $A^{\sigma}$ is a Manin matrix of type ${\bf s}^{\sigma}$.
The following proposition asserts that the Berezinian of a Manin matrix is invariant under the action of a permutation.
Proposition 3.3 ([HM, Proposition 3.6]).
Let $A$ be an $(m+n)\times(m+n)$ amply invertible Manin matrix of type ${\bf s}$ over $\mathcal{A}$. Then
$\hskip 28.45274pt{\rm Ber}^{{\bf s}^{\sigma}}A^{\sigma}={\rm Ber}^{\bf s}\,A\qquad\mbox{for any $\sigma\in\mathfrak{S}_{m+n}$.}$
We may decompose ${\rm Ber}^{\bf s}\,A$ in Proposition 3.1 in a different way if $A$ is a Manin matrix.
Proposition 3.4.
Let $A$ be an $(m+n)\times(m+n)$ amply invertible Manin matrix of type ${\bf s}$ over $\mathcal{A}$. Write $A$ as in (3.1). Then $W-XZ^{-1}Y$ and $Z$ are sufficiently invertible Manin matrices of types ${\bf s}^{\prime}:=(s_{1},\ldots,s_{k})$ and ${\bf s}^{\prime\prime}:=(s_{k+1},\ldots,s_{m+n})$, respectively. Moreover,
${\rm Ber}^{\bf s}\,A={\rm Ber}^{{\bf s}^{\prime\prime}}Z\cdot{\rm Ber}^{{\bf s}^{\prime}}\big{(}W-XZ^{-1}Y\big{)}.$ (3.2)
Proof.
The first assertion is clear from Proposition 3.2. Let $\sigma\in\mathfrak{S}_{m+n}$ be the permutation defined by $\sigma(k+i)=i$ for $i=1,\ldots,m+n-k$ and $\sigma(j)=m+n-k+j$ for $j=1,\ldots,k$. Then
$A^{\sigma}=\begin{bmatrix}Z&Y\\ X&W\end{bmatrix}.$
The equality (3.2) is now a direct consequence of Proposition 3.1 and Proposition 3.3. ∎
Proposition 3.4 will be used in Section 4.2. We also have the following proposition.
Proposition 3.5 ([CFR, Lemma 8]).
Let ${\bf s}=(0^{m})$. For any $m\times m$ sufficiently invertible Manin matrix $A$ of type ${\bf s}$ over $\mathcal{A}$, we have ${\rm Ber}^{\bf s}\,A={\rm cdet}\,\,A$.
3.2. Pseudo-differential operators
Let $\mathcal{A}$ be an associative unital superalgebra over ${\mathbb{C}}$ and $z$ an even variable. We denote by $\mathcal{A}[\kern-1.1pt[z]\kern-1.1pt]$ and $\mathcal{A}(\!(z)\!)$ the superalgebras of formal power series and Laurent series in $z$ with coefficients in $\mathcal{A}$, respectively. Let $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ denote the set of all formal series of the form
$\sum_{\ell=-\infty}^{s}\sum_{k=-\infty}^{r}a_{k\ell}z^{k}\partial_{z}^{\ell},$
where $r,s\in{\mathbb{Z}}$ and $a_{k\ell}\in\mathcal{A}$. We may endow $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ with a superalgebra structure using the rules:
$\partial_{z}\partial_{z}^{-1}=\partial_{z}^{-1}\partial_{z}=1,\quad\partial_{z}^{k}z^{\ell}=\sum_{i=0}^{\infty}\binom{k}{i}\binom{\ell}{i}i!\,z^{\ell-i}\,\partial_{z}^{k-i},\quad\mbox{for $k,\ell\in{\mathbb{Z}}$.}$ (3.3)
Here, e.g., ${\binom{k}{i}:=\frac{k(k-1)\ldots(k-i+1)}{i!}}$. The superalgebra $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ is called the superalgebra of pseudo-differential operators over $\mathcal{A}$.
Let $\omega$ be an anti-involution on $\mathcal{A}$. In particular, $\omega^{2}=1$ and $\omega(ab)=\omega(b)\omega(a)$ for $a,b\in\mathcal{A}$. We readily see that $\omega$ extends to a ${\mathbb{C}}$-linear automorphism $\omega:\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ defined by
$\omega(az^{k}\partial_{z}^{\ell})=\omega(a)z^{\ell}\partial_{z}^{k}$ (3.4)
for $a\in\mathcal{A}$ and $k,\ell\in{\mathbb{Z}}$.
Proposition 3.6.
The map $\omega$ is an anti-involution on $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$.
Proof.
Clearly, $\omega^{2}=1$. It remains to show that $\omega$ is an anti-homomorphism. For $a,b\in\mathcal{A}$ and $k,\ell,r,s\in{\mathbb{Z}}$, we have
$\omega((az^{k}\partial_{z}^{\ell})(bz^{r}\partial_{z}^{s}))$ $=$ $\omega(abz^{k}\partial_{z}^{\ell}z^{r}\partial_{z}^{s})$
$=$ $\sum_{i=0}^{\infty}\binom{\ell}{i}\binom{r}{i}i!\,\omega(abz^{k+r-i}\partial_{z}^{\ell+s-i})$
$=$ $\sum_{i=0}^{\infty}\binom{\ell}{i}\binom{r}{i}i!\,\omega(ab)z^{\ell+s-i}\partial_{z}^{k+r-i}$
$=$ $\omega(b)\omega(a)z^{s}\partial_{z}^{r}z^{\ell}\partial_{z}^{k}$
$=$ $\left(\omega(b)z^{s}\partial_{z}^{r}\right)\left(\omega(a)z^{\ell}\partial_{z}^{k}\right)$
$=$ $\omega(bz^{r}\partial_{z}^{s})\,\omega(az^{k}\partial_{z}^{\ell}),$
and the proposition follows. ∎
3.3. The Bethe algebra for ${\mathfrak{gl}}_{d}$ and Frobenius algebras
Let
${\bf w}=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{\ell}.$
Define $\mu_{\bf w}\in{\mathfrak{gl}}_{d}^{*}$ by
$\hskip 28.45274pt\mu_{\bf w}(e_{ab})=-\delta_{a,b}w_{a}\qquad\mbox{for $1\leq a\leq d$.}$
The choice of the minus sign will be clear below.
Recall the map $\Psi^{\mu_{\bf w}}$ defined in (2.3). We write
$\Psi^{\bf w}=\Psi^{\mu_{\bf w}}.$
Let $t$ be an even variable and
$\tau=-\partial_{t}.$
Here $t$ and $\partial_{t}$ satisfy the rules similar to (3.3). The map $\Psi^{\bf w}$ extends to an algebra homomorphism
$\Psi^{\bf w}:U(t^{-1}{\mathfrak{gl}}_{d}[t^{-1}])[\tau]\longrightarrow U({\mathfrak{gl}}_{d}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt][\partial_{z}]$
such that $\Psi^{\bf w}(\tau)=\partial_{z}$. Let
$\mathcal{T}_{d}=\Big{[}\delta_{a,b}\tau+e_{ab}[-1]\Big{]}_{a,b=1,\ldots,d},$
which is a $d\times d$ Manin matrix. We have
${\rm cdet}\,(\mathcal{T}_{d})=\tau^{d}+\sum_{i=1}^{d}S_{i}\tau^{i-1}.$ (3.5)
By [CM, Theorem 3.1] (see also [CT]), $S_{1},\ldots,S_{d}$ form a complete set of Segal–Sugawara vectors for the Feigin–Frenkel center ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$. That is, the set $\{{\rm T}^{r}S_{i}\,|\,i=1,\ldots,d,\,r\in{\mathbb{Z}}_{+}\}$ is algebraically independent, and
${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})={\mathbb{C}}[{\rm T}^{r}S_{i}\,|\,i=1,\ldots,d,\,r\in{\mathbb{Z}}_{+}].$
Now consider the $d\times d$ Manin matrix
${\mathcal{L}}_{d}^{\bf w}:=\Big{[}\Psi^{\bf w}\big{(}\delta_{a,b}\tau+e_{ab}[-1]\big{)}\Big{]}_{a,b=1,\ldots,d}=\Big{[}\delta_{a,b}(\partial_{z}-w_{a})-e_{ab}(z)\Big{]}_{a,b=1,\ldots,d}.$
Then in view of (3.5),
${\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})=\partial_{z}^{d}+\sum_{i=1}^{d}S_{i}^{{\bf w}}(z)\partial_{z}^{i-1},$ (3.6)
where $S_{i}^{{\bf w}}(z):=\Psi^{\bf w}(S_{i})$.
The following is well-known and matches the definition of the Bethe algebra in [MTV1].
Proposition 3.7.
The Bethe algebra $\mathcal{B}_{d}^{\bf w}$ is generated by the coefficients of $S_{i}^{{\bf w}}(z)$ for $i=1,\ldots,d$.
A Frobenius algebra $\mathcal{A}$ is a finite-dimensional commutative associative unital algebra which admits a nondegenerate symmetric bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) such that
$\hskip 28.45274pt(a_{1}a_{2},a_{3})=(a_{1},a_{2}a_{3})\quad\mbox{for all $a_{1},a_{2},a_{3}\in\mathcal{A}$.}$
Recall the notations (2.1) and (2.2). Let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible ${\mathfrak{gl}}_{d}$-modules, and let ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$. As ${\mathfrak{h}}_{d}\subseteq{\mathfrak{gl}}_{d}^{\mu_{\bf w}}$, the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ acts on any weight space of $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$. We will explain why $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a Frobenius algebra for ${\bf w}\in{\boldsymbol{X}}_{\!d}$.
Define an anti-involution $\boldsymbol{\theta}:{\mathfrak{gl}}_{d}\longrightarrow{\mathfrak{gl}}_{d}$ by
$\boldsymbol{\theta}(e_{ab})=e_{ba}\qquad\mbox{for $1\leq a,b\leq d$.}$ (3.7)
The map $\boldsymbol{\theta}$ is called the Cartan anti-involution on ${\mathfrak{gl}}_{d}$.
For $i=1,\ldots,\ell$, $V_{i}$ is an irreducible highest weight ${\mathfrak{gl}}_{d}$-module with a dominant integral highest weight. Let $v_{i}$ be a highest weight vector of $V_{i}$. There is a unique nondegenerate symmetric bilinear form $(\cdot,\cdot)_{i}$ on $V_{i}$ defined by
$(v_{i},v_{i})_{i}=1\quad\text{and}\quad(Av,v^{\prime})_{i}=(v,\boldsymbol{\theta}(A)v^{\prime})_{i}$
for all $v,v^{\prime}\in V_{i}$ and $A\in{\mathfrak{gl}}_{d}$ ([Sh]). The bilinear form $(\cdot,\cdot)_{i}$ is called the Shapovalov form on $V_{i}$. It is easy to see that if $v$ and $v^{\prime}$ are weight vectors of $V_{i}$ of distinct weights, then $(v,v^{\prime})_{i}=0$.
Let (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denote the tensor form of the bilinear forms $(\cdot,\cdot)_{1},\ldots,(\cdot,\cdot)_{\ell}$, which is called the tensor Shapovalov form. It is a nondegenerate symmetric bilinear form on $\underline{V\mkern-4.0mu}\mkern 4.0mu$. Let ${\bf w}\in{\mathbb{C}}^{d}$. By [MTV1, Theorem 9.1] and [Lu, Lemma 2.6], the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ is symmetric with respect to (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ), i.e.,
$\hskip 42.67912pt({\boldsymbol{b}}v,v^{\prime})=(v,{\boldsymbol{b}}v^{\prime})\quad\mbox{for ${\boldsymbol{b}}\in\mathcal{B}_{d}^{\bf w}$ and $v,v^{\prime}\in\underline{V\mkern-4.0mu}\mkern 4.0mu$}.$
Theorem 3.8.
For any ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, the Bethe algebra $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a Frobenius algebra. Moreover, $(\mathcal{B}_{d}^{\bf w})_{{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\gamma}}$ is a Frobenius algebra for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
Proof.
The first statement follows from the above discussion on the tensor Shapovalov form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on $\underline{V\mkern-4.0mu}\mkern 4.0mu$ in view of Theorem 2.1 and [Lu, Lemma 2.7]. As (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) restricts to a nondegenerate symmetric bilinear form on ${\underline{V\mkern-4.0mu}\mkern 4.0mu}_{\!\gamma}$ for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$, the second statement follows. ∎
The following is a consequence of [Lu, Lemma 1.3].
Theorem 3.9.
For any ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, the following properties hold:
  1. (i)
    The Bethe algebra $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a maximal commutative subalgebra of ${\rm End}(\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z}))$ of dimension $\dim\!\left({\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}\right)$.
  2. (ii)
    Every eigenspace of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is one-dimensional, and the set of eigenspaces of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is in bijective correspondence with the set of maximal ideals of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$.
  3. (iii)
    Every generalized eigenspace of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a cyclic $\mathcal{B}_{d}^{\bf w}$-module.
Remark 3.10.
Each statement of Theorem 3.9 holds if $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ is replaced with ${\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\gamma}$ for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
3.4. The Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$
Fix ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. Let
$\mathcal{T}_{p+m|q+n}=\Big{[}\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}[-1]\Big{]}_{i,j\in\mathbb{I}}.$
It is straightforward to verify that $\mathcal{T}_{p+m|q+n}$ is an amply invertible $(p+q+m+n)\times(p+q+m+n)$ Manin matrix of type ${\bf s}$ over $U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])(\!(\tau^{-1})\!)$. The verification is analogous to [MR, Lemma 3.1]. The following is an immediate consequence of Proposition 3.3.
Proposition 3.11.
For each $\sigma\in\mathfrak{S}_{p+q+m+n}$,
${\rm Ber}\,^{{\bf s}^{\sigma}}\!\left(\mathcal{T}_{p+m|q+n}^{\sigma}\right)={\rm Ber}^{\bf s}\,\!\left(\mathcal{T}_{p+m|q+n}\right).$
We have the expansion
${\rm Ber}^{\bf s}\,\!\left(\mathcal{T}_{p+m|q+n}\right)=\sum_{i=-\infty}^{p+m-q-n}b_{i}\tau^{i},$ (3.8)
for some $b_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$. According to [MR, Corollary 3.3] (see also [ChL2, Proposition 3.3]),
$b_{i}\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$
for all $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$.
Now let
${\bf z}=(z_{1},\ldots,z_{p+q+m+n})\in{\mathbb{C}}^{p+q+m+n}.$
Define $\mu_{\bf z}\in{\mathfrak{gl}}_{p+m|q+n}^{*}$ by
$\hskip 28.45274pt\mu_{\bf z}(E^{i}_{j})=(-1)^{|i|+1}\delta_{i,j}z_{i}\qquad\mbox{for $i,j\in\mathbb{I}$.}$
Similar to (2.3), we have a superalgebra homomorphism
$\Psi^{\bf z}:U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])\longrightarrow U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1})\!)$
given by
$\Psi^{\bf z}(A[-r])=A\otimes(t-z)^{-r}+\delta_{r,1}\mu_{\bf z}(A),\quad\mbox{for $A\in{\mathfrak{gl}}_{p+m|q+n}\,$ and $\,r\in{\mathbb{N}}$.}$
The map $\Psi^{\bf z}$ extends to a superalgebra homomorphism
$\Psi^{\bf z}:U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])(\!(\tau^{-1})\!)\longrightarrow U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)$
given by
$\Psi^{\bf z}\left(\sum_{i=-\infty}^{s}a_{i}\tau^{i}\right)=\sum_{i=-\infty}^{s}\Psi^{\bf z}(a_{i})\partial_{z}^{i}$
for $a_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$ and $s\in{\mathbb{Z}}$.
Now consider the $(p+q+m+n)\times(p+q+m+n)$ matrix
${\mathcal{L}}_{{p+m|q+n}}^{\bf z}:=\Big{[}\Psi^{\bf z}\big{(}\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}[-1]\big{)}\Big{]}_{i,j\in\mathbb{I}}=\Big{[}\delta_{i,j}(\partial_{z}-z_{i})-(-1)^{|i|}E^{i}_{j}(z)\Big{]}_{i,j\in\mathbb{I}}.$
Evidently, ${\mathcal{L}}_{{p+m|q+n}}^{\bf z}$ is an amply invertible Manin matrix of type ${\bf s}$ over the superalgebra $U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)$ of pseudo-differential operators over $U({\mathfrak{gl}}_{p+m|q+n}[t])$. Just like Proposition 3.11, we have the following.
Proposition 3.12.
For each $\sigma\in\mathfrak{S}_{p+q+m+n}$,
${\rm Ber}\,^{{\bf s}^{\sigma}}\big{(}\big{(}{\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big{)}^{\sigma}\big{)}={\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right).$
By (3.8), we have
${\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)=\sum_{i=-\infty}^{p+m-q-n}b_{i}^{{\bf z}}(z)\partial_{z}^{i},$ (3.9)
where $b_{i}^{{\bf z}}(z):=\Psi^{\bf z}(b_{i})$. Let ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ be the subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$ generated by the coefficients of the series $b_{i}^{{\bf z}}(z)$, for $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$. It is called the Bethe algebra of ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}$. If we take $q=n=0$, we recover the Bethe algebra of ${\mathfrak{gl}}_{p+m}$ by virtue of Proposition 3.5. According to [MR, Corallary 3.7], the algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ is a commutative subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$. It also commutes with the Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ of ${\mathfrak{gl}}_{p+m|q+n}$ and hence acts on any weight space of a ${\mathfrak{gl}}_{p+m|q+n}$-module.
For any ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module $V$, let $\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}_{V}$ denote the image of the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ in ${\rm End}(V)$. We call $\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}_{V}$ the Bethe algebra of $V$ with respect to ${\bf z}$.
4. Duality for Bethe algebras
This section is devoted to establishing a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ between the Bethe algebras $\mathcal{B}_{d}^{\bf w}\subseteq U({\mathfrak{gl}}_{d}[t])$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}\subseteq U({\mathfrak{gl}}_{p+m|q+n}[t])$ for any ${\bf w}\in{\mathbb{C}}^{d}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$.
4.1. The actions of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$
Let $\mathcal{D}$ be the associative unital superalgebra generated by
$x^{a}_{i},\,\eta^{a}_{j},\,y^{a}_{r},\,\zeta^{a}_{s}$
as well as their derivatives
$\partial_{x^{a}_{i}}:=\frac{\partial}{\partial x_{i}^{a}},\quad\partial_{\eta^{a}_{j}}:=\frac{\partial}{\partial\eta_{j}^{a}},\quad\partial_{y^{a}_{r}}:=\frac{\partial}{\partial y_{r}^{a}},\quad\partial_{\zeta^{a}_{s}}:=\frac{\partial}{\partial\zeta_{r}^{a}}$
for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$ and $a=1,\ldots,d$. Here $x^{a}_{i}$, $y^{a}_{r}$, $\partial_{x^{a}_{i}}$ and $\partial_{y^{a}_{r}}$ are even while $\eta^{a}_{j}$, $\zeta^{a}_{s}$, $\partial_{\eta^{a}_{j}}$, and $\partial_{\zeta^{a}_{s}}$ are odd. The superalgebra $\mathcal{D}$ is called the Weyl superalgebra.
There is an anti-involution $\omega:\mathcal{D}\longrightarrow\mathcal{D}$ defined by
${\begin{array}[]{lll}x^{a}_{i}\mapsto\partial_{x^{a}_{i}},&\quad\partial_{x^{a}_{i}}\mapsto x^{a}_{i},\quad\eta^{a}_{j}\mapsto\partial_{\eta^{a}_{j}},&\quad\partial_{\eta^{a}_{j}}\mapsto\eta^{a}_{j},\\ y^{a}_{r}\mapsto\partial_{y^{a}_{r}},&\quad\partial_{y^{a}_{r}}\mapsto y^{a}_{r},\quad\zeta^{a}_{s}\mapsto\partial_{\zeta^{a}_{s}},&\quad\partial_{\zeta^{a}_{s}}\mapsto\zeta^{a}_{s},\end{array}}$ (4.1)
for $1\leq i\leq m$, $1\leq j\leq n$, $1\leq r\leq p$, and $1\leq s\leq q$. In light of Proposition 3.6, the map $\omega$ extends to an anti-involution
$\omega:\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!).$
According to [CLZ, pp. 789–790], there is a superalgebra homomorphism $\phi:U({\mathfrak{gl}}_{d})\longrightarrow\mathcal{D}$, given by
$e_{ab}\mapsto\sum_{i=1}^{m}x_{i}^{a}\partial_{x^{b}_{i}}+\sum_{j=1}^{n}\eta_{j}^{a}\partial_{\eta^{b}_{j}}-\sum_{r=1}^{p}y_{r}^{b}\partial_{y^{a}_{r}}-\sum_{s=1}^{q}\zeta_{s}^{b}\partial_{\zeta^{a}_{s}},$ (4.2)
for $1\leq a,b\leq d$, and a superalgebra homomorphism $\overline{\phi}:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$, given by
${\begin{array}[]{cc}E^{r}_{r^{\prime}}\mapsto-\sum_{a=1}^{d}\partial_{y^{a}_{r}}y_{r^{\prime}}^{a},&E^{r}_{p+s}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\zeta_{s}^{a},\\ E^{p+s}_{r}\mapsto-\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}y_{r}^{a},&E^{p+s}_{p+s^{\prime}}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\zeta_{s^{\prime}}^{a},\\ E^{r}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{x^{a}_{i}},&E^{r}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{\eta^{a}_{j}},\\ E^{p+s}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{x^{a}_{i}},&E^{p+s}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{\eta^{a}_{j}},\\ E^{p+q+i}_{r}\mapsto-\sum_{a=1}^{d}x^{a}_{i}y_{r}^{a},&E^{p+q+i}_{p+s}\mapsto\sum_{a=1}^{d}x^{a}_{i}\zeta_{s}^{a},\\ E^{p+q+m+j}_{r}\mapsto-\sum_{a=1}^{d}\eta^{a}_{j}y_{r}^{a},&E^{p+q+m+j}_{p+s}\mapsto\sum_{a=1}^{d}\eta^{a}_{j}\zeta_{s}^{a},\\ E^{p+q+i}_{p+q+i^{\prime}}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{x^{a}_{i^{\prime}}},&E^{p+q+i}_{p+q+m+j}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{\eta^{a}_{j}},\\ E^{p+q+m+j}_{p+q+i}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{x^{a}_{i}},&E^{p+q+m+j}_{p+q+m+j^{\prime}}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{\eta^{a}_{j^{\prime}}}.\end{array}}$ (4.3)
where $1\leq i,i^{\prime}\leq m$, $1\leq j,j^{\prime}\leq n$, $1\leq r,r^{\prime}\leq p$, and $1\leq s,s^{\prime}\leq q$.
Let $\mathcal{F}$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$ and $\zeta^{a}_{s}$, for $1\leq i\leq m$, $1\leq j\leq n$, $1\leq r\leq p$, and $1\leq s\leq q$. With an appropriate inner product, $\mathcal{F}$ may be regarded as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators. The Weyl superalgebra $\mathcal{D}$ acts naturally on $\mathcal{F}$, and so it is a subalgebra of ${\rm End}(\mathcal{F})$.
We simplify notations for the rest of this section. Let
$x^{a}_{m+j}=\eta^{a}_{j},\quad\partial_{x^{a}_{m+j}}=\partial_{\eta^{a}_{j}},\quad y^{a}_{p+s}=\xi^{a}_{s},\quad\partial_{y^{a}_{p+s}}=\partial_{\zeta^{a}_{s}},$
where $1\leq j\leq n$ and $1\leq s\leq q$. Fix two sequences
${\bf w}:=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}\qquad\text{and}\qquad{\bf z}:=(z_{1},\ldots,z_{p+q+m+n})\in{\mathbb{C}}^{p+q+m+n}.$
The map $\phi$ induces a superalgebra homomorphism $\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])\longrightarrow\mathcal{D}$ defined by
$e_{ab}[k]\mapsto\sum_{i=1}^{m+n}z_{p+q+i}^{k}x^{a}_{i}\partial_{x^{b}_{i}}-\sum_{r=1}^{p+q}z_{r}^{k}y^{b}_{r}\partial_{y^{a}_{r}},$ (4.4)
for $1\leq a,b\leq d$ and $k\in{\mathbb{Z}}_{+}$, and $\overline{\phi}$ induces a superalgebra homomorphism $\overline{\phi}_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])\longrightarrow\mathcal{D}$ defined by
${\begin{array}[]{cll}E^{r}_{s}[k]&\mapsto\sum_{a=1}^{d}(-1)^{|s|+1}w_{a}^{k}\partial_{y^{a}_{r}}y^{a}_{s},\quad E^{r}_{p+q+j}[k]&\mapsto\sum_{a=1}^{d}w_{a}^{k}\partial_{y^{a}_{r}}\partial_{x^{a}_{j}},\\ E^{p+q+i}_{s}[k]&\mapsto\sum_{a=1}^{d}(-1)^{|s|+1}w_{a}^{k}x^{a}_{i}y^{a}_{s},\quad E^{p+q+i}_{p+q+j}[k]&\mapsto\sum_{a=1}^{d}w_{a}^{k}x^{a}_{i}\partial_{x^{a}_{j}},\end{array}}$ (4.5)
for $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$.
The Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ are respectively commutative subalgebras of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$. Thus, the maps $\phi_{\bf z}$ and $\overline{\phi}_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ and an action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$, respectively, on the polynomial superalgebra $\mathcal{F}$.
The Cartan anti-involution $\boldsymbol{\theta}$ on ${\mathfrak{gl}}_{d}$, defined in (3.7), gives rise to an anti-involution $\boldsymbol{\theta}:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$, defined by
$\hskip 42.67912pt\boldsymbol{\theta}(A[r])=\boldsymbol{\theta}(A)[r]\qquad\mbox{for $A\in{\mathfrak{gl}}_{d}$ and $r\in{\mathbb{Z}}_{+}$}.$
Proposition 4.1 ([MTV1, Proposition 8.4]).
The anti-involution $\boldsymbol{\theta}:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$ fixes each element of $\mathcal{B}_{d}^{\bf w}$.
Recall the anti-involution $\omega$ on $\mathcal{D}$ defined in (4.1). It is straightforward to verify that the diagram
[…]
commutes.
Proposition 4.2.
For any ${\boldsymbol{b}}\in\mathcal{B}_{d}^{\bf w}$, $\omega\big{(}\phi_{\bf z}({\boldsymbol{b}})\big{)}=\phi_{\bf z}({\boldsymbol{b}})$.
Proof.
This follows from (4.6) and Proposition 4.1. ∎
4.2. From Berezinians to column determinants
The map $\phi_{\bf z}$ and $\overline{\phi}_{\bf w}$, defined in (4.4) and (4.5), extend naturally to the superalgebra homomorphisms
$\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$
and
$\overline{\phi}_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!),$
respectively.
Let
$\mathcal{C}=\left[\delta_{a,b}(z-w_{a})-\left(\sum_{i=1}^{m+n}\frac{x^{b}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}-\sum_{r=1}^{p+q}\frac{y^{a}_{r}\partial_{y^{b}_{r}}}{\partial_{z}-z_{r}}\right)\right]_{a,b=1,\ldots,d}.$
Lemma 4.3.
We have
$\omega\big{(}\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\big{)}={\rm cdet}\,(\mathcal{C}).$
Proof.
For $1\leq a,b\leq d$,
$\phi_{\bf z}\big{(}e_{ab}(z)\big{)}=\sum_{i=1}^{m+n}\frac{x^{a}_{i}\partial_{x^{b}_{i}}}{z-z_{p+q+i}}-\sum_{r=1}^{p+q}\frac{y^{b}_{r}\partial_{y^{a}_{r}}}{z-z_{r}}.$ (4.7)
We define
$F^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(\partial_{z}-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{b}_{i}\partial_{y^{a}_{i}}}{z-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{a}_{i-p-q}\partial_{x^{b}_{i-p-q}}}{z-z_{i}}}&\mbox{if }\,p+q+1\leq i\leq p+q+m+n,\end{cases}$
and
$G^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(z-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{a}_{i}\partial_{y^{b}_{i}}}{\partial_{z}-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{b}_{i-p-q}\partial_{x^{a}_{i-p-q}}}{\partial_{z}-z_{i}}}&\mbox{if }\,p+q+1\leq i\leq p+q+m+n.\end{cases}$
We have
$\omega(F^{i}_{a,b})=G^{i}_{a,b}$ (4.8)
for all $0\leq i\leq{p+q+m+n}$ and $1\leq a,b\leq d$.
Note that ${\mathcal{L}}_{d}^{\bf w}$ is an amply invertible Manin matrix of type $(0^{d})$. Let $\pi\in\mathfrak{S}_{d}$ be the permutation given by $\pi(i)=d+1-i$ for $i=1,\ldots,d$. Clearly, $\pi^{2}=1$. By (4.7), Proposition 3.5 and Proposition 3.12, we obtain
$\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})$ $=$ $\phi_{\bf z}{\rm cdet}\,\big{(}({\mathcal{L}}_{d}^{\bf w})^{\pi}\big{)}$
$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\pi(\sigma(1)),\pi(1)}\ldots F^{i_{1}}_{\pi(\sigma(d)),\pi(d)}$
$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\pi\sigma(1),d}\ldots F^{i_{1}}_{\pi\sigma(d),1}$
$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\sigma(d),d}\ldots F^{i_{1}}_{\sigma(1),1}.$
Here the last equality holds since $\pi^{-1}\mathfrak{S}_{d}\pi=\mathfrak{S}_{d}$ and $l(\pi^{-1}\sigma\pi)=l(\sigma)$ for all $\sigma\in\mathfrak{S}_{d}$. Thus,
$\omega\big{(}\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\big{)}=\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}\omega\!\left(F^{i_{1}}_{\sigma(1),1}\right)\ldots\omega\!\left(F^{i_{d}}_{\sigma(d),d}\right).$
In view of (4.8), we deduce that $\omega\big{(}{\rm cdet}\,\phi_{\bf z}\,{\mathcal{L}}_{d}^{\bf w}\big{)}={\rm cdet}\,(\mathcal{C})$. ∎
Fix ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. For $u=z$ or $\partial_{z}$, let
$R_{p|q}(u)=\frac{(u-z_{1})\ldots(u-z_{p})}{(u-z_{p+1})\ldots(u-z_{p+q})}$
and
$R^{\prime}_{m|n}(u)=\frac{(u-z_{p+q+1})\ldots(u-z_{p+q+m})}{(u-z_{p+q+m+1})\ldots(u-z_{p+q+m+n})}.$
Proposition 4.4.
$(z-w_{1})\ldots(z-w_{d})\,\overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)=\omega\left(\!R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\cdot R_{p|q}(z)\!\right).$
Proof.
For $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$, we have
${\begin{array}[]{cll}\overline{\phi}_{\bf w}\big{(}E^{r}_{s}(z)\big{)}&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}\partial_{y^{a}_{r}}y^{a}_{s}}{z-w_{a}},\quad\overline{\phi}_{\bf w}\big{(}E^{r}_{p+q+j}(z)\big{)}&=\sum_{a=1}^{d}\frac{\partial_{y^{a}_{r}}\partial_{x^{a}_{j}}}{z-w_{a}},\\ \overline{\phi}_{\bf w}\big{(}E^{p+q+i}_{s}(z)\big{)}&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}x^{a}_{i}y^{a}_{s}}{z-w_{a}},\quad\overline{\phi}_{\bf w}\big{(}E^{p+q+i}_{p+q+j}(z)\big{)}&=\sum_{a=1}^{d}\frac{x^{a}_{i}\partial_{x^{a}_{j}}}{z-w_{a}}.\end{array}}$ (4.9)
Let
$D=\Big{[}\delta_{a,b}(z-w_{a})\Big{]}_{a,b=1,\ldots,d}\quad\text{and}\quad D^{\prime}=\Big{[}\delta_{i,j}(\partial_{z}-z_{i})\Big{]}_{i,j\in\mathbb{I}}.$
Consider the $(m+n)\times d$ matrices
$X:=\big{[}(-1)^{|i|}x^{a}_{i}\big{]}^{a=1,\ldots,d}_{i=1,\ldots,m+n}\quad\text{and}\quad P_{X}:=\big{[}\partial_{x^{a}_{i}}\big{]}^{a=1,\ldots,d}_{i=1,\ldots,m+n}$
and the $(p+q)\times d$ matrices
$Y:=\big{[}(-1)^{|r|+1}y^{a}_{r}\big{]}^{a=1,\ldots,d}_{r=1,\ldots,p+q}\qquad\text{and}\qquad P_{Y}:=\big{[}(-1)^{|r|}\partial_{y^{a}_{r}}\big{]}^{a=1,\ldots,d}_{r=1,\ldots,p+q}.$
Write
$T=\begin{bmatrix}Y^{t}\hskip 5.69054ptP_{X}^{t}\end{bmatrix}\quad\text{and}\quad T^{\prime}=\begin{bmatrix}P_{Y}\vspace{3mm}\\ X\end{bmatrix}.$
By (4.9), we get
$\overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)={\rm Ber}^{\bf s}\,(D^{\prime}-T^{\prime}D^{-1}T).$
Let
$\mathfrak{L}=\begin{bmatrix}D&T\vspace{2mm}\\ T^{\prime}&D^{\prime}\end{bmatrix}.$
By Proposition 3.1, we have
${\rm Ber}\,^{\!\hat{{\bf s}}}\,(\mathfrak{L})=(z-w_{1})\ldots(z-w_{d})\,\overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right),$
where $\hat{{\bf s}}:=(0^{d},{\bf s})\in\mathcal{S}_{d+p+m|q+n}$. As $D-T(D^{\prime})^{-1}T^{\prime}$ is a sufficiently invertible Manin matrix of type $(0^{d})$, Proposition 3.4 and Proposition 3.5 imply that
${\rm Ber}\,^{\!\hat{{\bf s}}}\,(\mathfrak{L})$ $=$ ${\rm Ber}^{\bf s}\,(D^{\prime})\cdot{\rm cdet}\,(D-T(D^{\prime})^{-1}T^{\prime})$
$=$ $R_{p|q}(\partial_{z})\cdot R^{\prime}_{m|n}(\partial_{z})\cdot{\rm cdet}\,(\mathcal{C}^{\prime}),$
where
$\mathcal{C}^{\prime}:=\left[\delta_{a,b}(z-w_{a})-\left(\sum_{i=1}^{m+n}\frac{(-1)^{|i|}\partial_{x^{a}_{i}}x^{b}_{i}}{\partial_{z}-z_{p+q+i}}-\sum_{r=1}^{p+q}\frac{y^{a}_{r}\partial_{y^{b}_{r}}}{\partial_{z}-z_{r}}\right)\right]_{a,b=1,\ldots,d}.$
We claim that
$R^{\prime}_{m|n}(\partial_{z})\cdot{\rm cdet}\,(\mathcal{C}^{\prime})={\rm cdet}\,(\mathcal{C})\cdot R^{\prime}_{m|n}(\partial_{z}).$
This yields
${\rm Ber}\,^{\!\hat{{\bf s}}}\,(\mathfrak{L})=R_{p|q}(\partial_{z})\cdot{\rm cdet}\,(\mathcal{C})\cdot R^{\prime}_{m|n}(\partial_{z}),$
which proves the proposition by Lemma 4.3.
To prove the claim, it suffices to show that
$R^{\prime}_{m|n}(\partial_{z})\cdot\mathcal{C}^{\prime}_{a,b}=\mathcal{C}_{a,b}\cdot R^{\prime}_{m|n}(\partial_{z}),\qquad\mbox{for $1\leq a,b\leq d$.}$ (4.10)
Here, for instance, $\mathcal{C}_{a,b}$ denotes the $(a,b)$ entry of $\mathcal{C}$. Since
$\partial_{x^{a}_{i}}x^{b}_{i}=(-1)^{|i|}x^{b}_{i}\partial_{x^{a}_{i}}+\delta_{a,b},\qquad\mbox{for $1\leq a,b\leq d$},$ (4.11)
the off-diagonal entries of $\mathcal{C}^{\prime}$ coincide with those of $\mathcal{C}$. For $1\leq i\leq m+n$, the term $\partial_{z}-z_{p+q+i}$ commutes with $\partial_{x^{a}_{j}}x^{b}_{j}$ and $y^{a}_{r}\partial_{y^{b}_{r}}$ for all $1\leq j\leq m+n$, $1\leq r\leq p+q$ and $1\leq a,b\leq d$. In particular, $R^{\prime}_{m|n}(\partial_{z})$ commutes with any off-diagonal entry of $\mathcal{C}^{\prime}$. This proves (4.10) for $a\not=b$. Moreover, for $1\leq i\leq m+n$ and $1\leq a\leq d$,
$(\partial_{z}-z_{p+q+i})(z-w_{a})=(z-w_{a})(\partial_{z}-z_{p+q+i})+1.$ (4.12)
If $1\leq i\leq m$, we see from (4.11) and (4.12) that
${(\partial_{z}-z_{p+q+i})}\left(z-w_{a}-\frac{\partial_{x^{a}_{i}}x^{a}_{i}}{\partial_{z}-z_{p+q+i}}\right)=\left(z-w_{a}-\frac{x^{a}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}\right)(\partial_{z}-z_{p+q+i}).$
Similarly, if $m+1\leq i\leq m+n$, we have
${(\partial_{z}-z_{p+q+i})^{-1}}\left(z-w_{a}+\frac{\partial_{x^{a}_{i}}x^{a}_{i}}{\partial_{z}-z_{p+q+i}}\right)=\left(z-w_{a}-\frac{x^{a}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}\right)(\partial_{z}-z_{p+q+i})^{-1}.$
The above discussion proves (4.10) for $a=b$. Thus, the proof of (4.10) is now complete. ∎
We relate ${\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)$ to ${\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})$ as follows.
Proposition 4.5.
We may write
$R_{p|q}(z)\cdot R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})=\sum_{\ell=0}^{d}\sum_{k=-\infty}^{p+m-q-n}b_{k,\ell}^{{\bf w},{\bf z}}z^{k}\partial_{z}^{\ell},$
where $b_{k,\ell}^{{\bf w},{\bf z}}:=\phi_{\bf z}(b_{k,\ell}^{\bf w})$ for some $b_{k,\ell}^{\bf w}\in\mathcal{B}_{d}^{\bf w}$. Then
$(z-w_{1})\ldots(z-w_{d})\ \overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)$
$\hskip 85.35826pt=R_{p|q}(\partial_{z})\left(\sum_{k=-\infty}^{p+m-q-n}\sum_{\ell=0}^{d}b_{k,\ell}^{{\bf w},{\bf z}}z^{\ell}\partial_{z}^{k}\right)R_{p|q}(\partial_{z})^{-1}.$
Proof.
By Proposition 4.4,
$(z-w_{1})\ldots(z-w_{d})\,\overline{\phi}_{\bf w}{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)$
$\hskip 42.67912pt=R_{p|q}(\partial_{z})\cdot\omega\left(R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\right)\!$
$\hskip 42.67912pt=R_{p|q}(\partial_{z})\cdot\omega\left(R_{p|q}(z)\cdot R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\right)\cdot R_{p|q}(\partial_{z})^{-1}.$
Therefore,
$(z-w_{1})\ldots(z-w_{d})\ \overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)$
$\hskip 85.35826pt=R_{p|q}(\partial_{z})\left(\sum_{k=-\infty}^{p+m-q-n}\sum_{\ell=0}^{d}\omega\big{(}b_{k,\ell}^{{\bf w},{\bf z}}\big{)}z^{\ell}\partial_{z}^{k}\right)R_{p|q}(\partial_{z})^{-1}.$
By Proposition 4.2,
$\omega\big{(}b_{k,\ell}^{{\bf w},{\bf z}}\big{)}=b_{k,\ell}^{{\bf w},{\bf z}}$
for all $k\leq p+m-q-n$ and $0\leq\ell\leq d$. This completes the proof. ∎
4.3. Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$
Lemma 4.6.
Let $c\in\{z_{1},\ldots,z_{p+q}\}$. Given $\alpha_{r,s}^{{\bf w},{\bf z}}\in\mathcal{D}$ for $r,s\in{\mathbb{Z}}$ with $0\leq r\leq d$ and $s\leq p+m-q-n$. Let $\mathcal{D}_{1}$ be the subalgebra of $\mathcal{D}$ generated by all $\alpha_{r,s}^{{\bf w},{\bf z}}$’s, and let $\mathcal{D}_{2}$ be the subalgebra of $\mathcal{D}$ generated by the coefficients of the series
$(\partial_{z}-c)\left(\sum_{s=-\infty}^{p+m-q-n}\sum_{r=0}^{d}\alpha_{r,s}^{{\bf w},{\bf z}}z^{r}\partial_{z}^{s}\right)(\partial_{z}-c)^{-1}.$ (4.13)
Then $\mathcal{D}_{1}=\mathcal{D}_{2}$.
Proof.
For any $r\in{\mathbb{Z}}_{+}$,
$(\partial_{z}-c)z^{r}(\partial_{z}-c)^{-1}=z^{r}+rz^{r-1}(\partial_{z}-c)^{-1}.$
Let $f$ denote the series (4.13). We find that
$f=\sum_{s=-\infty}^{p+m-q-n}\sum_{r=0}^{d}\left(\alpha_{r,s}^{{\bf w},{\bf z}}z^{r}\partial_{z}^{s}+r\alpha_{r,s}^{{\bf w},{\bf z}}z^{r-1}\partial_{z}^{s}(\partial_{z}-c)^{-1}\right)\!,$
and hence
$f=\sum_{\ell=-\infty}^{p+m-q-n}\sum_{k=0}^{d}\beta_{k,\ell}^{{\bf w},{\bf z}}z^{k}\partial_{z}^{\ell},$
where
$\beta_{d,\ell}^{{\bf w},{\bf z}}=\alpha_{d,\ell}^{{\bf w},{\bf z}},$ (4.14)
$\beta_{k,\ell}^{{\bf w},{\bf z}}=\alpha_{k,\ell}^{{\bf w},{\bf z}}+\sum_{i=0}^{\infty}(k+1)\alpha_{k+1,\ell+i+1}^{{\bf w},{\bf z}}c^{i},$ (4.15)
for $10\leq k\leq d-$ 1 and $\ell\leq p+m-q-n$. Note that the right-hand side of (4.15) has only finitely many nonzero terms.
The algebra $\mathcal{D}_{2}$ is, by definition, generated by $\beta_{k,\ell}^{{\bf w},{\bf z}}$’s. So, it is a subalgebra of $\mathcal{D}_{1}$ by (4.14) and (4.15). Also, (4.14) means that $\alpha_{d,\ell}^{{\bf w},{\bf z}}\in\mathcal{D}_{2}$ for all $\ell\leq p+m-q-n$. Using (4.15), we see, by induction, that $\alpha_{r,\ell}^{{\bf w},{\bf z}}\in\mathcal{D}_{2}$ for all $0\leq r\leq d$ and $\ell\leq p+m-q-n$. It follows that $\mathcal{D}_{1}\subseteq\mathcal{D}_{2}$, and thus $\mathcal{D}_{1}=\mathcal{D}_{2}$. ∎
We have the following duality between the Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$. We call it the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$.
Theorem 4.7.
$\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}$.
Proof.
This follows from Proposition 4.5 and repeated use of Lemma 4.6. ∎
Remark 4.8.
We will demonstrate another duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for Gaudin models with irregular singularities, which specializes to a variant of Theorem 4.7, in [ChL3]. The classical version of the duality will also be investigated there.
If some of $p,q,m,n$ are 0, we will suppress (most of) them in the symbol ${\mathcal{B}_{p+m|q+n}^{\bf z}}$. For instance,
$\mathcal{B}^{\bf z}_{p+m|n}:=\mathcal{B}^{\bf z}_{p+m|0+n},\quad\mathcal{B}^{\bf z}_{m|n}:=\mathcal{B}^{\bf z}_{0+m|0+n},\quad\mathcal{B}^{\bf z}_{0|n}:=\mathcal{B}^{\bf z}_{0+0|0+n},\quad\text{etc.}$. (4.16)
Theorem 4.7 gives the Bethe dualities of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|n})$ and $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m})$.
Corollary 4.9.
  1. (i)
    Suppose $q=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{p+m|n}\big{)}$.
  2. (ii)
    Suppose $q=n=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{p+m}\big{)}$.
Theorem 4.7 recovers the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$ due to Huang and Mukhin as well as the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ due to Mukhin, Tarasov, and Varchenko.
Corollary 4.10 ([HM, Theorem 5.2]).
Suppose $p=q=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{m|n}\big{)}$.
Corollary 4.11 ([MTV3, Theorem 6]).
Suppose $p=q=n=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{m}\big{)}$.
We also obtain a variant of the duality of Bethe algebras established by Tarasov and Uvarov (cf. [TU, Theorem 5.2]).
Corollary 4.12.
Suppose $p=q=m=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{0|n}\big{)}$.
5. An application of Bethe duality
In this section, we recall *-structures and unitarizable modules. We apply the Bethe duality to the Bethe algebra of ${\mathfrak{gl}}_{p+m|q+n}$ acting on a tensor product of infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules.
5.1. *-structures and unitarizable modules
A *-superalgebra is an associative superalgebra $A$ with an even anti-linear anti-involution ${\boldsymbol{\sigma}}:A\longrightarrow A$. A homomorphism $f:(A,{\boldsymbol{\sigma}})\longrightarrow(A^{\prime},{\boldsymbol{\sigma}}^{\prime})$ of *-superalgebras is a homomorphism of superalgebras satisfying ${\boldsymbol{\sigma}}^{\prime}\circ f=f\circ{\boldsymbol{\sigma}}$.
[…]
$\langle av|v^{\prime}\rangle=\langle v|{\boldsymbol{\sigma}}(a)v^{\prime}\rangle,$
for all $a\in A$ and $v,v^{\prime}\in V$. An $A$-module equipped with a positive definite contravariant Hermitian form is called a unitarizable $A$-module (with respect to ${\boldsymbol{\sigma}}$).
Let ${\mathfrak{g}}$ be a Lie superalgebra. An anti-linear anti-involution ${\boldsymbol{\sigma}}:{\mathfrak{g}}\longrightarrow{\mathfrak{g}}$ is called a *-structure on ${\mathfrak{g}}$. A homomorphism $f:({\mathfrak{g}},{\boldsymbol{\sigma}})\longrightarrow({\mathfrak{g}}^{\prime},{\boldsymbol{\sigma}}^{\prime})$ of Lie superalgebras with *-structures is a homomorphism of Lie superalgebras satisfying ${\boldsymbol{\sigma}}^{\prime}\circ f=f\circ{\boldsymbol{\sigma}}$. Moreover, ${\boldsymbol{\sigma}}$ is a *-structure on ${\mathfrak{g}}$ if and only if the natural extension of ${\boldsymbol{\sigma}}$ to $U({\mathfrak{g}})$ is an anti-linear anti-involution.
[…]
The Lie superalgebra ${\mathfrak{gl}}_{p+m|q+n}$ admits a *-structure ${\boldsymbol{\sigma}}:{\mathfrak{gl}}_{p+m|q+n}\longrightarrow{\mathfrak{gl}}_{p+m|q+n}$ (see [CLZ, Section 3.2]), defined by
$\hskip 42.67912pt\sum_{i,j\in\mathbb{I}}a_{ij}E^{i}_{j}\mapsto\sum_{i,j\in\mathbb{I}}(-1)^{[i]+[j]}\bar{a}_{ij}E^{j}_{i}\qquad\mbox{for $1\leq i,j\leq p+q+m+n$},$
where
$[i]:=\begin{cases}\,0&\,\,\mbox{if}\ \ 1\leq i\leq p,\\ \,1&\,\,\mbox{otherwise},\end{cases}$
and $\bar{a}_{ij}$ denotes the complex conjugate of $a_{ij}\in{\mathbb{C}}$. Let $\pi:{\mathfrak{gl}}_{d}\longrightarrow{\mathfrak{gl}}_{d}$ be the *-structure on ${\mathfrak{gl}}_{d}$ defined by
$\pi(e_{ab})=e_{ba}\quad\mbox{for $1\leq a,b\leq d$.}$
The maps $\pi$ and ${\boldsymbol{\sigma}}$ give rise to a *-structure $(\pi,{\boldsymbol{\sigma}})$ on ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$.
5.2. Bethe algebras of unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules
The Weyl superalgebra $\mathcal{D}$ is a subalgebra of the endomorphism algebra ${\rm End}(\mathcal{F})$ of the polynomial superalgebra $\mathcal{F}$. For ${\bf w}\in{\mathbb{C}}^{d}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$, Theorem 4.7 can be rephrased as saying that the image of $\mathcal{B}_{d}^{\bf w}$ in ${\rm End}(\mathcal{F})$ under $\phi_{\bf z}$ coincides with the image of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ in ${\rm End}(\mathcal{F})$ under $\overline{\phi}_{\bf w}$. This makes it possible to study the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on $\mathcal{F}$ via the action of $\mathcal{B}_{d}^{\bf w}$ on $\mathcal{F}$, and vice versa.
[…]
for $1\leq a\leq d$. As ${\mathfrak{gl}}_{d}$-modules,
$\mathcal{F}\cong\left(\!\bigotimes_{r=1}^{p}{\mathbb{C}}[{\bf y}_{r}]\!\right)\otimes\left(\!\bigotimes_{s=1}^{q}{\mathbb{C}}[{\boldsymbol{\zeta}}_{s}]\!\right)\otimes\left(\!\bigotimes_{i=1}^{m}{\mathbb{C}}[{\bf x}_{i}]\!\right)\otimes\left(\!\bigotimes_{j=1}^{n}{\mathbb{C}}[{\boldsymbol{\eta}}_{j}]\!\right)\!,$ (5.1)
and as ${\mathfrak{gl}}_{p+m|q+n}$-modules,
$\mathcal{F}\cong{\mathbb{C}}[{\boldsymbol{\Sigma}}^{1}]\otimes\cdots\otimes{\mathbb{C}}[{\boldsymbol{\Sigma}}^{d}].$ (5.2)
The Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ asserts that $\mathcal{F}$ decomposes into a direct sum of unitarizable ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-modules (with respect to $(\pi,{\boldsymbol{\sigma}})$). We refer the reader to [CLZ, Theorem 3.3] for further details.
For $\xi\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$, we denote by $L_{p+m|q+n}(\xi)$ the irreducible highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module with highest weight $\xi$ with respect to the Borel subalgebra $\mathfrak{b}_{{p+m|q+n}}$. Let
$\mathbbm{1}_{p|q}=\sum_{i=1}^{p}\epsilon_{i}-\sum_{j=1}^{q}\epsilon_{p+j}\qquad\text{and}\qquad\mathbbm{1}_{p}=\mathbbm{1}_{p|0}.$
Suppose $p\not=0\not=m$. For $r\in{\mathbb{Z}}$, let
$V_{r}=\begin{cases}\displaystyle{L_{{p+m|q+n}}\big{(}r\epsilon_{p+q+1}-\mathbbm{1}_{p|q}\big{)}}&\quad\mbox{if $r\geq 0$,}\\ \displaystyle{L_{{p+m|q+n}}\left(\!-\sum_{i=q+r+1}^{q}\epsilon_{p+i}-\mathbbm{1}_{p|q}\!\right)}&\quad\mbox{if $-q\leq r<0$ and $q\not=0$,}\\ \displaystyle{L_{{p+m|q+n}}\left(\!(r+q)\epsilon_{p}-\sum_{i=1}^{q}\epsilon_{p+i}-\mathbbm{1}_{p|q}\!\right)}&\quad\mbox{if $r<-q$ and $q\not=0$,}\\ \displaystyle{L_{{p+m|q+n}}\big{(}r\epsilon_{p}-\mathbbm{1}_{p}\big{)}}&\quad\mbox{if $r<0$ and $q=0$}.\end{cases}$
Note that $V_{r}$ is an infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module (with respect to ${\boldsymbol{\sigma}}$), and that the highest weight of $V_{r}$ corresponds to a generalized partition of depth 1; see [CCL2, Section 2.5] and also [CLZ]. (Such a partition is called a generalized partition of length 1 in [CLZ].)
For any weight $\mu$ of a ${\mathfrak{gl}}_{p+m|q+n}$-module $M$, we denote by $M_{\mu}$ the $\mu$-weight space of $M$. For ${\mathfrak{gl}}_{p+m|q+n}$-modules $M_{1},\ldots,M_{d}$ and ${\bf w}=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$, we write
$\underline{M\mkern-4.0mu}\mkern 4.0mu=M_{1}\otimes\cdots\otimes M_{d}\qquad\text{and}\qquad\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})=M_{1}(w_{1})\otimes\cdots\otimes M_{d}(w_{d}).$
Analogous to (4.16), we write
$L_{p+m}(\xi)=L_{p+m|0+0}(\xi)\qquad\text{and}\qquad L_{m}(\xi)=L_{0+m|0+0}(\xi).$
Theorem 5.1.
Suppose $p\not=0\not=m$. For $a=1,\ldots,d$, let $M_{a}=V_{\gamma_{a}}$ for some $\gamma_{a}\in{\mathbb{Z}}$. For each weight $\mu$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!p+q+m+n}$, we have:
  1. (i)
    $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}$ is a cyclic ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module.
  2. (ii)
[…]
  3. (iii)
[…]
  4. (iv)
[…]
Proof.
Fix $1\leq a\leq d$. If $\gamma_{a}\geq 0$, then we may identify $M_{a}$ with the ${\mathfrak{gl}}_{p+m|q+n}$-submodule of the polynomial superalgebra ${\mathbb{C}}[{\boldsymbol{\Sigma}}^{a}]$ generated by the highest weight vector $(x^{a}_{1})^{\gamma_{a}}$. If $\gamma_{a}<0$, then we may identify $M_{a}$ with the ${\mathfrak{gl}}_{p+m|q+n}$-submodule of ${\mathbb{C}}[{\boldsymbol{\Sigma}}^{a}]$ generated by the highest weight vector
$\begin{cases}\zeta^{a}_{q+\gamma_{a}+1}\ldots\zeta^{a}_{q}&\quad\text{if }\mbox{$\gamma_{a}\geq-q$ and $q\not=0$,}\\ (y^{a}_{p})^{-\gamma_{a}-q}\zeta^{a}_{1}\ldots\zeta^{a}_{q}&\quad\text{if }\mbox{$\gamma_{a}<-q$ and $q\not=0$,}\\ (y^{a}_{p})^{-\gamma_{a}}&\quad\text{if }q=0.\end{cases}$
Under these identifications, the tensor product $\underline{M\mkern-4.0mu}\mkern 4.0mu$ is a ${\mathfrak{gl}}_{p+m|q+n}$-submodule of ${\mathbb{C}}[{\boldsymbol{\Sigma}}^{1}]\otimes\cdots\otimes{\mathbb{C}}[{\boldsymbol{\Sigma}}^{d}]$ and hence a ${\mathfrak{gl}}_{p+m|q+n}$-submodule of $\mathcal{F}$ via (5.2). We will see in a moment that for each weight $\mu:=\sum_{i=1}^{p+q+m+n}\mu_{i}\epsilon_{i}$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, the $\mu$-weight space $\underline{M\mkern-4.0mu}\mkern 4.0mu_{\!\mu}$ is also a weight space of a tensor product of finite-dimensional irreducible ${\mathfrak{gl}}_{d}$-modules.
For $1\leq i\leq p$ (resp., $p+1\leq i\leq p+q$, $p+q+1\leq i\leq p+q+m$ and $p+q+m+1\leq i\leq p+q+m+n$), let $N_{i}$ denote the ${\mathfrak{gl}}_{d}$-submodule of ${\mathbb{C}}[y^{1}_{i},\cdots,y^{d}_{i}]$ (resp., ${\mathbb{C}}[\zeta^{1}_{i-p},\cdots,\zeta^{d}_{i-p}]$, ${\mathbb{C}}[x^{1}_{i-p-q},\cdots,x^{d}_{i-p-q}]$ and ${\mathbb{C}}[\eta^{1}_{i-p-q-m},\cdots,\eta^{d}_{i-p-q-m}]$) generated by the highest weight vector $(y^{d}_{i})^{-\mu_{i}-d}$ (resp., $\zeta^{\mu_{i}+1}_{i-p}\cdots\zeta^{d}_{i-p}$, $(x^{1}_{i-p-q})^{\mu_{i}}$ and $\eta^{1}_{i-p-q-m}\cdots\eta^{\mu_{i}}_{i-p-q-m}$). Then
$N_{i}\cong\begin{cases}L_{d}\big{(}(\mu_{i}+d)\epsilon_{d}\big{)}&\quad\text{if }1\leq i\leq p,\\ \displaystyle{L_{d}\left(\!-\sum_{a=\mu_{i}+1}^{d}\epsilon_{a}\!\right)}&\quad\text{if }p+1\leq i\leq p+q,\\ L_{d}(\mu_{i}\epsilon_{1})&\quad\text{if }p+q+1\leq i\leq p+q+m,\\ \displaystyle{L_{d}\left(\sum_{a=1}^{\mu_{i}}\epsilon_{a}\right)}&\quad\text{if }p+q+m+1\leq i\leq p+q+m+n,\end{cases}$
and so $N_{i}$ is an irreducible highest weight ${\mathfrak{gl}}_{d}$-module with a dominant integral highest weight. Write ${\bf z}=(z_{1},\ldots,z_{p+q+m+n})$. Let
$\underline{N\mkern-4.0mu}\mkern 4.0mu=N_{1}\otimes\cdots\otimes N_{p+q+m+n}\quad\text{and}\quad\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})=N_{1}(z_{1})\otimes\cdots\otimes N_{p+q+m+n}(z_{p+q+m+n}).$
The tensor product $\underline{N\mkern-4.0mu}\mkern 4.0mu$ is a ${\mathfrak{gl}}_{d}$-submodule of $\mathcal{F}$ via (5.1). By [CLZ, Theorem 3.3], we have
$\underline{M\mkern-4.0mu}\mkern 4.0mu_{\mu}\subset\underline{N\mkern-4.0mu}\mkern 4.0mu\qquad\hbox{and}\qquad\underline{N\mkern-4.0mu}\mkern 4.0mu_{\gamma}\subset\underline{M\mkern-4.0mu}\mkern 4.0mu,$
and no nonzero element in the $\nu$-weight space $\underline{N\mkern-4.0mu}\mkern 4.0mu_{\nu}$ is contained in $\underline{M\mkern-4.0mu}\mkern 4.0mu$ for all $\nu\not=\gamma$, where $\gamma:=\sum_{a=1}^{d}\gamma_{a}\epsilon_{a}\in{\mathfrak{h}}_{d}^{*}$. We deduce that $\underline{M\mkern-4.0mu}\mkern 4.0mu_{\mu}\subseteq\underline{N\mkern-4.0mu}\mkern 4.0mu_{\gamma}$. Similarly, $\underline{N\mkern-4.0mu}\mkern 4.0mu_{\gamma}\subseteq\underline{M\mkern-4.0mu}\mkern 4.0mu_{\mu}$. Thus we have
$\underline{M\mkern-4.0mu}\mkern 4.0mu_{\mu}=\underline{N\mkern-4.0mu}\mkern 4.0mu_{\gamma}.$
This gives
[…]
in view of Theorem 4.7. By Theorem 2.1, $\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}$ is a cyclic $\mathcal{B}_{d}^{\bf w}$-module as $\mathcal{B}_{d}^{\bf w}$ preserves weight spaces, and by Theorem 3.8, $(\mathcal{B}_{d}^{\bf w})_{\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}}$ is a Frobenius algebra. Hence, (i) and (ii) follow from (5.3). Also, (iii) and (iv) are immediate consequences of [Lu, Lemma 1.3] (cf. Remark 3.10). The last statement follows from (5.3) and Theorem 2.2. ∎
Remark 5.2.
A similar statement corresponding to $p=0$ or $m=0$ can be formulated. The details will be left to the reader.
Suppose $p\not=0\not=m$. For $r\in{\mathbb{Z}}$, let
$V_{r}=\begin{cases}L_{p+m}(r\epsilon_{p+1}-\mathbbm{1}_{p})&\quad\mbox{if $r\geq 0$,}\\ L_{p+m}(r\epsilon_{p}-\mathbbm{1}_{p})&\quad\mbox{if $r<0$}.\end{cases}$
We have the following result.
Corollary 5.3.
For $a=1,\ldots,d$, let $M_{a}=V_{\gamma_{a}}$ for some $\gamma_{a}\in{\mathbb{Z}}$. For each weight $\mu$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!p+m}$, we have:
  1. (i)
    $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}$ is a cyclic $\mathcal{B}_{p+m}^{\bf z}$-module.
  2. (ii)
    $(\mathcal{B}_{p+m}^{\bf z})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is a Frobenius algebra.
  3. (iii)
    $(\mathcal{B}_{p+m}^{\bf z})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is a maximal commutative subalgebra of ${\rm End}(\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu})$ of dimension $\dim\!\left(\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}\right)$.
  4. (iv)
    Every eigenspace of $(\mathcal{B}_{p+m}^{\bf z})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is one-dimensional.
Moreover, $(\mathcal{B}_{p+m}^{\bf z})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is diagonalizable with a simple spectrum for generic ${\bf w}$ and generic ${\bf z}$.
Proof.
This is an immediate consequence of Theorem 5.1 by taking $q=n=0$. ∎
The structures of the Bethe algebras of finite-dimensional modules over a general linear Lie algebra have been well studied (see [FFRy, MTV2]). However, very little is known about the Bethe algebras of infinite-dimensional modules. Corollary 5.3, which concerns the action of the Bethe algebra of ${\mathfrak{gl}}_{p+m}$ on a tensor product of infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m}$-modules, makes good progress toward the infinite-dimensional cases.
We expect that Theorem 2.1 and Theorem 2.2 should have unitarizable analogs. More generally, we have the following conjecture.
Conjecture 5.4.
Let $d\in{\mathbb{N}}$ and $p,q,m,n\in{\mathbb{Z}}_{+}$ that are not all zero. For $a=1,\ldots,d$, let $M_{a}$ be a unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module. For each weight $\mu$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!p+q+m+n}$, parts (i)–(iv) of Theorem 5.1 hold for the Bethe algebra $({\mathcal{B}_{p+m|q+n}^{\bf z}})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$. Moreover, $({\mathcal{B}_{p+m|q+n}^{\bf z}})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is diagonalizable with a simple spectrum for generic ${\bf w}$ and generic ${\bf z}$.
Remark 5.5.
An analog of the conjecture when the entries of ${\bf z}$ are all zero will be studied in [CCL2].
Acknowledgments. The first author was partially supported by NSTC grant 113-2115-M-006-010 of Taiwan. The second author was partially supported by NSTC grant 112-2115-M-006-015-MY2 of Taiwan.

### v2

@@
-3cm3cm3cm3cm
+Bethe algebras for unitarizable modules over classical Lie (super)algebras and a duality
 Abstract.
-We study the joint action of the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ for the general linear Lie algebra ${\mathfrak{gl}}_{d}$ with respect to ${\bf w}\in{\mathbb{C}}^{d}$ and the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ for the general linear Lie superalgebra ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$ on the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators.
-We establish a duality, called the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$, which is an equivalence between the actions of the algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on the Fock space.
-As an application, we show that the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on each weight space of the evaluation module $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})$, where $\underline{M\mkern-4.0mu}\mkern 4.0mu$ is a $d$-fold tensor product of certain classes of infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules, is a cyclic ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module, and that ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ is diagonalizable with a simple spectrum on the weight space for generic ${\bf w}$ and ${\bf z}$.
+Let $\mathfrak{g}$} denote the classical Lie algebra ${\mathfrak{gl}}_{d}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$ with a fixed *-structure $\bm{\sigma}$}.
+Let $M_{1},\ldots,M_{\ell}$ be unitarizable $\mathfrak{g}$}-modules (with respect to $\bm{\sigma}$}), and let ${\bf z}=(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}$.
+We investigate the action of the Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ for $\mathfrak{g}$} with respect to $\mu\in{\mathfrak{g}}^{*}$ on the tensor product $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z}):=M_{1}(z_{1})\otimes\cdots\otimes M_{\ell}(z_{\ell})$ of evaluation ${\mathfrak{g}}[t]$-modules.
+We show that if $\mu\circ{\bm{\sigma}}$ equals the complex conjugation of $\mu$, then $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable on any finite-dimensional $\mathcal{B}_{\mathfrak{g}}^{\mu}$-submodule of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for $\bf z$$\in{\mathbb{R}}^{\ell}$.
+This, together with the result derived from the duality of Bethe algebras (see below), suggests that a simple spectrum conjecture for $\mathcal{B}_{\mathfrak{g}}^{\mu}$ should hold.
+We establish a duality of Bethe algebras for the general linear Lie (super)algebras ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$.
+As an application, we show that under a generic condition, the Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$ is diagonalizable with a simple spectrum on any weight space of $L_{1}(w_{1})\otimes\cdots\otimes L_{d}(w_{d})$, where the $L_{i}$ are (infinite-dimensional) unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules corresponding to generalized partitions of depth 1, and $w_{1},\ldots,w_{d}\in{\mathbb{C}}$.
+We also obtain the corresponding result for ${\mathfrak{gl}}_{p+m}$ by setting $q=n=0$.
 1. Introduction
-Let $\mathcal{F}:={\mathbb{C}}[{\bf x},{\bf y},{\boldsymbol{\eta}},{\boldsymbol{\zeta}}]$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$ and $\zeta^{a}_{s}$, for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$ and $a=1,\ldots,d$.
+Let $\mathfrak{g}$} be a finite-dimensional semisimple Lie algebra over $\mathbb{C}$}.
+The Bethe algebra for $\mathfrak{g}$} is a commutative subalgebra of $U({\mathfrak{g}}[t])$ induced by the Feigin–Frenkel center (see [FFR, FFTL, Ry1]), where $U({\mathfrak{g}}[t])$ denotes the universal enveloping algebra of the current algebra ${\mathfrak{g}}[t]$.
+It produces (higher) Gaudin Hamiltonians, including the quadratic ones originally introduced by Gaudin [G1, G2] to describe a completely integrable quantum spin chain model.
+Also, the notion of the Bethe algebra can be extended to the general linear Lie (super)algebras (see [HM, MR, MTV1, Ta]).
+Let $d$ be a positive integer, and let $p$, $q$, $m$, and $n$ be non-negative integers that are not all zero.
+The aim of the paper is two-fold.
+Firstly, we investigate the action of the Bethe algebra for $\mathfrak{g}$} on unitarizable $\mathfrak{g}$}-modules, where $\mathfrak{g}$} is the classical Lie algebra of type $\mathfrak{a,c,d}$ of rank $d$ (i.e., ${\mathfrak{g}}={\mathfrak{gl}}_{d}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$).
+Secondly, we establish a duality of Bethe algebras for the general linear Lie (super)algebras ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$ and use it to study the action of the Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$ on some infinite-dimensional unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules.
+1.1. Unitarizable modules and diagonalization of Bethe algebras
+Let $\mathfrak{g}$} be a simple Lie algebra or the general linear Lie algebra over $\mathbb{C}$}.
+For $\mu\in{\mathfrak{g}}^{*}$, let $\mathcal{B}_{\mathfrak{g}}^{\mu}$ be the Bethe algebra for $\mathfrak{g}$} with respect to $\mu$; see Section 3.2.
+The action of $\mathcal{B}_{\mathfrak{g}}^{\mu}$ on a tensor product of finite-dimensional $\mathfrak{g}$}-modules has been extensively studied (see [FFRy, MTV1, MTV2, MTV4, MTV5, Ry3]).
+Let ${\bf z}=(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}$, and let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible $\mathfrak{g}$}-modules.
+Feigin, Frenkel, and Rybnikov [FFRy] show that the algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ acts cyclically on the tensor product $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z}):=V_{1}(z_{1})\otimes\cdots\otimes V_{\ell}(z_{\ell})$ of evaluation ${\mathfrak{g}}[t]$-modules for any regular $\mu$ and any pairwise distinct $z_{1},\ldots,z_{\ell}$.
+They also show that $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable with a simple spectrum on $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for generic $\mu$ and generic $\bf z$}.
+Now let ${\mathfrak{g}}={\mathfrak{gl}}_{d}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$.
+It possesses a *-structure ${\bm{\sigma}}:{\mathfrak{g}}\longrightarrow{\mathfrak{g}}$, i.e., $\bm{\sigma}$} is an even anti-linear anti-involution (see Section 3.3).
+Let $M_{1},\ldots,M_{\ell}$ be unitarizable $\mathfrak{g}$}-modules with respect to $\bm{\sigma}$}, and let $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})=M_{1}(z_{1})\otimes\cdots\otimes M_{\ell}(z_{\ell})$.
+Define ${\mathfrak{g}}^{*}_{{\bm{\sigma}}}=\left\{\,\mu\in{\mathfrak{g}}^{*}\;\middle|\;\mu\circ{\bm{\sigma}}=\overline{\mu}\,\right\}$, where $\overline{\mu}\in{\mathfrak{g}}^{*}$ is the complex conjugation of $\mu$.
+Using the properties of *-structures and Feigin–Frenkel centers, we obtain the following.
+Theorem 1.1 (Theorem 3.10).
+The Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable on any finite-dimensional $\mathcal{B}_{\mathfrak{g}}^{\mu}$-submodule of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for any $\mu\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$ and $\bf z$$\in{\mathbb{R}}^{\ell}$.
+For ${\bf w}:=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$, let $\mu_{\bf w}\in{\mathfrak{g}}^{*}$ be the element defined by (3.12) and (3.13).
+In other words, $\mu_{\bf w}$ corresponds to a diagonal matrix.
+We write $\mathcal{B}_{\mathfrak{g}}^{\bf w}:=\mathcal{B}_{\mathfrak{g}}^{\mu_{\bf w}}$.
+Let $L_{1},\ldots,L_{\ell}$ be unitarizable highest weight $\mathfrak{g}$}-modules with respect to $\bm{\sigma}$}.
+Set $\underline{L}=L_{1}\otimes\cdots\otimes L_{\ell}$ and $\underline{L}({\bf z})=L_{1}(z_{1})\otimes\cdots\otimes L_{\ell}(z_{\ell})$.
+Let $\mathfrak{h}$} denote the standard Cartan subalgebra of $\mathfrak{g}$} (see Section 2.1.4).
+Note that for any weight $\gamma\in{\mathfrak{h}}^{*}$ of $\underline{L}$, the $\gamma$-weight space $\underline{L}({\bf z})_{\gamma}$ of $\underline{L}({\bf z})$ is a finite-dimensional $\mathcal{B}_{\mathfrak{g}}^{\bf w}$-module, and that $\mu_{\bf w}\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$ for ${\bf w}\in{\mathbb{R}}^{d}$.
+We have the following as a consequence of Theorem 1.1.
+Corollary 1.2 (Corollary 3.12).
+Let $\gamma\in{\mathfrak{h}}^{*}$ be any weight of $\underline{L}$.
+Then $\mathcal{B}_{\mathfrak{g}}^{\bf w}$ is diagonalizable on $\underline{L}({\bf z})_{\gamma}$ for any ${\bf w}\in{\mathbb{R}}^{d}$ and $\bf z$$\in{\mathbb{R}}^{\ell}$.
+A natural question is whether there are unitarizable analogs of the results of [FFRy] mentioned above.
+Conjecture 1.3 (Conjecture 3.13).
+Let $\gamma\in{\mathfrak{h}}^{*}$ be any weight of $\underline{L}$.
+Then the weight space $\underline{L}({\bf z})_{\gamma}$ is a cyclic $\mathcal{B}_{\mathfrak{g}}^{\bf w}$-module for any pairwise distinct $w_{1},\ldots,w_{d}\in{\mathbb{C}}$ and any pairwise distinct $z_{1},\ldots,z_{\ell}\in{\mathbb{C}}$.
+Moreover, $\mathcal{B}_{\mathfrak{g}}^{\bf w}$ is diagonalizable with a simple spectrum on $\underline{L}({\bf z})_{\gamma}$ for generic $\bf w$} and generic $\bf z$}.
+Theorem 1.5 (see below) provides positive evidence for Conjecture 1.3 when $q$ and $n$ are set to 0.
+1.2. A duality for the general linear Lie (super)algebras
+For ${\bf w}:=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$ and ${\bf z}:=(z_{1},\ldots,z_{p+q+m+n})\in{\mathbb{C}}^{p+q+m+n}$, we consider the Bethe algebra $\mathcal{B}_{d}^{\bf w}:=\mathcal{B}_{{\mathfrak{gl}}_{d}}^{\bf w}$ for the general linear Lie algebra ${\mathfrak{gl}}_{d}$ with respect to $\bf w$} and the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}:={\mathcal{B}_{{\mathfrak{gl}}_{p+m|q+n}}^{\bf z}}$ for the general linear Lie superalgebra ${\mathfrak{gl}}_{p+m|q+n}$ with respect to $\bf z$}; see Section 4.2.
+Note that $\mathcal{B}_{d}^{\bf w}$ is a subalgebra of $U({\mathfrak{gl}}_{d}[t])$ generated by the coefficients of the series ${\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ while $\mathcal{B}_{p+m|q+n}^{\bf z}$} is a subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$ generated by the coefficients of the series ${\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$.
+Here $\rm cdet$} and ${\rm Ber}^{\bf s}$ are respectively the column determinant and the Berezinian of type ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})$ (see Section 4.1), and ${\mathcal{L}}_{d}^{\bf w}$ and ${\mathcal{L}}_{{p+m|q+n}}^{\bf z}$ are respectively the matrices defined in (4.4) and (4.7).
+Let $\mathcal{F}$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$, and $\zeta^{a}_{s}$, for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$, and $a=1,\ldots,d$.
 Here $x^{a}_{i}$ and $y^{a}_{r}$ are even variables while $\eta^{a}_{j}$ and $\zeta^{a}_{s}$ are odd.
-The superalgebra $\mathcal{F}$ can be realized as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators.
-Let $\mathcal{D}:={\mathbb{D}}[{\bf x},{\bf y},{\boldsymbol{\eta}},{\boldsymbol{\zeta}}]$ be the corresponding Weyl superalgebra, i.e., $\mathcal{D}$ is the superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$ and $\zeta^{a}_{s}$ as well as their derivatives $\frac{\partial}{\partial x_{i}^{a}}$, $\frac{\partial}{\partial\eta_{j}^{a}}$, $\frac{\partial}{\partial y_{r}^{a}}$ and $\frac{\partial}{\partial\zeta_{r}^{a}}$.
-There is a natural action of $\mathcal{D}$ on $\mathcal{F}$, and hence $\mathcal{D}$ is a subalgebra of the endomorphism algebra ${\rm End}(\mathcal{F})$ of $\mathcal{F}$.
-Moreover, there exist superalgebra homomorphisms $\phi:U({\mathfrak{gl}}_{d})\longrightarrow\mathcal{D}$ and $\overline{\phi}:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$ from the universal enveloping algebras of ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$ to $\mathcal{D}$ defined by (4.2) and (4.3), respectively.
-These maps induce a ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-action on $\mathcal{F}$.
-It is well-known that $\mathcal{F}$ decomposes into a direct sum of unitarizable ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-modules.
-This is called the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ ([CLZ]).
-Such unitarizable modules are, in general, infinite-dimensional.
-Taking $p=q=0$, the maps $\phi$ and $\overline{\phi}$ specialize to the superalgebra homomorphisms $U({\mathfrak{gl}}_{d})\longrightarrow{\mathbb{D}}[{\bf x},{\boldsymbol{\eta}}]$ and $U({\mathfrak{gl}}_{m|n})\longrightarrow{\mathbb{D}}[{\bf x},{\boldsymbol{\eta}}]$, respectively, and the duality reduces to the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$, which yields a decomposition of ${\mathbb{C}}[{\bf x},{\boldsymbol{\eta}}]$ into a direct sum of finite-dimensional ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{m|n}$-modules ([CW, Se1, Se2]).
-Assuming further that $n=0$, it gives the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ ([H]).
-Fix ${\bf w}\in{\mathbb{C}}^{\ell}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$.
-There are natural extensions of $\phi$ and $\overline{\phi}$ to the universal enveloping algebras of current algebras, and we have superalgebra homomorphisms $\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])\longrightarrow\mathcal{D}$ and $\overline{\phi}_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])\longrightarrow\mathcal{D}$; see (4.4) and (4.5).
-It is natural to ask if dualities exist in the context of current algebras.
-When $p=q=n=0$, a duality phenomenon is noticed in [TL] (see also [TV]).
-In [MTV3], a precise statement of the duality between $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{m}[t])$ is made and proved.
-So far, various generalizations have emerged (see [HM, TU, VY]).
-All of these are related to the Gaudin models for general linear Lie (super)algebras.
-We will establish a duality between $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$ in this paper.
-We consider the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ for ${\mathfrak{gl}}_{d}$ with respect to ${\bf w}$ and the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ for ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}$; see Section 3.3 and Section 3.4.
-As the algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ are respectively subalgebras of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$, the maps $\phi_{\bf z}$ and $\overline{\phi}_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ and an action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on the polynomial superalgebra $\mathcal{F}$, respectively.
-We show that these two actions are equivalent.
-More precisely, we have:
+The superalgebra $\mathcal{F}$ can be realized as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators (cf.
+[CL, LZ]).
+Let $\mathcal{D}$ be the corresponding Weyl superalgebra, i.e., $\mathcal{D}$ is the associative unital superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$, and $\zeta^{a}_{s}$ as well as their derivatives $\frac{\partial}{\partial x_{i}^{a}}$, $\frac{\partial}{\partial\eta_{j}^{a}}$, $\frac{\partial}{\partial y_{r}^{a}}$, and $\frac{\partial}{\partial\zeta_{r}^{a}}$.
+There exist superalgebra homomorphisms $\phi:U({\mathfrak{gl}}_{d})\longrightarrow\mathcal{D}$ and $\varphi:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$; see (4.9) and (4.10).
+They extend to superalgebra homomorphisms $\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$ and $\varphi_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$, respectively.
+Here, for instance, $\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$ is the superalgebra of pseudo-differential operators over $\mathcal{D}$; see Section 2.3.
+The maps $\phi_{\bf z}$ and $\varphi_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ and an action of $\mathcal{B}_{p+m|q+n}^{\bf z}$} on $\mathcal{F}$, respectively.
+The following theorem asserts that the actions are equivalent.
+We call it the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$.
 Theorem 1.4 (Theorem 4.10).
-$\phi_{\bf z}\big{(}\mathcal{B}_{d}^{\bf w}\big{)}=\overline{\phi}_{\bf w}\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}$.
-We name Theorem 1.1 the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$.
-It recovers several dualities established earlier.
+The coefficients of $\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ are determined by those of $\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$, and vice versa.
+Consequently, $\phi_{\bf z}\big(\mathcal{B}_{d}^{\bf w}\big)=\varphi_{\bf w}\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)$.
+Theorem 1.4 recovers several dualities established earlier.
 After specializing to $p=q=0$, it recovers the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$ due to Huang and Mukhin [HM].
 If we assume further that $n=0$ (resp., $m=0$), Theorem 4.10 gives the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ obtained by Mukhin, Tarasov, and Varchenko [MTV3] (resp., a variant of the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{n})$ obtained by Tarasov and Uvarov [TU]).
-Furthermore, we will show in [ChL3] that there is another duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for Gaudin models with irregular singularities, which specializes to a variant of Theorem 1.1.
-The Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ enables us to study the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on $\mathcal{F}$ via the action of $\mathcal{B}_{d}^{\bf w}$ on $\mathcal{F}$, and vice versa.
-Using the properties of the action of $\mathcal{B}_{d}^{\bf w}$ on finite-dimensional ${\mathfrak{gl}}_{d}$-modules, we will better understand the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules.
-Let $\underline{M\mkern-4.0mu}\mkern 4.0mu:=M_{1}\otimes\cdots\otimes M_{d}$, where each $M_{i}$ is an infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module with highest weight corresponding to a generalized partition of depth 1 (see Section 5.2 for a detailed description).
-We will see in Theorem 5.1 that each weight space of the evaluation ${\mathfrak{gl}}_{p+m|q+n}[t]$-module $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})$ is a cyclic ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module, and that ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ is diagonalizable with a simple spectrum on the weight space for generic ${\bf w}$ and ${\bf z}$.
-Let us explain a little how we prove Theorem 1.1.
-The Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ are determined by the column determinant ${\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})$ and the Berezinian ${\rm Ber}^{\bf s}\,({\mathcal{L}}_{{p+m|q+n}}^{\bf z})$ of type ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})$, respectively; see (3.6) and (3.9).
-In the special case where $p=q=0$, the specializations of the series $\phi_{\bf z}({\rm cdet}\,({\mathcal{L}}_{d}^{\bf w}))$ and $\overline{\phi}_{\bf w}({\rm Ber}^{\bf s}\,({\mathcal{L}}_{{p+m|q+n}}^{\bf z}))$ admit Capelli-type identities [HM, Proposition 5.4 and Proposition 5.6], and the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$ is obtained by matching the identities.
-However, the situation is somewhat complicated in our case as no such identities are known to exist.
-Based on the properties of Berezinians, we can relate $\overline{\phi}_{\bf w}({\rm Ber}^{\bf s}\,({\mathcal{L}}_{{p+m|q+n}}^{\bf z}))$ to $\phi_{\bf z}({\rm cdet}\,({\mathcal{L}}_{d}^{\bf w}))$ (for general $p,q,m,n$) by examining an anti-involution on the Weyl superalgebra $\mathcal{D}$ and an anti-involution on the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ (see Proposition 4.5).
-Theorem 1.1 will then follow.
+Also, another duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ is obtained by Vicedo and Young [VY] for Gaudin models with irregular singularities.
+It is extended by the authors [ChL3] to a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$, which specializes to a variant of Theorem 1.4.
+Using the properties of the actions of $\mathcal{B}_{d}^{\bf w}$ on tensor products of finite-dimensional ${\mathfrak{gl}}_{d}$-modules, Theorem 1.4 enables us to better understand the actions of $\mathcal{B}_{p+m|q+n}^{\bf z}$} on tensor products of infinite-dimensional unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules, as shown in the following theorem.
+Theorem 1.5 (Theorem 4.13).
+Let $\underline{L}=L_{1}\otimes\cdots\otimes L_{d}$ and $\underline{L}({\bf w})=L_{1}(w_{1})\otimes\cdots\otimes L_{d}(w_{d})$, where the $L_{i}$ are (infinite-dimensional) unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules corresponding to generalized partitions of depth 1 (see (4.17)).
+Let $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ be any weight of $\underline{L}$ (where ${\mathfrak{h}}_{{p+m|q+n}}$ is the Cartan subalgebra of ${\mathfrak{gl}}_{p+m|q+n}$ defined in Section 2.1.1).
+Then the weight space $\underline{L}({\bf w})_{\mu}$ is a cyclic $\mathcal{B}_{p+m|q+n}^{\bf z}$}-module for any pairwise distinct $z_{1},\ldots,z_{p+q+m+n}\in{\mathbb{C}}$ and any pairwise distinct $w_{1},\ldots,w_{d}\in{\mathbb{C}}$.
+Moreover, $\mathcal{B}_{p+m|q+n}^{\bf z}$} is diagonalizable with a simple spectrum on $\underline{L}({\bf w})_{\mu}$ for generic $\bf z$} and generic $\bf w$}.
+1.3. Outline
 We organize the paper as follows.
-In Section 2, we set up our notations.
-We recall the notions of the universal affine vertex algebra at the critical level, the Feigin–Frenkel center and the Bethe algebra.
-In Section 3, we discuss the Bethe algebras for general linear Lie (super)algebras and their basic properties.
-In Section 4, we introduce the joint action of the Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on the polynomial superalgebra $\mathcal{F}$ for any sequences ${\bf w}\in{\mathbb{C}}^{\ell}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$.
-We prove Theorem 1.1, i.e., the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$.
-In Section 5, we apply the Bethe duality to the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules and obtain Theorem 5.1.
+Section 2 is devoted to reviewing the classical Lie (super)algebras of types $\mathfrak{a,c,d}$ and some background material needed in this paper.
+In Section 3, we investigate the action of the Bethe algebra for $\mathfrak{g}$} on a tensor product of evaluation ${\mathfrak{g}}[t]$-modules, where $\mathfrak{g}$} is the classical Lie algebra of type $\mathfrak{a,c,d}$.
+We prove Theorem 1.1 and obtain Corollary 1.2.
+In Section 4, we establish Theorem 1.4, i.e., the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$.
+We give an application of the duality and demonstrate Theorem 1.5, which provides positive evidence for Conjecture 1.3.
+In Appendix A, we prove Proposition 4.7, which gives an equality relating $\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ to $\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$.
 Notations.
-Throughout the paper, the symbol ${\mathbb{Z}}$ (resp., ${\mathbb{N}}$ and ${\mathbb{Z}}_{+}$) stands for the set of all (resp., positive and non-negative) integers, the symbol ${\mathbb{C}}$ for the field of complex numbers, and the symbol ${\mathbb{Z}}_{2}:=\{\bar{0},\bar{1}\}$ for the field of integers modulo 2.
-All vector spaces, algebras, tensor products, etc., are over ${\mathbb{C}}$.
+Throughout the paper, the symbol $\mathbb{Z}$} (resp., $\mathbb{N}$} and $\mathbb{Z}$$_{+}$) stands for the set of all (resp., positive and non-negative) integers, the symbol $\mathbb{C}$} (resp., $\mathbb{R}$}) for the field of complex (resp., real) numbers, and the symbol ${\mathbb{Z}}_{2}:=\{\bar{0},\bar{1}\}$ for the field of integers modulo 2.
+Unless otherwise stated, all vector spaces, algebras, tensor products, etc., are over $\mathbb{C}$}.
 We fix $d\in{\mathbb{N}}$ and $p,q,m,n\in{\mathbb{Z}}_{+}$.
 2. Preliminaries
-In this section, we set up our notations for the general linear Lie (super)algebras and their related concepts.
-We review several notions needed in this paper, including current algebras, the Feigin–Frenkel centers and the Bethe algebras.
-2.1. The general linear Lie (super)algebras
+In this section, we review the classical Lie (super)algebras of types $\mathfrak{a,c,d}$, evaluation modules, and pseudo-differential operators.
+2.1. Classical Lie (super)algebras
+2.1.1. The general linear Lie (super)algebra
 For $p,q,m,n\in{\mathbb{Z}}_{+}$ that are not all zero, let
-[…]
+$\mathbb{I}=\{i\in{\mathbb{N}}\,|\,1\leq i\leq p+q+m+n\}.$
 For $i\in\mathbb{I}$, define $|i|\in{\mathbb{Z}}_{2}$ by
 $|i|=\begin{cases}\,\bar{0}&\,\,\mbox{if}\ \ i\in\{1,\ldots,p\}\cup\{p+q+1,\ldots,p+q+m\},\\ \,\bar{1}&\,\,\mbox{otherwise}.\end{cases}$
-[…]
+Let $\{e_{i}\,|\,i\in\mathbb{I}\}$ be a basis for the superspace ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ such that $\{e_{i}\,|\,1\leq i\leq p+q\}$ and $\{e_{p+q+i}\,|\,1\leq i\leq m+n\}$ are respectively the standard homogeneous bases for ${\mathbb{C}}^{p|q}$ and ${\mathbb{C}}^{m|n}$.
+In other words, the parity of $e_{i}$ is given by $|e_{i}|=|i|$ for $i\in\mathbb{I}$.
 For any $i,j\in\mathbb{I}$, let $E^{i}_{j}$ denote the $\mathbb{C}$}-linear endomorphism on ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ defined by
@@
 The parity of $E^{i}_{j}$ is given by $|E^{i}_{j}|=|i|+|j|$.
-The superspace of ${\mathbb{C}}$-linear endomorphisms on ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ has a natural structure of a Lie superalgebra, called a general linear Lie (super)algebra and denoted by ${\mathfrak{gl}}_{p+m|q+n}$.
+The superspace of $\mathbb{C}$}-linear endomorphisms on ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ has a natural structure of a Lie superalgebra, called the general linear Lie (super)algebra and denoted by ${\mathfrak{gl}}_{p+m|q+n}$.
 The commutation relations of ${\mathfrak{gl}}_{p+m|q+n}$ are given by
-$[E^{i}_{j},E^{k}_{\ell}]=\delta_{j,k}E^{i}_{\ell}-(-1)^{(|i|+|j|)(|k|+|\ell|)}\delta_{i,\ell}E^{k}_{j}\qquad\mbox{for $i,j,k,\ell\in\mathbb{I}$.}$
+$E^{i}_{j},E^{k}_{l}]=\delta_{j,k}E^{i}_{l}-(-1)^{(|i|+|j|)(|k|+|l|)}\delta_{i,l}E^{k}_{j}\qquad\mbox{for $i,j,k,l\in\mathbb{I}$.}$
 Note that $\{E^{i}_{j}\,|\,i,j\in\mathbb{I}\}$ is a homogeneous basis for ${\mathfrak{gl}}_{p+m|q+n}$.
 Let $\mathfrak{b}_{{p+m|q+n}}=\bigoplus_{\begin{subarray}{c}i,j\in\mathbb{I},i\leq j\end{subarray}}{\mathbb{C}}E^{i}_{j}$ be a Borel subalgebra of ${\mathfrak{gl}}_{p+m|q+n}$.
-The corresponding Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ has a basis $\{E^{i}_{i}\,|\,i\in\mathbb{I}\}$, and the dual basis in ${\mathfrak{h}}_{{p+m|q+n}}^{*}$ is denoted by $\{\epsilon_{i}\,|\,i\in\mathbb{I}\}$, where the parity of $\epsilon_{i}$ is given by $|\epsilon_{i}|=|i|$.
-Note that the Borel subalgebra $\mathfrak{b}_{{p+m|q+n}}$ is not the standard one unless some of $p,q,m,n$ are set to 0.
-For $m=d$ and $p=q=n=0$, we write
-${\mathfrak{gl}}_{d}={\mathfrak{gl}}_{d|0},\quad{\mathfrak{b}}_{d}={\mathfrak{b}}_{d|0},\quad{\mathfrak{h}}_{d}={\mathfrak{h}}_{d|0},$
-and
-$\hskip 28.45274pte_{ij}=E^{i}_{j}\qquad\mbox{for $i,j=1,\ldots,d$.}$
-2.2. Current algebras and evaluation modules
-For any Lie (super)algebra ${\mathfrak{g}}$ and an even variable $t$, let ${\mathfrak{g}}[t,t^{-1}]:={\mathfrak{g}}\otimes{\mathbb{C}}[t,t^{-1}]$ denote the loop algebra of ${\mathfrak{g}}$, which is a Lie superalgebra with commutation relations
+The corresponding Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ has a basis $\{E^{i}_{i}\,|\,i\in\mathbb{I}\}$.
+We denote the dual basis in ${\mathfrak{h}}_{{p+m|q+n}}^{*}$ by $\{\epsilon_{i}\,|\,i\in\mathbb{I}\}$.
+The parity of $\epsilon_{i}$ is given by $|\epsilon_{i}|=|i|$.
+For $m=d$ and $p=q=n=0$, we define
+${\mathfrak{gl}}_{d}:={\mathfrak{gl}}_{d|0}\qquad\text{and}\qquad e^{\mathfrak{a}}_{ij}:=e_{ij}:=E^{i}_{j}\qquad\mbox{for $i,j=1,\ldots,d$.}$ (2.1)
+2.1.2. The symplectic Lie algebra
+The subalgebra of ${\mathfrak{gl}}_{2d}$ that preserves the nondegenerate skew-symmetric bilinear form on ${\mathbb{C}}^{2d}$ corresponding to the matrix
+$\left[\begin{matrix}0&J_{d}\\ -J_{d}&0\end{matrix}\right]$
+is called the symplectic Lie algebra and is denoted by $\mathfrak{sp}_{2d}$.
+Hereafter, $J_{d}$ is the $d\times d$ reversion matrix defined by $(J_{d})_{ij}=1$ if $i+j=d+1$ and $(J_{d})_{ij}=0$ otherwise.
+The Lie algebra $\mathfrak{sp}_{2d}$ is simple and is spanned by
+$e^{\mathfrak{c}}_{ij}:=e_{ij}-\varepsilon_{i}\varepsilon_{j}e_{j^{\prime}i^{\prime}}$ (2.2)
+for $i,j=1,\ldots,2d$, where $i^{\prime}:=2d-i+1$, and $\varepsilon_{i}:=1$ if $1\leq i\leq d$ and $\varepsilon_{i}:=-1$ if $d+1\leq i\leq 2d$.
+2.1.3. The orthogonal Lie algebra
+The subalgebra of ${\mathfrak{gl}}_{2d}$ that preserves the nondegenerate symmetric bilinear form on ${\mathbb{C}}^{2d}$ corresponding to the matrix
+$\left[\begin{matrix}0&J_{d}\\ J_{d}&0\end{matrix}\right]$
+is called the orthogonal Lie algebra and is denoted by $\mathfrak{so}_{2d}$. It is spanned by
+$e^{\mathfrak{d}}_{ij}:=e_{ij}-e_{j^{\prime}i^{\prime}}$ (2.3)
+for $i,j=1,\ldots,2d$, where $i^{\prime}:=2d-i+1$, and is a simple Lie algebra if $d\geq 3$.
+2.1.4. Triangular decompositions
+The general linear Lie (super)algebras ${\mathfrak{gl}}_{p+m|q+n}$ and ${\mathfrak{gl}}_{d}$ are called the classical Lie (super)algebras of types $\mathfrak{a}$ while the Lie algebras $\mathfrak{sp}_{2d}$ and $\mathfrak{so}_{2d}$ are called the classical Lie algebras of types $\mathfrak{c}$ and $\mathfrak{d}$, respectively.
+Let
+${\mathfrak{g}}^{\mathfrak{a}}={\mathfrak{gl}}_{d},\qquad{\mathfrak{g}}^{\mathfrak{c}}=\mathfrak{sp}_{2d},\qquad{\mathfrak{g}}^{\mathfrak{d}}=\mathfrak{so}_{2d}.$
+For ${\mathfrak{x}}=\mathfrak{a,c,d}$, the classical Lie algebra ${\mathfrak{g}}^{\mathfrak{x}}$ has the triangular decomposition
+${\mathfrak{g}}^{\mathfrak{x}}=\mathfrak{n}_{-}^{\mathfrak{x}}\oplus{\mathfrak{h}}^{\mathfrak{x}}\oplus\mathfrak{n}_{+}^{\mathfrak{x}},$ (2.4)
+where ${\mathfrak{h}}^{\mathfrak{x}}:=\bigoplus_{i=1}^{d}{\mathbb{C}}e^{\mathfrak{x}}_{ii}$ is the standard Cartan subalgebra of ${\mathfrak{g}}^{\mathfrak{x}}$, and $\mathfrak{n}_{+}^{\mathfrak{x}}:=\sum_{\begin{subarray}{c}i<j\end{subarray}}e^{\mathfrak{x}}_{ij}$ and $\mathfrak{n}_{-}^{\mathfrak{x}}:=\sum_{\begin{subarray}{c}i<j\end{subarray}}{\mathbb{C}}e^{\mathfrak{x}}_{ji}$ are nilpotent subalgebras of ${\mathfrak{g}}^{\mathfrak{x}}$.
+We will suppress the superscript $\mathfrak{x}$} if there is no danger of confusion.
+2.2. Evaluation modules
+For any Lie superalgebra $\mathfrak{g}$}, we denote by $U({\mathfrak{g}})$ the universal enveloping algebra of $\mathfrak{g}$}.
+For any even variable $t$, the loop algebra ${\mathfrak{g}}[t,t^{-1}]:={\mathfrak{g}}\otimes{\mathbb{C}}[t,t^{-1}]$ is defined to be the Lie superalgebra with commutation relations
 $\hskip 28.45274pt\big[A\otimes t^{r},B\otimes t^{s}\big]=[A,B]\otimes t^{r+s}\qquad\mbox{for $A,B\in{\mathfrak{g}}\,$ and $\,r,s\in{\mathbb{Z}}$.}$
@@
 $\hskip 28.45274ptA[r]:=A\otimes t^{r}\qquad\mbox{for $A\in{\mathfrak{g}}\,$ and $\,r\in{\mathbb{Z}}$.}$
-The current algebra ${\mathfrak{g}}[t]:={\mathfrak{g}}\otimes{\mathbb{C}}[t]$ is a subalgebra of ${\mathfrak{g}}[t,t^{-1}]$ defined in an obvious way.
-We identify the Lie (super)algebra ${\mathfrak{g}}$ with the subalgebra ${\mathfrak{g}}\otimes 1$ of constant polynomials in ${\mathfrak{g}}[t]$ and hence $U({\mathfrak{g}})\subseteq U({\mathfrak{g}}[t])$.
-Hereafter $U(\mathcal{G})$ denotes the universal enveloping algebra of $\mathcal{G}$ for any Lie (super)algebra $\mathcal{G}$.
+The current algebra ${\mathfrak{g}}[t]:={\mathfrak{g}}\otimes{\mathbb{C}}[t]$ and the superalgebra $t^{-1}{\mathfrak{g}}[t^{-1}]:={\mathfrak{g}}\otimes t^{-1}{\mathbb{C}}[t^{-1}]$ are subalgebras of ${\mathfrak{g}}[t,t^{-1}]$ defined in an obvious way.
+We identify $\mathfrak{g}$} with the subalgebra ${\mathfrak{g}}\otimes 1$ of constant polynomials in ${\mathfrak{g}}[t]$ and hence $U({\mathfrak{g}})\subseteq U({\mathfrak{g}}[t])$.
 For $\alpha\in{\mathbb{C}}$, we have an evaluation homomorphism ${\rm ev}_{\alpha}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}})$ given by
 ${\rm ev}_{\alpha}(A[r])=\alpha^{r}A\qquad\mbox{for $A\in{\mathfrak{g}}\,$ and $\,r\in{\mathbb{Z}}_{+}$}.$
-Any ${\mathfrak{g}}$-module $M$ is a ${\mathfrak{g}}[t]$-module via ${\rm ev}_{\alpha}$, which is called the evaluation module and denoted by $M(\alpha)$.
+Any $\mathfrak{g}$}-module $M$ is a ${\mathfrak{g}}[t]$-module via ${\rm ev}_{\alpha}$, called the evaluation module and denoted by $M(\alpha)$.
 For any $A\in{\mathfrak{g}}$ and any even variable $z$, we set
@@
 $A(z)v=\frac{Av}{z-\alpha}\qquad\mbox{for $v\in M(\alpha)$.}$
-Here every rational function in $z$ represents its power series expression at $\infty$.
+Here each rational function in $z$ represents its power series expression at $\infty$.
 Let $\ell\in{\mathbb{N}}$ and ${\bf z}=(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}$.
@@
 We have the homomorphism
-$\underline{{\rm ev}}_{\bf z}:=({\rm ev}_{z_{1}}\otimes\ldots\otimes{\rm ev}_{z_{\ell}})\circ\Delta^{(\ell-1)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}})^{\otimes{\ell}},$
+$\underline{{\rm ev}}_{\bf z}:=({\rm ev}_{z_{1}}\otimes\ldots\otimes{\rm ev}_{z_{\ell}})\circ\Delta^{(\ell-1)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}})^{\otimes{\ell}},$ (2.5)
 given by
@@
 $\hskip 42.67912ptA(z)v=\sum_{i=1}^{\ell}\frac{A^{(i)}v}{z-z_{i}}\qquad\mbox{for $v\in\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$}.$
-2.3. The Feigin–Frenkel center
-The Lie algebra ${\mathfrak{gl}}_{d}$ admits a nondegenerate invariant symmetric bilinear form
-$(A,B)={\rm tr}(AB)-\frac{1}{d}{\rm tr}(A){\rm tr}(B),\quad\mbox{for $A,B\in{\mathfrak{gl}}_{d}$},$
-where ${\rm tr}$ denotes the trace. The affine Lie superalgebra
-$\widehat{{\mathfrak{gl}}_{d}}:={\mathfrak{gl}}_{d}[t,t^{-1}]\oplus{\mathbb{C}}K,$
-where $K$ is even and central, is a Lie superalgebra with commutation relations
-$\big{[}A[r],B[s]\big{]}=[A,B][r+s]+r\delta_{r+s,0}(A,B)K$
-for $A,B\in{\mathfrak{gl}}_{d}$ and $r,s\in{\mathbb{Z}}$.
-The vacuum module $V_{{\rm crit}}({\mathfrak{gl}}_{d})$ at the critical level is
-$V_{{\rm crit}}({\mathfrak{gl}}_{d}):=U(\widehat{{\mathfrak{gl}}_{d}})/I,$
-where $I$ is the left ideal of $U(\widehat{{\mathfrak{gl}}_{d}})$ generated by ${\mathfrak{gl}}_{d}[t]$ and $K+d$.
-It possesses a unique vertex algebra structure, called the universal affine vertex algebra associated to ${\mathfrak{gl}}_{d}$ at the critical level; see [FBZ, Theorem 4.4.1].
-The center of the vertex algebra $V_{{\rm crit}}({\mathfrak{gl}}_{d})$ is given by
-${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d}):=\left\{\,\!v\in V_{{\rm crit}}({\mathfrak{gl}}_{d})\;\middle|\;{\mathfrak{gl}}_{d}[t]v=0\!\,\right\},$
-called the Feigin–Frenkel center. Any element of ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$ is called a Segal–Sugawara vector.
-We also have the universal affine vertex algebra $V_{{\rm crit}}({\mathfrak{gl}}_{p+m|q+n})$ at the critical level and the Feigin–Frenkel center ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$ for ${\mathfrak{gl}}_{p+m|q+n}$.
-We skip details here; see [MR, Section 1.3] (and also [ChL2, Section 3.1]).
-Let ${\mathfrak{g}}$ be a general linear Lie (super)algebra. The algebra $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ is equipped with the (even) derivation ${\rm T}:=-d/dt$ defined by
-${\rm T}(1)=0\qquad\text{and}\qquad{\rm T}(A[-r])=rA[-r-1]$
-for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{N}}$.
-Note that $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ is isomorphic to $V_{{\rm crit}}({\mathfrak{g}})$ as superspaces by the Poincaré–Birkhoff–Witt theorem, and that the derivation ${\rm T}$ corresponds to the translator operator on $V_{{\rm crit}}({\mathfrak{g}})$.
-It is known that ${\mathfrak{z}}(\widehat{{\mathfrak{g}}})$ can be viewed as a commutative subalgebra of $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ and is ${\rm T}$-invariant.
-We refer the reader to [Fr, FBZ, Mo, MR] for further details.
+2.3. Pseudo-differential operators
+Let $\mathcal{A}$ be an associative unital superalgebra and $z$ an even variable.
+We denote by $\mathcal{A}[\kern-1.1pt[z]\kern-1.1pt]$ (resp., $\mathcal{A}(\!(z)\!)$) the superalgebra of formal power series (resp., formal Laurent series) in $z$ with coefficients in $\mathcal{A}$.
+Let $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ denote the set of all formal series of the form
+$\sum_{j=-\infty}^{s}\sum_{i=-\infty}^{r}a_{ij}z^{i}\partial_{z}^{j},$
+where $r,s\in{\mathbb{Z}}$ and $a_{ij}\in\mathcal{A}$. We may endow $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ with a superalgebra structure using the rules:
+$\partial_{z}\partial_{z}^{-1}=\partial_{z}^{-1}\partial_{z}=1,\quad\partial_{z}^{i}z^{j}=\sum_{k=0}^{\infty}\binom{i}{k}\binom{j}{k}k!\,z^{j-k}\,\partial_{z}^{i-k},\quad\mbox{for $i,j\in{\mathbb{Z}}$.}$ (2.8)
+Here, e.g., ${\binom{i}{k}:=\frac{i(i-1)\ldots(i-k+1)}{k!}}$.
+The superalgebra $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ is called the superalgebra of pseudo-differential operators over $\mathcal{A}$.
+Let $\omega$ be an anti-involution on $\mathcal{A}$.
+In particular, $\omega^{2}=1$ and $\omega(ab)=\omega(b)\omega(a)$ for $a,b\in\mathcal{A}$.
+We readily see that $\omega$ extends to a $\mathbb{C}$}-linear automorphism $\omega:\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ defined by
+$\omega(az^{i}\partial_{z}^{j})=\omega(a)z^{j}\partial_{z}^{i}$ (2.9)
+for $a\in\mathcal{A}$ and $i,j\in{\mathbb{Z}}$. It is straightforward to prove the following.
+Proposition 2.1.
+The map $\omega$ is an anti-involution on $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$.
+3. Bethe algebras for classical Lie algebras
+In this section, we examine the behavior of the Feigin–Frenkel center and the Bethe algebra for the classical Lie algebra $\mathfrak{g}$} of type $\mathfrak{a,c,d}$ under a *-structure.
+Our primary purpose is to study the diagonalization of the Bethe algebra for $\mathfrak{g}$} on a tensor product of unitarizable $\mathfrak{g}$}-modules.
+3.1. Feigin–Frenkel centers
+Let $t$ be an even variable, and let $\mathfrak{g}$} be a simple Lie algebra or the general linear Lie (super)algebra.
+We denote by $V_{{\rm crit}}({\mathfrak{g}})$ the vacuum module at the critical level over the central extension $\widehat{{\mathfrak{g}}}:={\mathfrak{g}}[t,t^{-1}]\oplus{\mathbb{C}}K$ of the loop algebra ${\mathfrak{g}}[t,t^{-1}]$.
+By the Poincaré–Birkhoff–Witt theorem, $V_{{\rm crit}}({\mathfrak{g}})$ can be identified with $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ as (super)spaces.
+Moreover, the $\widehat{{\mathfrak{g}}}-$module structure on $V_{{\rm crit}}({\mathfrak{g}})$ allows us to define a vertex algebra structure on $V_{{\rm crit}}({\mathfrak{g}})$, called the universal affine vertex algebra at the critical level (see [Fr, FBZ, Mo2, MR] for details).
+[…]
+${\mathfrak{z}}(\widehat{{\mathfrak{g}}})=\left\{\,\!v\in V_{{\rm crit}}({\mathfrak{g}})\;\middle|\;{\mathfrak{g}}[t]v=0\!\,\right\}.$
+[…]
+${\rm T}(1)=0\qquad\text{and}\qquad{\rm T}(A[-r])=rA[-r-1]$ (3.1)
+[…]
+The following result is a famous theorem of Feigin and Frenkel.
+Theorem 3.1 ([FF]).
+[…]
+Remark 3.2.
+Theorem 3.1 holds as well for the general linear Lie algebra (see [CM, Theorem 3.1] (and also [CT]).
+Let ${\mathfrak{g}}={\mathfrak{gl}}_{d}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$, which has rank $d$.
+For ${\mathfrak{x}}=\mathfrak{a,c,d}$, recall the elements $e^{{\mathfrak{x}}}_{ij}$ defined in (2.1), (2.2) and (2.3), respectively.
+We denote by $\mathfrak{g}$$_{\mathbb{R}}$ the $\mathbb{R}$}-subspace of $\mathfrak{g}$} spanned by all the $e^{\mathfrak{x}}_{ij}$ and by $\mathfrak{h}$$_{\mathbb{R}}$ the $\mathbb{R}$}-subspace of $\mathfrak{h}$} spanned by all the $e^{\mathfrak{x}}_{ii}$.
+Let $U({\mathfrak{g}}[t])_{\mathbb{R}}$ be the $\mathbb{R}$}-subalgebra of $U({\mathfrak{g}}[t])$ generated by $A[r]$ for $A\in{\mathfrak{g}}_{\mathbb{R}}$ and $r\in{\mathbb{Z}}_{+}$, and let $U(t^{-1}{\mathfrak{g}}[t^{-1}])_{\mathbb{R}}$ be the $\mathbb{R}$}-subalgebra of $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ generated by A[r]A[-r] for $A\in{\mathfrak{g}}_{\mathbb{R}}$ and $r\in{\mathbb{N}}$.
+The $\mathbb{R}$}-algebras $U({\mathfrak{h}}[t])_{\mathbb{R}}$ and $U(t^{-1}{\mathfrak{h}}[t^{-1}])_{\mathbb{R}}$ are defined analogously.
+By [CM, Theorem 3.1] for type $\mathfrak{a}$ and [Mo3, Theorem 2.3] for type $\mathfrak{c,d}$ (see also [Mo1, Ya]), there exist
+[…]
+[…]
 3.2. Bethe algebras
-For $\mu\in{\mathfrak{gl}}_{d}^{*}$, there is an algebra homomorphism
-$\Psi^{\mu}:U(t^{-1}{\mathfrak{gl}}_{d}[t^{-1}])\longrightarrow U({\mathfrak{gl}}_{d}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$
+Let $\mathfrak{g}$} be a simple Lie algebra or the general linear Lie algebra. For any $\mu\in{\mathfrak{g}}^{*}$, there is an algebra homomorphism
+$\Psi^{\mu}:U(t^{-1}{\mathfrak{g}}[t^{-1}])\longrightarrow U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$
 given by
-$\hskip 28.45274pt\Psi^{\mu}(A[-r])=A\otimes(t-z)^{-r}+\delta_{r,1}\mu(A),\quad\mbox{for $A\in{\mathfrak{gl}}_{d}\,$ and $\,r\in{\mathbb{N}}$.}$ (2.3)
-The Bethe algebra $\mathcal{B}_{d}^{\mu}$ of ${\mathfrak{gl}}_{d}$ with respect to $\mu$ is defined to be the subalgebra of $U({\mathfrak{gl}}_{d}[t])$ generated by the coefficients of $\Psi^{\mu}(S)$ for $S\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$.
-Since ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$ is commutative, so is $\mathcal{B}_{d}^{\mu}$ ([FFTL, Ry]).
-For any $\mathcal{B}_{d}^{\mu}$-module $V$, let $(\mathcal{B}_{d}^{\mu})_{V}$ denote the image of the Bethe algebra $\mathcal{B}_{d}^{\mu}$ in the endomorphism algebra ${\rm End}(V)$ of $V$.
-We call $(\mathcal{B}_{d}^{\mu})_{V}$ the Bethe algebra of $V$ with respect to $\mu$.
-For $\alpha\in{\mathbb{C}}$, any ${\mathfrak{gl}}_{d}$-module $M$ is the evaluation module $M(\alpha)$ over ${\mathfrak{gl}}_{d}[t]$, and so $\mathcal{B}_{d}^{\mu}$ acts on $M(\alpha)$.
-If the centralizer ${\mathfrak{gl}}_{d}^{\mu}$ of $\mu$ in ${\mathfrak{gl}}_{d}$ contains the Cartan subalgebra ${\mathfrak{h}}_{d}$, then $\mathcal{B}_{d}^{\mu}$ commutes with ${\mathfrak{h}}_{d}$ and hence acts on any $\gamma$-weight space $M(\alpha)_{\gamma}$ of $M(\alpha)$, where $\gamma\in{\mathfrak{h}}_{d}^{*}$ ([Ry, Proposition 4]).
-Let
-${\boldsymbol{X}}_{\!\ell}=\left\{\,\!(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}\;\middle|\;z_{i}\not=z_{j}\,\,\mbox{for any $i\not=j$}\!\,\right\}$
-be the configuration space of $\ell$ distinct points on ${\mathbb{C}}^{\ell}$, and let ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$.
-Let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible ${\mathfrak{gl}}_{d}$-modules.
-The Bethe algebra $\mathcal{B}_{d}^{\mu}\subseteq U({\mathfrak{gl}}_{d}[t])$ acts on $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ (for the notation, see (2.2)) via the evaluation map $\underline{{\rm ev}}_{\bf z}$.
-An element $\mu\in{\mathfrak{gl}}_{d}^{*}$ is called regular if the dimension of ${\mathfrak{gl}}_{d}^{\mu}$ is $d$.
-The following remarkable results for $\mathcal{B}_{d}^{\mu}$, due to Feigin, Frenkel, and Rybnikov, will be used in this paper.
-Theorem 2.1 (cf. [FFRy, Corollary 5]).
-For any regular $\mu\in{\mathfrak{gl}}_{d}^{*}$ and ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ is a cyclic $\mathcal{B}_{d}^{\mu}$-module.
-Theorem 2.2 (cf. [FFRy, Corollary 6]).
-For generic $\mu\in{\mathfrak{gl}}_{d}^{*}$ and generic ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, the Bethe algebra $(\mathcal{B}_{d}^{\mu})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is diagonalizable with a simple spectrum.
-3. Bethe algebras for general linear Lie (super)algebras
-In this section, we recall column determinants, Berezinians and pseudo-differential operators. We discuss the Bethe algebras of the general linear Lie (super)algebras and their properties.
+$\Psi^{\mu}(A[-r])=A\otimes(t-z)^{-r}+\delta_{r,1}\mu(A),\qquad\mbox{for $A\in{\mathfrak{g}}\,$ and $\,r\in{\mathbb{N}}$.}$ (3.3)
+[…]
+[…]
+$\underline{{\rm ev}}_{\bf z}:U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]\longrightarrow U({\mathfrak{g}})^{\otimes\ell}[\kern-1.1pt[z^{-1}]\kern-1.1pt],$
+such that $\underline{{\rm ev}}_{\bf z}(z^{-1})=z^{-1}$. We have the commutative diagram
+$U(t^{-1}{\mathfrak{g}}[t^{-1}])$$U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$$U({\mathfrak{g}})^{\otimes\ell}[\kern-1.1pt[z^{-1}]\kern-1.1pt]$$\scriptstyle{\Psi^{\mu}_{\bf z}}$$\scriptstyle{\Psi^{\mu}}$$\scriptstyle{\underline{{\rm ev}}_{\bf z}}$
+where $\Psi^{\mu}_{\bf z}$ is defined by
+$\Psi^{\mu}_{\bf z}(A[-r])=\sum_{i=1}^{\ell}\frac{A^{(i)}}{(z_{i}-z)^{r}}+\delta_{r,1}\mu(A),\qquad\mbox{for $A\in{\mathfrak{g}}\,$ and $\,r\in{\mathbb{N}}$.}$
+[…]
+[…]
+$\big(\mathcal{A}_{{\mathfrak{g}}}^{\mu}({\bf z})\big)_{\underline{M\mkern-4.0mu}\mkern 4.0mu}=\big(\mathcal{B}_{\mathfrak{g}}^{\mu}\big)_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})}.$
+In this paper, we will restrict our attention to $\mathcal{B}_{\mathfrak{g}}^{\mu}$.
+For $\mu\in{\mathfrak{g}}^{*}$, we denote by ${\mathfrak{g}}^{\mu}$ the centralizer of $\mu$ in $\mathfrak{g}$}.
+Moreover, $\mu$ is called regular if ${\mathfrak{g}}^{\mu}$ has dimension equal to the rank of $\mathfrak{g}$}.
+Set
+${\bm{X}}_{\!\ell}=\left\{\,\!(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}\;\middle|\;z_{i}\not=z_{j}\,\,\mbox{for any $i\not=j$}\!\,\right\},$
+called the configuration space of $\ell$ distinct points on $\mathbb{C}$$^{\ell}$.
+Let $M_{1},\ldots,M_{\ell}$ be $\mathfrak{g}$}-modules, and let ${\bf z}\in{\bm{X}}_{\!\ell}$.
+If ${\mathfrak{h}}\subseteq{\mathfrak{g}}^{\mu}$, then $\mathcal{B}_{\mathfrak{g}}^{\mu}$ commutes with $\mathfrak{h}$} and hence acts on the $\gamma$-weight space $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for any $\gamma\in{\mathfrak{h}}^{*}$ ([Ry1, Proposition 4]).
+Let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible $\mathfrak{g}$}-modules.
+We have the following remarkable results about the action of $\mathcal{B}_{\mathfrak{g}}^{\mu}$ on $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$.
+Theorem 3.3 ([FFRy, Corollary 5]).
+The space $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ is a cyclic $\mathcal{B}_{\mathfrak{g}}^{\mu}$-module for any regular $\mu\in{\mathfrak{g}}^{*}$ and ${\bf z}\in{\bm{X}}_{\!\ell}$.
+Theorem 3.4 ([FFRy, Corollary 6]).
+The Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable with a simple spectrum on $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for generic $\mu\in{\mathfrak{g}}^{*}$ and generic ${\bf z}\in{\bm{X}}_{\!\ell}$.
+$3.$ *-structures
+A *-superalgebra is a pair $(\mathcal{A},{\bm{\sigma}})$, where $\mathcal{A}$ is an associative superalgebra, and ${\bm{\sigma}}:\mathcal{A}\longrightarrow\mathcal{A}$ is an even anti-linear anti-involution, called a *-structure on $\mathcal{A}$.
+A homomorphism $f:(\mathcal{A},{\bm{\sigma}})\longrightarrow(\mathcal{A}^{\prime},{\bm{\sigma}}^{\prime})$ of *-superalgebras is a homomorphism of superalgebras satisfying ${\bm{\sigma}}^{\prime}\circ f=f\circ{\bm{\sigma}}$.
+Let $(\mathcal{A},{\bm{\sigma}})$ be a *-superalgebra, and let $V$ be an $\mathcal{A}$-module.
+A Hermitian form $\langle\cdot|\cdot\rangle$ on $V$ is said to be contravariant if $\langle av_{1}|v_{2}\rangle=\langle v_{1}|{\bm{\sigma}}(a)v_{2}\rangle$ for all $a\in\mathcal{A}$ and $v_{1},v_{2}\in V$.
+An $\mathcal{A}$-module equipped with a positive definite contravariant Hermitian form is called a unitarizable $\mathcal{A}$-module.
+A Lie superalgebra with *-structure is a pair $(\mathcal{G},{\bm{\sigma}})$, where $\mathcal{G}$ is a Lie superalgebra, and ${\bm{\sigma}}:\mathcal{G}\longrightarrow\mathcal{G}$ is an even anti-linear anti-involution, called a *-structure on $\mathcal{G}$.
+Note that $\bm{\sigma}$} is a *-structure on $\mathcal{G}$ if and only if the natural extension of $\bm{\sigma}$} to $U(\mathcal{G})$ is a *-structure on $U(\mathcal{G})$.
+A homomorphism $f:(\mathcal{G},{\bm{\sigma}})\longrightarrow(\mathcal{G}^{\prime},{\bm{\sigma}}^{\prime})$ of Lie superalgebras with *-structures is a homomorphism of Lie superalgebras satisfying ${\bm{\sigma}}^{\prime}\circ f=f\circ{\bm{\sigma}}$.
+Let $(\mathcal{G},{\bm{\sigma}})$ be a Lie superalgebra with *-structure, and let $V$ be a $\mathcal{G}$-module.
+A Hermitian form $\langle\cdot|\cdot\rangle$ on $V$ is said to be contravariant if $\langle Av_{1}|v_{2}\rangle=\langle v_{1}|{\bm{\sigma}}(A)v_{1}\rangle$ for all $A\in\mathcal{G}$ and $v_{1},v_{2}\in V$.
+A $\mathcal{G}$-module equipped with a positive definite contravariant Hermitian form is called a unitarizable $\mathcal{G}$-module.
+It is clear that all unitarizable $\mathcal{G}$-modules are unitarizable $U(\mathcal{G})$-modules, and vice versa, and that all unitarizable $\mathcal{G}$-modules are completely reducible.
+We refer the reader to [ChL1, CLZ, LZ] for related discussions.
+The Lie superalgebra ${\mathfrak{gl}}_{p+m|q+n}$ admits a *-structure ${\bm{\sigma}}_{p}:{\mathfrak{gl}}_{p+m|q+n}\longrightarrow{\mathfrak{gl}}_{p+m|q+n}$ (see [CLZ, Section 3.2]), defined by
+[…]
+where $\bar{c}_{ij}$ denotes the complex conjugate of $c_{ij}\in{\mathbb{C}}$, and
+$i]:=\begin{cases}0&\mbox{if}\ \ 1\leq i\leq p,\\ 1&\mbox{otherwise}.\end{cases}$
+After specializing to $q=n=0$, we obtain a *-structure
+${\bm{\sigma}}_{p}:{\mathfrak{gl}}_{p+m}\longrightarrow{\mathfrak{gl}}_{p+m}.$ (3.5)
+Taking $p=m=d$, we obtain a *-structure ${\bm{\sigma}}_{d}$ on ${\mathfrak{gl}}_{2d}$. Evidently, the map ${\bm{\sigma}}_{d}$ restricts to *-structures
+${\bm{\sigma}}_{d}:\mathfrak{sp}_{2d}\longrightarrow\mathfrak{sp}_{2d}\qquad\text{and}\qquad{\bm{\sigma}}_{d}:\mathfrak{so}_{2d}\longrightarrow\mathfrak{so}_{2d}.$ (3.6)
+We remark that the Lie algebras with *-structures $({\mathfrak{gl}}_{p+m},{\bm{\sigma}}_{p})$, $(\mathfrak{sp}_{2d},{\bm{\sigma}}_{d})$, and $(\mathfrak{so}_{2d},{\bm{\sigma}}_{d})$ correspond to the Hermitian symmetric spaces of the noncompact type referred to as HS.1, HS.3, and HS.5, respectively, in [ES, p.
+37].
+For the rest of this section, we let ${\mathfrak{g}}={\mathfrak{gl}}_{p+m}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$. Set
+${\bm{\sigma}}=\begin{cases}{\bm{\sigma}}_{p}&\quad\mbox{if ${\mathfrak{g}}={\mathfrak{gl}}_{p+m}$,}\\ {\bm{\sigma}}_{d}&\quad\mbox{if ${\mathfrak{g}}=\mathfrak{sp}_{2d}$ or $\mathfrak{so}_{2d}$}.\end{cases}$ (3.7)
+For ${\mathfrak{x}}=\mathfrak{a,c,d}$, we clearly have ${\bm{\sigma}}(e^{{\mathfrak{x}}}_{ij})=(-1)^{[i]+[j]}e^{{\mathfrak{x}}}_{ji}$ for all $i,j$.
+In particular, ${\bm{\sigma}}(e^{{\mathfrak{x}}}_{ii})=e^{{\mathfrak{x}}}_{ii}$ for all $i$.
+Note that every non-trivial unitarizable highest weight $\mathfrak{g}$}-module with respect to $\bm{\sigma}$} is infinite-dimensional (see [EHW]).
+Let $k$ denote the dimension of $\mathfrak{g}$}. Consider the Segal–Sugawara vector
+[…]
+where $\{v_{1},\ldots,v_{k}\}$ is a basis for $\mathfrak{g}$}, and $\{v^{1},\ldots,v^{k}\}$ is the dual basis with respect to a nondegenerate invariant symmetric bilinear form on $\mathfrak{g}$}.
+Theorem 3.5 ([Ry2, Theorem 1]).
+[…]
+We denote by $U(t^{-1}{\mathfrak{g}}[t^{-1}])^{\mathfrak{h}}$ the $\mathfrak{h}$}-centralizer of $U(t^{-1}{\mathfrak{g}}[t^{-1}])$. It admits a direct sum decomposition
+$U(t^{-1}{\mathfrak{g}}[t^{-1}])^{\mathfrak{h}}=U(t^{-1}{\mathfrak{h}}[t^{-1}])\oplus\mathcal{N},$ (3.8)
+where
+$\mathcal{N}$ $:=$ $U(t^{-1}{\mathfrak{g}}[t^{-1}])^{\mathfrak{h}}\cap U(t^{-1}{\mathfrak{g}}[t^{-1}])t^{-1}\mathfrak{n}_{-}[t^{-1}]$ (3.9)
+$\hskip 2.84526pt=$ $U(t^{-1}{\mathfrak{g}}[t^{-1}])^{\mathfrak{h}}\cap t^{-1}\mathfrak{n}_{+}[t^{-1}]U(t^{-1}{\mathfrak{g}}[t^{-1}]).$ (3.10)
+The projection of $U(t^{-1}{\mathfrak{g}}[t^{-1}])^{\mathfrak{h}}$ onto $U(t^{-1}{\mathfrak{h}}[t^{-1}])$ yields an affine version of the Harish-Chandra homomorphism
+$\mathfrak{f}:U(t^{-1}{\mathfrak{g}}[t^{-1}])^{\mathfrak{h}}\longrightarrow U(t^{-1}{\mathfrak{h}}[t^{-1}]).$
+[…]
+$\mathfrak{f}:{\mathfrak{z}}(\widehat{{\mathfrak{g}}})\longrightarrow\mathcal{W}(^{L}{\mathfrak{g}})$
+is an isomorphism (see [Fr, Theorem 8.1.5] and [Mo2, Theorem 13.1.1]). It is called the Feigin–Frenkel isomorphism.
+The *-structure $\bm{\sigma}$} on $\mathfrak{g}$}, defined in (3.7), extends naturally to a *-structure ${\bm{\sigma}}:U({\mathfrak{g}})\longrightarrow U({\mathfrak{g}})$, which gives rise to a *-structure ${\bm{\sigma}}:U({\mathfrak{g}}[t,t^{-1}])\longrightarrow U({\mathfrak{g}}[t,t^{-1}])$, defined by
+$\hskip 42.67912pt{\bm{\sigma}}(A[r])={\bm{\sigma}}(A)[r]\qquad\mbox{for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{Z}}$}.$
+Evidently, $\bm{\sigma}$} restricts to *-structures on $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ and on $U({\mathfrak{g}}[t])$.
+[…]
+We have the following proposition, which is proved using the same strategy as the proof of [Lu, Proposition 2.4].
+Proposition 3.6.
+The *-structure ${\bm{\sigma}}:U(t^{-1}{\mathfrak{g}}[t^{-1}])\longrightarrow U(t^{-1}{\mathfrak{g}}[t^{-1}])$ fixes $S_{1},\ldots,S_{d}$.
+Proof.
+Let $S\in\{S_{1},\ldots,S_{d}\}$. Since ${\bm{\sigma}}\big(S_{0}\big)=S_{0}$, we have
+${\bm{\sigma}}(S)S_{0}={\bm{\sigma}}(S){\bm{\sigma}}(S_{0})={\bm{\sigma}}(S_{0}S)={\bm{\sigma}}(SS_{0})={\bm{\sigma}}(S_{0}){\bm{\sigma}}(S)=S_{0}{\bm{\sigma}}(S).$
+[…]
+Recall that $S_{1},\ldots,S_{d}\in U(t^{-1}{\mathfrak{g}}[t^{-1}])_{\mathbb{R}}$.
+According to (3.8), $S=H+N$ for some $H\in U(t^{-1}{\mathfrak{h}}[t^{-1}])_{\mathbb{R}}$ and $N\in\mathcal{N}$.
+Clearly, ${\bm{\sigma}}(H)=H$.
+By (3.9) and (3.10), ${\bm{\sigma}}(N)\in\mathcal{N}$.
+Thus, $\mathfrak{f}({\bm{\sigma}}(S))=H$, and hence $\mathfrak{f}({\bm{\sigma}}(S))=\mathfrak{f}(S)$.
+We have ${\bm{\sigma}}(S)=S$ by the injectivity of $\mathfrak{f}:{\mathfrak{z}}(\widehat{{\mathfrak{g}}})\longrightarrow\mathcal{W}(^{L}{\mathfrak{g}})$.
+∎
+Recall the map $\Psi^{\mu}:U(t^{-1}{\mathfrak{g}}[t^{-1}])\longrightarrow U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$ defined by (3.3). For $\mu\in{\mathfrak{g}}^{*}$, let
+$S^{\mu}_{i}(z)=\Psi^{\mu}(S_{i})\qquad\mbox{for $i=1,\ldots,d$.}$ (3.11)
+We define $\overline{\mu}\in{\mathfrak{g}}^{*}$ by $\overline{\mu}(A)=\overline{\mu(A)}$ for $A\in{\mathfrak{g}}$. That is, $\overline{\mu}$ is the complex conjugation of $\mu$. Set
+${\mathfrak{g}}^{*}_{{\bm{\sigma}}}=\left\{\,\mu\in{\mathfrak{g}}^{*}\;\middle|\;\mu\circ{\bm{\sigma}}=\overline{\mu}\,\right\}.$
+Proposition 3.7.
+For any $\mu\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$, the *-structure ${\bm{\sigma}}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}}[t])$ fixes all coefficients of the series $S^{\mu}_{1}(z),\ldots,S^{\mu}_{d}(z)$.
+Proof.
+We may extend ${\bm{\sigma}}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}}[t])$ to a *-structure ${\bm{\sigma}}:U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]\longrightarrow U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$ with ${\bm{\sigma}}(z^{-1})=z^{-1}$.
+For any $\mu\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$, the diagram
+[…]
+commutes. This, together with Proposition 3.6, yields ${\bm{\sigma}}\big(S^{\mu}_{i}(z)\big)=S^{\mu}_{i}(z)$ for $i=1,\ldots,d$. Thus, the proposition follows. ∎
+3.4. Unitarizable modules and diagonalization
+[…]
+Proposition 3.8.
+For any $\mu\in{\mathfrak{g}}^{*}$, the Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ equals the subalgebra of $U({\mathfrak{g}}[t])$ generated by the coefficients of $S^{\mu}_{i}(z)$ for $i=1,\ldots,d$.
+Let $M_{1},\ldots,M_{\ell}$ be unitarizable $\mathfrak{g}$}-modules (with respect to the *-structure $\bm{\sigma}$}).
+By definition, $M_{i}$ possesses a positive definite contravariant Hermitian form $\langle\cdot|\cdot\rangle_{i}$ for each $i=1,\ldots,\ell$.
+Let $\langle\cdot|\cdot\rangle$ denote the tensor form of $\langle\cdot|\cdot\rangle_{1},\ldots,\langle\cdot|\cdot\rangle_{\ell}$.
+It is given by
+$\big\langle v_{1}\otimes\ldots\otimes v_{\ell}\,|\,v^{\prime}_{1}\otimes\ldots\otimes v^{\prime}_{\ell}\big\rangle=\prod_{i=1}^{\ell}\big\langle v_{i}\,|\,v^{\prime}_{i}\big\rangle_{i}$
+for $v_{i},v^{\prime}_{i}\in M_{i}$, $1\leq i\leq\ell$.
+Lemma 3.9.
+Let $\mu\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$ and $\bf z$$\in{\mathbb{R}}^{\ell}$.
+For $i=1,\ldots,d$, let $\bm{b}$} be any coefficient of the series $S^{\mu}_{i}(z)$.
+Then $\bm{b}$} is a Hermitian operator on $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ with respect to $\langle\cdot|\cdot\rangle$.
+Proof.
+Using the contravariance of $\langle\cdot|\cdot\rangle_{1},\ldots,\langle\cdot|\cdot\rangle_{\ell}$ and the hypothesis that $\bf z$$\in{\mathbb{R}}^{\ell}$, we can check directly that
+$\big\langle A_{1}[r_{1}]\ldots A_{k}[r_{k}]v\,|\,v^{\prime}\big\rangle\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\big\langle v\,|\,{\bm{\sigma}}(A_{k})[r_{k}]\ldots{\bm{\sigma}}(A_{1})[r_{1}]v^{\prime}\big\rangle$
+$=$ $\hskip-5.69054pt\big\langle v\,|\,{\bm{\sigma}}\big(A_{1}[r_{1}]\ldots A_{k}[r_{k}]\big)v^{\prime}\big\rangle$
+for $v,v^{\prime}\in\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$, $A_{1},\ldots,A_{k}\in{\mathfrak{g}}$, $r_{1},\ldots,r_{k}\in{\mathbb{Z}}_{+}$, and $k\in{\mathbb{N}}$. This gives
+$\big\langle{\bm{b}}v\,|\,v^{\prime}\big\rangle=\big\langle v\,|\,{\bm{\sigma}}({\bm{b}})v^{\prime}\big\rangle=\big\langle v\,|\,{\bm{b}}v^{\prime}\big\rangle,$
+where the first equality is due to the sesquilinearity of $\langle\cdot|\cdot\rangle$ and the anti-linearity of $\bm{\sigma}$}, and the second equality follows from Proposition 3.7.
+This proves the lemma.
+∎
+Theorem 3.10.
+For any finite-dimensional $\mathcal{B}_{\mathfrak{g}}^{\mu}$-submodule $V$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$, the Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable on $V$ for any $\mu\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$ and $\bf z$$\in{\mathbb{R}}^{\ell}$.
+Proof.
+For $i=1,\ldots,\ell$, let $\bm{b}$} be any coefficient of the series $S^{\mu}_{i}(z)$.
+By Lemma 3.9, $\bm{b}$} is a Hermitian operator on $V$ and hence is diagonalizable on $V$.
+This proves the theorem in view of Proposition 3.8.
+∎
+We expect the following conjecture to be true.
+Conjecture 3.11.
+For any finite-dimensional $\mathcal{B}_{\mathfrak{g}}^{\mu}$-submodule $V$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$, the Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable on $V$ for generic $\mu\in{\mathfrak{g}}^{*}$ and generic $\bf z$$\in{\mathbb{C}}^{\ell}$.
+Let ${\mathfrak{g}}=\mathfrak{n}_{-}\oplus{\mathfrak{h}}\oplus\mathfrak{n}_{+}$ be the triangular decomposition in (2.4).
+For any ${\bf w}:=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$, we define $\mu_{\bf w}\in{\mathfrak{g}}^{*}$ by
+$\mu_{\bf w}(e^{{\mathfrak{x}}}_{ii})=-w_{i},\quad\mbox{for $1\leq i\leq d$;}$ (3.12)
+$\mu_{\bf w}$ vanishes on $\mathfrak{n}_{-}\oplus\mathfrak{n}_{+}$. (3.13)
+That is, $\mu_{\bf w}$ corresponds to a diagonal matrix. The choice of the minus sign will be clear in Section 4.2. Set
+$\mathcal{B}_{\mathfrak{g}}^{\bf w}=\mathcal{B}_{\mathfrak{g}}^{\mu_{\bf w}}.$
+It is evident that ${\mathfrak{h}}\subseteq{\mathfrak{g}}^{\mu_{\bf w}}$.
+Thus for any weight $\gamma\in{\mathfrak{h}}^{*}$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, the weight space $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}$ is a $\mathcal{B}_{\mathfrak{g}}^{\mu}$-submodule of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$.
+Corollary 3.12.
+Let $L_{1},\ldots,L_{\ell}$ be unitarizable highest weight $\mathfrak{g}$}-modules (with respect to $\bm{\sigma}$}), and let $\gamma\in{\mathfrak{h}}^{*}$ be any weight of $\underline{L}$.
+Then $\mathcal{B}_{\mathfrak{g}}^{\bf w}$ is diagonalizable on $\underline{L}({\bf z})_{\gamma}$ for any ${\bf w}\in{\mathbb{R}}^{d}$ and $\bf z$$\in{\mathbb{R}}^{\ell}$.
+Proof.
+Since $\underline{L}({\bf z})_{\gamma}$ is finite-dimensional and $\mu_{\bf w}\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$ for ${\bf w}\in{\mathbb{R}}^{d}$, the corollary follows from Theorem 3.10. ∎
+Theorem 3.3 and Theorem 3.4 are statements for finite-dimensional irreducible $\mathfrak{g}$}-modules.
+Their analogs for unitarizable highest weight $\mathfrak{g}$}-modules (with respect to $\bm{\sigma}$}) are conjectured to be true.
+Conjecture 3.13.
+Let $\underline{L}$ and $\gamma$ be as in Corollary 3.12.
+Then the weight space $\underline{L}({\bf z})_{\gamma}$ is a cyclic $\mathcal{B}_{\mathfrak{g}}^{\bf w}$-module for any ${\bf w}\in{\bm{X}}_{\!d}$ and ${\bf z}\in{\bm{X}}_{\!\ell}$.
+Moreover, $\mathcal{B}_{\mathfrak{g}}^{\bf w}$ is diagonalizable with a simple spectrum on $\underline{L}({\bf z})_{\gamma}$ for generic $\bf w$} and generic $\bf z$}.
+Positive evidence for Conjecture 3.13 (and hence for Conjecture 3.11), in the case where $\mathfrak{g}$} is the general linear Lie algebra, will be given in Corollary 4.17.
+3.5. Frobenius algebras
+A Frobenius algebra is a finite-dimensional commutative associative unital algebra $\mathcal{A}$ which admits a nondegenerate symmetric bilinear form $(\cdot,\cdot)$ such that
+$\hskip 28.45274pt(a_{1}a_{2},a_{3})=(a_{1},a_{2}a_{3})\quad\mbox{for all $a_{1},a_{2},a_{3}\in\mathcal{A}$.}$
+Let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible $\mathfrak{g}$}-modules, and let ${\bf z}\in{\bm{X}}_{\!\ell}$.
+We will give a brief account of the fact that $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a Frobenius algebra for ${\bf w}\in{\bm{X}}_{\!d}$ (cf.
+[MTV1, Lu]).
+For ${\mathfrak{x}}=\mathfrak{a,c,d}$, we define an anti-involution $\varpi:{\mathfrak{g}}\longrightarrow{\mathfrak{g}}$ by
+$\varpi(e^{\mathfrak{x}}_{ij})=e^{\mathfrak{x}}_{ji}\qquad\mbox{for all $i,j$.}$ (3.14)
+The map $\varpi$ is called the Cartan anti-involution on $\mathfrak{g}$}.
+For $i=1,\ldots,\ell$, $V_{i}$ is an irreducible highest weight $\mathfrak{g}$}-module with a dominant integral highest weight.
+Let $v_{i}$ be a highest weight vector of $V_{i}$.
+There is a unique nondegenerate symmetric bilinear form $(\cdot,\cdot)_{i}$ on $V_{i}$ defined by
+$(v_{i},v_{i})_{i}=1\quad\text{and}\quad(Av,v^{\prime})_{i}=(v,\varpi(A)v^{\prime})_{i}$
+for all $v,v^{\prime}\in V_{i}$ and $A\in{\mathfrak{g}}$ ([Sh]).
+The bilinear form $(\cdot,\cdot)_{i}$ is called the Shapovalov form on $V_{i}$.
+It is easy to see that if $v$ and $v^{\prime}$ are weight vectors of $V_{i}$ of distinct weights, then $(v,v^{\prime})_{i}=0$.
+Let $(\cdot,\cdot)$ denote the tensor form of the bilinear forms $(\cdot,\cdot)_{1},\ldots,(\cdot,\cdot)_{\ell}$, which is called the tensor Shapovalov form.
+It is a nondegenerate symmetric bilinear form on $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
+Let ${\bf w}\in{\mathbb{C}}^{d}$.
+By [MTV1, Theorem 9.1] and [Lu, Lemma 2.6], the Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\bf w}$ is symmetric with respect to $(\cdot,\cdot)$, i.e.,
+$\hskip 42.67912pt({\bm{b}}v,v^{\prime})=(v,{\bm{b}}v^{\prime})\quad\mbox{for ${\bm{b}}\in\mathcal{B}_{\mathfrak{g}}^{\bf w}\,$ and $\,v,v^{\prime}\in\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$}.$
+Theorem 3.14.
+For any ${\bf w}\in{\bm{X}}_{\!d}$ and ${\bf z}\in{\bm{X}}_{\!\ell}$, the Bethe algebra $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a Frobenius algebra.
+Moreover, $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\gamma}}$ is a Frobenius algebra for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
+Proof.
+The first statement follows from the above discussion on the tensor Shapovalov form $(\cdot,\cdot)$ on $\underline{V\mkern-4.0mu}\mkern 4.0mu$ in view of Theorem 3.3 and [Lu, Lemma 2.7].
+As $(\cdot,\cdot)$ restricts to a nondegenerate symmetric bilinear form on ${\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\!\gamma}$ for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$, the second statement follows.
+∎
+The following is a consequence of [Lu, Lemma 1.3].
+Theorem 3.15.
+For any ${\bf w}\in{\bm{X}}_{\!d}$ and ${\bf z}\in{\bm{X}}_{\!\ell}$, the following properties hold:
+  1. (i)
+    The Bethe algebra $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a maximal commutative subalgebra of ${\rm End}(\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z}))$ of dimension $\dim\!\left({\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}\right)$.
+  2. (ii)
+    Every eigenspace of $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is one-dimensional, and the set of eigenspaces of $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is in bijective correspondence with the set of maximal ideals of $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$.
+  3. (iii)
+    Every generalized eigenspace of $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a cyclic $\mathcal{B}_{\mathfrak{g}}^{\bf w}$-module.
+Remark 3.16.
+Each statement of Theorem 3.15 holds if $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ is replaced with ${\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\gamma}$ for any weight $\mu$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
+4. A Duality of Bethe algebras and an application
+In this section, we establish a duality of Bethe algebras for the general linear Lie (super)algebras ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$.
+We apply it to study the action of the Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$ on a tensor product of infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules and obtain evidence for Conjecture 3.13 if $q$ and $n$ are set to 0.
 4.1. Column determinants and Berezinians
-Let $\mathcal{A}$ be an associative unital superalgebra over ${\mathbb{C}}$.
+Let us review the notions of column determinants and Berezinians.
+Let $\mathcal{A}$ be an associative unital superalgebra.
 The parity of a homogeneous element $a\in\mathcal{A}$ is denoted by $|a|$, which lies in ${\mathbb{Z}}_{2}$.
@@
 For any $k\times k$ matrix $A=\big[a_{i,j}\big]_{i,j=1,\ldots,k}$ over $\mathcal{A}$, the column determinant of $A$ is defined to be
-${\rm cdet}\,A=\sum_{\sigma\in\mathfrak{S}_{k}}(-1)^{l(\sigma)}\,a_{\sigma(1),1}\ldots a_{\sigma(k),k}.$
+${\rm cdet}(A)=\sum_{\sigma\in\mathfrak{S}_{k}}(-1)^{l(\sigma)}\,a_{\sigma(1),1}\ldots a_{\sigma(k),k}.$
 Here $\mathfrak{S}_{k}$ denotes the symmetric group on $\{1,\ldots,k\}$ and $l(\sigma)$ denotes the length of $\sigma$.
-For any nonempty subset $P=\{i_{1}<\ldots<i_{\ell}\}$ of $\{1,\ldots,k\}$, the matrix $A_{P}:=\big{[}a_{i,j}\big{]}_{i,j\in P}$ is called a standard submatrix of $A$.
 Assume that $A$ has a two-sided inverse $A^{-1}=\big[\widetilde{a}_{i,j}\big]$.
 For all $i,j=1,\ldots,k$, the $(i,j)$th quasideterminant of $A$ is defined to be $|A|_{ij}:=\widetilde{a}_{j,i}^{-1}$ provided that $\widetilde{a}_{j,i}$ has an inverse in $\mathcal{A}$.
-According to [GGRW, Proposition 1.2.6], if $k\geq 2$, then
-$|A|_{ij}=a_{i,j}-r_{i}^{j}(A^{ij})^{-1}c_{j}^{i},$
-where $A^{ij}$ is the $(k-1)\times(k-1)$ submatrix of $A$ obtained by deleting the $i$th row and the $j$th column of $A$ and is assumed to be invertible, $r_{i}^{j}$ is the $i$th row of $A$ with $a_{i,j}$ removed, and $c_{j}^{i}$ is the $j$th column of $A$ with $a_{i,j}$ removed.
 Following the notation of [GGRW], it is convenient to write
@@
 For $i=1,\ldots,k$, we define
-$d_{i}(A)=\begin{vmatrix}a_{1,1}&\ldots&a_{1,i}\\ \ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,i}$}\end{vmatrix},$
+$d_{i}(A)=\begin{vmatrix}a_{1,1}&\ldots&a_{1,i}\\ \ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,i}$}\end{vmatrix}$,\
 which are called the principal quasiminors of $A$.
-A $k\times k$ matrix $A$ over $\mathcal{A}$ is called sufficiently invertible if every principal quasiminor of $A$ is well defined, and $A$ is called amply invertible if each of its standard submatrices is sufficiently invertible.
-Let ${{\bf s}}=(s_{1},\ldots,s_{m+n})$ be a sequence of 0’s and 1’s such that exactly $m$ of the $s_{i}$’s are 0 and the others are 1.
-We call such a sequence a $0^{m}1^{n}$-sequence.
-Every $0^{m}1^{n}$-sequence can be written in the form $(0^{m_{1}},1^{n_{1}},\ldots,0^{m_{r}},1^{n_{r}}),$ where the sequence begins with $m_{1}$ copies of 00’s, followed by $n_{1}$ copies of $1$’s, and so on.
+Let $A=\big[a_{i,j}\big]_{i,j=1,\ldots,k}$ be a $k\times k$ matrix over $\mathcal{A}$.
+For any nonempty subset $P=\{i_{1}<\ldots<i_{p}\}$ of $\{1,\ldots,k\}$, the matrix $A_{P}:=\big[a_{i,j}\big]_{i,j\in P}$ is called a standard submatrix of $A$.
+We say that $A$ is sufficiently invertible if every principal quasiminor of $A$ is well defined, and that $A$ is amply invertible if each of its standard submatrices is sufficiently invertible.
+A sequence $(s_{1},\ldots,s_{m+n})$ of 0’s and 1’s is called a $0^{m}1^{n}$-sequence if exactly $m$ of the $s_{i}$’s are 0 and the others are 1.
+Every $0^{m}1^{n}$-sequence can be written in the form $(0^{m_{1}},1^{n_{1}},\ldots,0^{m_{r}},1^{n_{r}}),$ where the sequence begins with $m_{1}$ copies of 0’s, followed by $n_{1}$ copies of $1$’s, and so on.
 The set of all $0^{m}1^{n}$-sequences is denoted by $\mathcal{S}_{{m|n}}$.
-Let ${\bf s}\in\mathcal{S}_{{m|n}}$.
-For any $\sigma\in\mathfrak{S}_{m+n}$ and any $(m+n)\times(m+n)$ matrix $A=[a_{i,j}]$ over $\mathcal{A}$, we define ${\bf s}^{\sigma}=\left(s_{\sigma^{-1}(1)},s_{\sigma^{-1}(2)},\ldots,s_{\sigma^{-1}(m+n)}\right)$ and $A^{\sigma}=\big{[}a_{\sigma^{-1}(i),\sigma^{-1}(j)}\big{]}$.
-We say that $A$ is of type ${\bf s}$ if $a_{i,j}$ is a homogeneous element of parity $|a_{i,j}|=\bar{s}_{i}+\bar{s}_{j}$ for any $i,j=1,\ldots,m+n$.
-For any $(m+n)\times(m+n)$ sufficiently invertible matrix $A$ of type ${\bf s}$ over $\mathcal{A}$, the Berezinian of type ${\bf s}$ of $A$ is defined to be
-${\rm Ber}^{\bf s}\,A=d_{1}(A)^{\hat{s}_{1}}\ldots d_{m+n}(A)^{\hat{s}_{m+n}},$
-where $\hat{s}_{i}:=(-1)^{s_{i}}$ for $1\leq i\leq m+n$ (see [HM, (3.3)]). We refer the reader to [Ber, Na, MR] for earlier definitions of Berezinians. We have the following proposition.
+Fix ${\bf s}:=(s_{1},\ldots,s_{m+n})\in\mathcal{S}_{{m|n}}$.
+For any $\sigma\in\mathfrak{S}_{m+n}$ and any $(m+n)\times(m+n)$ matrix $A=[a_{i,j}]_{i=1,\ldots,m+n}$ over $\mathcal{A}$, we define ${\bf s}^{\sigma}=\left(s_{\sigma^{-1}(1)},s_{\sigma^{-1}(2)},\ldots,s_{\sigma^{-1}(m+n)}\right)$ and $A^{\sigma}=\big[a_{\sigma^{-1}(i),\sigma^{-1}(j)}\big]_{i=1,\ldots,m+n}$.
+We say that $A$ is of type $\bf s$} if $a_{i,j}$ is homogeneous of parity $|a_{i,j}|=\bar{s}_{i}+\bar{s}_{j}$ for any $i,j=1,\ldots,m+n$.
+For any sufficiently invertible matrix $A$ of type $\bf s$} over $\mathcal{A}$, the Berezinian of type $\bf s$} of $A$ is defined to be
+${\rm Ber}^{\bf s}(A)=d_{1}(A)^{\hat{s}_{1}}\ldots d_{m+n}(A)^{\hat{s}_{m+n}},$
+where $\hat{s}_{i}:=(-1)^{s_{i}}$ for $i=1,\ldots,m+n$ (see [HM, (3.3)]). We refer the reader to [Ber, Na, MR] for earlier definitions of Berezinians. The following proposition is useful.
 Proposition 4.1 (cf. [HM, Proposition 3.5]).
-Let $A$ be an $(m+n)\times(m+n)$ amply invertible matrix of type ${\bf s}$ over $\mathcal{A}$. Fix $k\in\{1,\ldots,m+n-1\}$. We write
+Let $A$ be an amply invertible matrix of type $\bf s$} over $\mathcal{A}$. Fix $k\in\{1,\ldots,m+n-1\}$. We write
 $A=\begin{bmatrix}W&X\\ Y&Z\end{bmatrix}$, (4.1)
@@
 Moreover,
-${\rm Ber}^{\bf s}\,A={\rm Ber}^{{\bf s}^{\prime}}W\cdot{\rm Ber}^{{\bf s}^{\prime\prime}}\big{(}Z-YW^{-1}X\big{)}.$
-An $(m+n)\times(m+n)$ matrix $A=\big{[}a_{i,j}\big{]}$ over $\mathcal{A}$ is called a Manin matrix of type ${\bf s}$ if $A$ is a matrix of type ${\bf s}$ satisfying the following relations
-$[a_{i,j},a_{k,l}]=(-1)^{s_{i}s_{j}+s_{i}s_{k}+s_{j}s_{k}}[a_{k,j},a_{i,l}]\qquad\text{for all}\quad i,j,k,l=1,\ldots,m+n,$
-where $[a,b]:=ab-(-1)^{|a||b|}ba$ for any homogeneous elements $a,b\in\mathcal{A}$. We recall some basic facts about Manin matrices.
+${\rm Ber}^{\bf s}(A)={\rm Ber}^{{\bf s}^{\prime}}(W)\cdot{\rm Ber}^{{\bf s}^{\prime\prime}}\big(Z-YW^{-1}X\big).$
+A matrix $A=[a_{i,j}]_{i=1,\ldots,m+n}$ over $\mathcal{A}$ is called a Manin matrix of type $\bf s$} if $A$ is a matrix of type $\bf s$} satisfying the following relations
+$a_{i,j},a_{k,l}]=(-1)^{s_{i}s_{j}+s_{i}s_{k}+s_{j}s_{k}}[a_{k,j},a_{i,l}]\qquad\text{for all}\quad i,j,k,l=1,\ldots,m+n,$
+where $a,b]:=ab-(-1)^{|a||b|}ba$ for any homogeneous elements $a,b\in\mathcal{A}$. We recall some basic facts about Manin matrices.
 Proposition 4.2 (cf. [HM, Section 3]).
-Let $A$ be an $(m+n)\times(m+n)$ Manin matrix of type ${\bf s}$ over $\mathcal{A}$. Then
+Let $A$ be a Manin matrix of type $\bf s$} over $\mathcal{A}$. Then
   1. (i)
-    $A^{-1}$ is an $(m+n)\times(m+n)$ Manin matrix of type ${\bf s}$ over $\mathcal{A}$.
+    If $P=\{i_{1}<\ldots<i_{p}\}$ is a nonempty subset of $\{1,\ldots,m+n\}$, then the standard submatrix $A_{P}$ is a Manin matrix of type $(s_{i_{1}},\ldots s_{i_{p}})$.
   2. (ii)
-    If $P=\{i_{1}<\ldots<i_{\ell}\}$ is a nonempty subset of $\{1,\ldots,m+n\}$, then the standard submatrix $A_{P}$ of $A$ is a Manin matrix of type ${{\bf s}}_{P}:=(s_{i_{1}},\ldots s_{i_{\ell}})$.
-  3. (iii)
     For any $\sigma\in\mathfrak{S}_{m+n}$, $A^{\sigma}$ is a Manin matrix of type ${\bf s}^{\sigma}$.
-The following proposition asserts that the Berezinian of a Manin matrix is invariant under the action of a permutation.
+The following proposition asserts that the Berezinian of a Manin matrix is invariant under any permutation.
 Proposition 4.3 ([HM, Proposition 3.6]).
-Let $A$ be an $(m+n)\times(m+n)$ amply invertible Manin matrix of type ${\bf s}$ over $\mathcal{A}$. Then
-$\hskip 28.45274pt{\rm Ber}^{{\bf s}^{\sigma}}A^{\sigma}={\rm Ber}^{\bf s}\,A\qquad\mbox{for any $\sigma\in\mathfrak{S}_{m+n}$.}$
-We may decompose ${\rm Ber}^{\bf s}\,A$ in Proposition 3.1 in a different way if $A$ is a Manin matrix.
+Let $A$ be an amply invertible Manin matrix of type $\bf s$} over $\mathcal{A}$. Then
+$\hskip 28.45274pt{\rm Ber}^{{\bf s}^{\sigma}}(A^{\sigma})={\rm Ber}^{\bf s}(A)\qquad\mbox{for any $\sigma\in\mathfrak{S}_{m+n}$.}$
+We may decompose ${\rm Ber}^{\bf s}(A)$ in Proposition 4.1 in a different way if $A$ is a Manin matrix.
 Proposition 4.4.
-Let $A$ be an $(m+n)\times(m+n)$ amply invertible Manin matrix of type ${\bf s}$ over $\mathcal{A}$.
+Let $A$ be an amply invertible Manin matrix of type $\bf s$} over $\mathcal{A}$.
 Write $A$ as in (4.1).
-Then $W-XZ^{-1}Y$ and $Z$ are sufficiently invertible Manin matrices of types ${\bf s}^{\prime}:=(s_{1},\ldots,s_{k})$ and ${\bf s}^{\prime\prime}:=(s_{k+1},\ldots,s_{m+n})$, respectively.
+Then $W-XZ^{-1}Y$ and $Z$ are sufficiently invertible matrices of types ${\bf s}^{\prime}:=(s_{1},\ldots,s_{k})$ and ${\bf s}^{\prime\prime}:=(s_{k+1},\ldots,s_{m+n})$, respectively.
 Moreover,
-${\rm Ber}^{\bf s}\,A={\rm Ber}^{{\bf s}^{\prime\prime}}Z\cdot{\rm Ber}^{{\bf s}^{\prime}}\big{(}W-XZ^{-1}Y\big{)}.$ (3.2)
+${\rm Ber}^{\bf s}(A)={\rm Ber}^{{\bf s}^{\prime\prime}}(Z)\cdot{\rm Ber}^{{\bf s}^{\prime}}\big(W-XZ^{-1}Y\big).$ (4.2)
 Proof.
-The first assertion is clear from Proposition 3.2.
-Let $\sigma\in\mathfrak{S}_{m+n}$ be the permutation defined by $\sigma(k+i)=i$ for $i=1,\ldots,m+n-k$ and $\sigma(j)=m+n-k+j$ for $j=1,\ldots,k$.
-Then
+The first assertion is clear. Let $\sigma\in\mathfrak{S}_{m+n}$ be the permutation defined by $\sigma(k+i)=i$ for $i=1,\ldots,m+n-k$ and $\sigma(j)=m+n-k+j$ for $j=1,\ldots,k$. Then
 $A^{\sigma}=\begin{bmatrix}Z&Y\\ X&W\end{bmatrix}$.
 The equality (4.2) is now a direct consequence of Proposition 4.1 and Proposition 4.3. ∎
-Proposition 3.4 will be used in Section 4.2. We also have the following proposition.
+We also have the following.
 Proposition 4.5 ([CFR, Lemma 8]).
-Let ${\bf s}=(0^{m})$. For any $m\times m$ sufficiently invertible Manin matrix $A$ of type ${\bf s}$ over $\mathcal{A}$, we have ${\rm Ber}^{\bf s}\,A={\rm cdet}\,\,A$.
-3.2. Pseudo-differential operators
-Let $\mathcal{A}$ be an associative unital superalgebra over ${\mathbb{C}}$ and $z$ an even variable.
-We denote by $\mathcal{A}[\kern-1.1pt[z]\kern-1.1pt]$ and $\mathcal{A}(\!(z)\!)$ the superalgebras of formal power series and Laurent series in $z$ with coefficients in $\mathcal{A}$, respectively.
-Let $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ denote the set of all formal series of the form
-$\sum_{\ell=-\infty}^{s}\sum_{k=-\infty}^{r}a_{k\ell}z^{k}\partial_{z}^{\ell},$
-where $r,s\in{\mathbb{Z}}$ and $a_{k\ell}\in\mathcal{A}$. We may endow $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ with a superalgebra structure using the rules:
-$\partial_{z}\partial_{z}^{-1}=\partial_{z}^{-1}\partial_{z}=1,\quad\partial_{z}^{k}z^{\ell}=\sum_{i=0}^{\infty}\binom{k}{i}\binom{\ell}{i}i!\,z^{\ell-i}\,\partial_{z}^{k-i},\quad\mbox{for $k,\ell\in{\mathbb{Z}}$.}$ (3.3)
-Here, e.g., ${\binom{k}{i}:=\frac{k(k-1)\ldots(k-i+1)}{i!}}$.
-The superalgebra $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ is called the superalgebra of pseudo-differential operators over $\mathcal{A}$.
-Let $\omega$ be an anti-involution on $\mathcal{A}$.
-In particular, $\omega^{2}=1$ and $\omega(ab)=\omega(b)\omega(a)$ for $a,b\in\mathcal{A}$.
-We readily see that $\omega$ extends to a ${\mathbb{C}}$-linear automorphism $\omega:\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ defined by
-$\omega(az^{k}\partial_{z}^{\ell})=\omega(a)z^{\ell}\partial_{z}^{k}$ (3.4)
-for $a\in\mathcal{A}$ and $k,\ell\in{\mathbb{Z}}$.
-Proposition 3.6.
-The map $\omega$ is an anti-involution on $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$.
-Proof.
-Clearly, $\omega^{2}=1$. It remains to show that $\omega$ is an anti-homomorphism. For $a,b\in\mathcal{A}$ and $k,\ell,r,s\in{\mathbb{Z}}$, we have
-$\omega((az^{k}\partial_{z}^{\ell})(bz^{r}\partial_{z}^{s}))$ $=$ $\omega(abz^{k}\partial_{z}^{\ell}z^{r}\partial_{z}^{s})$
-$=$ $\sum_{i=0}^{\infty}\binom{\ell}{i}\binom{r}{i}i!\,\omega(abz^{k+r-i}\partial_{z}^{\ell+s-i})$
-$=$ $\sum_{i=0}^{\infty}\binom{\ell}{i}\binom{r}{i}i!\,\omega(ab)z^{\ell+s-i}\partial_{z}^{k+r-i}$
-$=$ $\omega(b)\omega(a)z^{s}\partial_{z}^{r}z^{\ell}\partial_{z}^{k}$
-$=$ $\left(\omega(b)z^{s}\partial_{z}^{r}\right)\left(\omega(a)z^{\ell}\partial_{z}^{k}\right)$
-$=$ $\omega(bz^{r}\partial_{z}^{s})\,\omega(az^{k}\partial_{z}^{\ell}),$
-and the proposition follows. ∎
-3.3. The Bethe algebra for ${\mathfrak{gl}}_{d}$ and Frobenius algebras
-Let
-${\bf w}=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{\ell}.$
-Define $\mu_{\bf w}\in{\mathfrak{gl}}_{d}^{*}$ by
-$\hskip 28.45274pt\mu_{\bf w}(e_{ab})=-\delta_{a,b}w_{a}\qquad\mbox{for $1\leq a\leq d$.}$
-The choice of the minus sign will be clear below.
-Recall the map $\Psi^{\mu_{\bf w}}$ defined in (2.3). We write
-$\Psi^{\bf w}=\Psi^{\mu_{\bf w}}.$
+Let $A$ be a sufficiently invertible Manin matrix of type ${\bf s}_{0}:=(0^{m})$ over $\mathcal{A}$. Then ${\rm Ber}^{{\bf s}_{0}}(A)={\rm cdet}(A)$.
+4.2. The Bethe algebras for ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$
 Let $t$ be an even variable and
 $\tau=-\partial_{t}.$
-Here $t$ and $\partial_{t}$ satisfy the rules similar to (3.3). The map $\Psi^{\bf w}$ extends to an algebra homomorphism
+Here $t$ and $\partial_{t}$ satisfy the rules similar to (2.8).
+Let ${\bf w}=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$, and recall $\mu_{\bf w}\in{\mathfrak{gl}}_{d}^{*}$ defined by (3.12) and (3.13).
+Write
+$\Psi^{\bf w}=\Psi^{\mu_{\bf w}}.$
+The map $\Psi^{\bf w}$ extends to an algebra homomorphism
 $\Psi^{\bf w}:U(t^{-1}{\mathfrak{gl}}_{d}[t^{-1}])[\tau]\longrightarrow U({\mathfrak{gl}}_{d}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt][\partial_{z}]$
-such that $\Psi^{\bf w}(\tau)=\partial_{z}$. Let
+such that $\Psi^{\bf w}(\tau)=\partial_{z}$.
+Let
 $\mathcal{T}_{d}=\Big[\delta_{a,b}\tau+e_{ab}[-1]\Big]_{a,b=1,\ldots,d},$
-which is a $d\times d$ Manin matrix. We have
-${\rm cdet}\,(\mathcal{T}_{d})=\tau^{d}+\sum_{i=1}^{d}S_{i}\tau^{i-1}.$ (3.5)
-By [CM, Theorem 3.1] (see also [CT]), $S_{1},\ldots,S_{d}$ form a complete set of Segal–Sugawara vectors for the Feigin–Frenkel center ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$.
-That is, the set $\{{\rm T}^{r}S_{i}\,|\,i=1,\ldots,d,\,r\in{\mathbb{Z}}_{+}\}$ is algebraically independent, and
-${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})={\mathbb{C}}[{\rm T}^{r}S_{i}\,|\,i=1,\ldots,d,\,r\in{\mathbb{Z}}_{+}].$
-Now consider the $d\times d$ Manin matrix
-${\mathcal{L}}_{d}^{\bf w}:=\Big{[}\Psi^{\bf w}\big{(}\delta_{a,b}\tau+e_{ab}[-1]\big{)}\Big{]}_{a,b=1,\ldots,d}=\Big{[}\delta_{a,b}(\partial_{z}-w_{a})-e_{ab}(z)\Big{]}_{a,b=1,\ldots,d}.$
+which is a Manin matrix of type $(0^{d})$.
+The vectors $S_{1},\ldots,S_{d}\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$, mentioned in (3.2), are the coefficients of ${\rm cdet}(\mathcal{T}_{d})\in U(t^{-1}{\mathfrak{gl}}_{d}[t^{-1}])[\tau]$; more precisely,
+${\rm cdet}(\mathcal{T}_{d})=\tau^{d}+\sum_{i=1}^{d}S_{i}\tau^{i-1}$ (4.3)
+(see [CM, Theorem 3.1]). Now consider the Manin matrix
+${\mathcal{L}}_{d}^{\bf w}:=\Big[\Psi^{\bf w}\big(\delta_{a,b}\tau+e_{ab}[-1]\big)\Big]_{a,b=1,\ldots,d}=\Big[\delta_{a,b}(\partial_{z}-w_{a})-e_{ab}(z)\Big]_{a,b=1,\ldots,d}.$ (4.4)
 Then in view of (4.3),
 ${\rm cdet}({\mathcal{L}}_{d}^{\bf w})=\partial_{z}^{d}+\sum_{i=1}^{d}S_{i}^{{\bf w}}(z)\partial_{z}^{i-1},$ (4.5)
-where $S_{i}^{{\bf w}}(z):=\Psi^{\bf w}(S_{i})$.
-The following is well-known and matches the definition of the Bethe algebra in [MTV1].
-Proposition 3.7.
-The Bethe algebra $\mathcal{B}_{d}^{\bf w}$ is generated by the coefficients of $S_{i}^{{\bf w}}(z)$ for $i=1,\ldots,d$.
-A Frobenius algebra $\mathcal{A}$ is a finite-dimensional commutative associative unital algebra which admits a nondegenerate symmetric bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) such that
-$\hskip 28.45274pt(a_{1}a_{2},a_{3})=(a_{1},a_{2}a_{3})\quad\mbox{for all $a_{1},a_{2},a_{3}\in\mathcal{A}$.}$
-Recall the notations (2.1) and (2.2).
-Let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible ${\mathfrak{gl}}_{d}$-modules, and let ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$.
-As ${\mathfrak{h}}_{d}\subseteq{\mathfrak{gl}}_{d}^{\mu_{\bf w}}$, the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ acts on any weight space of $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$.
-We will explain why $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a Frobenius algebra for ${\bf w}\in{\boldsymbol{X}}_{\!d}$.
-Define an anti-involution $\boldsymbol{\theta}:{\mathfrak{gl}}_{d}\longrightarrow{\mathfrak{gl}}_{d}$ by
-$\boldsymbol{\theta}(e_{ab})=e_{ba}\qquad\mbox{for $1\leq a,b\leq d$.}$ (3.7)
-The map $\boldsymbol{\theta}$ is called the Cartan anti-involution on ${\mathfrak{gl}}_{d}$.
-For $i=1,\ldots,\ell$, $V_{i}$ is an irreducible highest weight ${\mathfrak{gl}}_{d}$-module with a dominant integral highest weight.
-Let $v_{i}$ be a highest weight vector of $V_{i}$.
-There is a unique nondegenerate symmetric bilinear form $(\cdot,\cdot)_{i}$ on $V_{i}$ defined by
-$(v_{i},v_{i})_{i}=1\quad\text{and}\quad(Av,v^{\prime})_{i}=(v,\boldsymbol{\theta}(A)v^{\prime})_{i}$
-for all $v,v^{\prime}\in V_{i}$ and $A\in{\mathfrak{gl}}_{d}$ ([Sh]).
-The bilinear form $(\cdot,\cdot)_{i}$ is called the Shapovalov form on $V_{i}$.
-It is easy to see that if $v$ and $v^{\prime}$ are weight vectors of $V_{i}$ of distinct weights, then $(v,v^{\prime})_{i}=0$.
-Let (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denote the tensor form of the bilinear forms $(\cdot,\cdot)_{1},\ldots,(\cdot,\cdot)_{\ell}$, which is called the tensor Shapovalov form.
-It is a nondegenerate symmetric bilinear form on $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
-Let ${\bf w}\in{\mathbb{C}}^{d}$.
-By [MTV1, Theorem 9.1] and [Lu, Lemma 2.6], the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ is symmetric with respect to (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ), i.e.,
-$\hskip 42.67912pt({\boldsymbol{b}}v,v^{\prime})=(v,{\boldsymbol{b}}v^{\prime})\quad\mbox{for ${\boldsymbol{b}}\in\mathcal{B}_{d}^{\bf w}$ and $v,v^{\prime}\in\underline{V\mkern-4.0mu}\mkern 4.0mu$}.$
-Theorem 3.8.
-For any ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, the Bethe algebra $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a Frobenius algebra.
-Moreover, $(\mathcal{B}_{d}^{\bf w})_{{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\gamma}}$ is a Frobenius algebra for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
-Proof.
-The first statement follows from the above discussion on the tensor Shapovalov form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on $\underline{V\mkern-4.0mu}\mkern 4.0mu$ in view of Theorem 2.1 and [Lu, Lemma 2.7].
-As (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) restricts to a nondegenerate symmetric bilinear form on ${\underline{V\mkern-4.0mu}\mkern 4.0mu}_{\!\gamma}$ for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$, the second statement follows.
-∎
-The following is a consequence of [Lu, Lemma 1.3].
-Theorem 3.9.
-For any ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, the following properties hold:
-  1. (i)
-    The Bethe algebra $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a maximal commutative subalgebra of ${\rm End}(\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z}))$ of dimension $\dim\!\left({\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}\right)$.
-  2. (ii)
-    Every eigenspace of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is one-dimensional, and the set of eigenspaces of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is in bijective correspondence with the set of maximal ideals of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$.
-  3. (iii)
-    Every generalized eigenspace of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a cyclic $\mathcal{B}_{d}^{\bf w}$-module.
-Remark 3.10.
-Each statement of Theorem 3.9 holds if $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ is replaced with ${\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\gamma}$ for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
-3.4. The Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$
+where $S_{i}^{{\bf w}}(z):=\Psi^{\bf w}(S_{i})$. Write
+$\mathcal{B}_{d}^{\bf w}=\mathcal{B}_{{\mathfrak{gl}}_{d}}^{\bf w}.$
+Recall from Proposition 3.8 that $\mathcal{B}_{d}^{\bf w}$ is equal to the subalgebra of $U({\mathfrak{gl}}_{d}[t])$ generated by the coefficients of $S_{i}^{{\bf w}}(z)$ for $i=1,\ldots,d$.
 Fix ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. Let
 $\mathcal{T}_{p+m|q+n}=\Big[\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}[-1]\Big]_{i,j\in\mathbb{I}}.$
-It is straightforward to verify that $\mathcal{T}_{p+m|q+n}$ is an amply invertible $(p+q+m+n)\times(p+q+m+n)$ Manin matrix of type ${\bf s}$ over $U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])(\!(\tau^{-1})\!)$.
-The verification is analogous to [MR, Lemma 3.1].
-The following is an immediate consequence of Proposition 3.3.
-Proposition 3.11.
-For each $\sigma\in\mathfrak{S}_{p+q+m+n}$,
-${\rm Ber}\,^{{\bf s}^{\sigma}}\!\left(\mathcal{T}_{p+m|q+n}^{\sigma}\right)={\rm Ber}^{\bf s}\,\!\left(\mathcal{T}_{p+m|q+n}\right).$
-We have the expansion
+It is straightforward to verify that $\mathcal{T}_{p+m|q+n}$ is an amply invertible Manin matrix of type $\bf s$} over $U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])(\!(\tau^{-1})\!)$ (cf.
+[MR, Lemma 3.1]).
+We have the following expansion:
 ${\rm Ber}^{\bf s}\!\left(\mathcal{T}_{p+m|q+n}\right)=\sum_{i=-\infty}^{p+m-q-n}b_{i}\tau^{i},$ (4.6)
-for some $b_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$. According to [MR, Corollary 3.3] (see also [ChL2, Proposition 3.3]),
-$b_{i}\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$
-for all $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$.
-Now let
-${\bf z}=(z_{1},\ldots,z_{p+q+m+n})\in{\mathbb{C}}^{p+q+m+n}.$
-Define $\mu_{\bf z}\in{\mathfrak{gl}}_{p+m|q+n}^{*}$ by
+for some $b_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$.
+Let $\widehat{{\mathfrak{z}}}_{p+m|q+n}$ be the subalgebra of $U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$ generated by
+$\{{\rm T}^{r}b_{i}\,|\,i\leq p+m-q-n,\,i\in{\mathbb{Z}},\,r\in{\mathbb{Z}}_{+}\},$
+where $\rm T$} is defined by (3.1).
+According to [MR, Corollary 3.3] (see also [ChL2, Proposition 3.3]), $b_{i}\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$ for all $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$.
+Since ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$ is $\rm T$}-invariant, we see that $\widehat{{\mathfrak{z}}}_{p+m|q+n}$ is a commutative subalgebra of ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$.
+There is a conjecture asserting that ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})=\widehat{{\mathfrak{z}}}_{p+m|q+n}$ (see [MR]).
+The conjecture holds if $q=n=0$ (see [CM, Theorem 3.1] and also Theorem 3.1).
+It is also true that ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{1|1})=\widehat{{\mathfrak{z}}}_{1|1}$ ([MM]) and ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{2|1})=\widehat{{\mathfrak{z}}}_{2|1}$ ([AN]).
+However, it is still unknown whether the conjecture is valid in general.
+Now let ${\bf z}=(z_{1},\ldots,z_{p+q+m+n})\in{\mathbb{C}}^{p+q+m+n}$. Define $\mu_{\bf z}\in{\mathfrak{gl}}_{p+m|q+n}^{*}$ by
 $\hskip 28.45274pt\mu_{\bf z}(E^{i}_{j})=(-1)^{|i|+1}\delta_{i,j}z_{i}\qquad\mbox{for $i,j\in\mathbb{I}$.}$
@@
 $\Psi^{\bf z}\left(\sum_{i=-\infty}^{s}a_{i}\tau^{i}\right)=\sum_{i=-\infty}^{s}\Psi^{\bf z}(a_{i})\partial_{z}^{i}$
-for $a_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$ and $s\in{\mathbb{Z}}$.
-Now consider the $(p+q+m+n)\times(p+q+m+n)$ matrix
-${\mathcal{L}}_{{p+m|q+n}}^{\bf z}:=\Big{[}\Psi^{\bf z}\big{(}\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}[-1]\big{)}\Big{]}_{i,j\in\mathbb{I}}=\Big{[}\delta_{i,j}(\partial_{z}-z_{i})-(-1)^{|i|}E^{i}_{j}(z)\Big{]}_{i,j\in\mathbb{I}}.$
-Evidently, ${\mathcal{L}}_{{p+m|q+n}}^{\bf z}$ is an amply invertible Manin matrix of type ${\bf s}$ over the superalgebra $U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)$ of pseudo-differential operators over $U({\mathfrak{gl}}_{p+m|q+n}[t])$.
-Just like Proposition 3.11, we have the following.
+for $a_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$ and $s\in{\mathbb{Z}}$. Consider the matrix
+${\mathcal{L}}_{{p+m|q+n}}^{\bf z}:=\Big[\Psi^{\bf z}\big(\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}[-1]\big)\Big]_{i,j\in\mathbb{I}}=\Big[\delta_{i,j}(\partial_{z}-z_{i})-(-1)^{|i|}E^{i}_{j}(z)\Big]_{i,j\in\mathbb{I}}.$ (4.7)
+It is an amply invertible Manin matrix of type $\bf s$} over $U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)$. The following is an immediate consequence of Proposition 4.3.
 Proposition 4.6.
 For each $\sigma\in\mathfrak{S}_{p+q+m+n}$,
-${\rm Ber}\,^{{\bf s}^{\sigma}}\big{(}\big{(}{\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big{)}^{\sigma}\big{)}={\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right).$
+${\rm Ber}^{{\bf s}^{\sigma}}\big(\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)^{\sigma}\big)={\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big).$
 By (4.6), we have
-${\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)=\sum_{i=-\infty}^{p+m-q-n}b_{i}^{{\bf z}}(z)\partial_{z}^{i},$ (3.9)
+${\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)=\sum_{i=-\infty}^{p+m-q-n}b_{i}^{{\bf z}}(z)\partial_{z}^{i},$ (4.8)
 where $b_{i}^{{\bf z}}(z):=\Psi^{\bf z}(b_{i})$.
-Let ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ be the subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$ generated by the coefficients of the series $b_{i}^{{\bf z}}(z)$, for $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$.
-It is called the Bethe algebra of ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}$.
-If we take $q=n=0$, we recover the Bethe algebra of ${\mathfrak{gl}}_{p+m}$ by virtue of Proposition 3.5.
+Let ${\mathcal{B}_{p+m|q+n}^{\bf z}}:={\mathcal{B}_{{\mathfrak{gl}}_{p+m|q+n}}^{\bf z}}$ be the subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$ generated by the coefficients of the series $b_{i}^{{\bf z}}(z)$, for $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$.
+It is called the Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$ with respect to $\bf z$}.
+We remark that $\mathcal{B}_{p+m|q+n}^{\bf z}$} coincides with the subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$ generated by the coefficients of $\Psi^{\bf z}(S)$ for $S\in\widehat{{\mathfrak{z}}}_{p+m|q+n}$ (as $\Psi^{\bf z}({\rm T}(S))=\frac{d}{dz}(\Psi^{\bf z}(S))$).
+If we take $q=n=0$, we recover the Bethe algebra for ${\mathfrak{gl}}_{p+m}$ by virtue of Proposition 4.5.
 According to [MR, Corallary 3.7], the algebra $\mathcal{B}_{p+m|q+n}^{\bf z}$} is a commutative subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$.
@@
 For any $\mathcal{B}_{p+m|q+n}^{\bf z}$}-module $V$, let $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{V}$ denote the image of the Bethe algebra $\mathcal{B}_{p+m|q+n}^{\bf z}$} in ${\rm End}(V)$.
-We call $\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}_{V}$ the Bethe algebra of $V$ with respect to ${\bf z}$.
-4. Duality for Bethe algebras
-This section is devoted to establishing a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ between the Bethe algebras $\mathcal{B}_{d}^{\bf w}\subseteq U({\mathfrak{gl}}_{d}[t])$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}\subseteq U({\mathfrak{gl}}_{p+m|q+n}[t])$ for any ${\bf w}\in{\mathbb{C}}^{d}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$.
-4.1. The actions of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$
-Let $\mathcal{D}$ be the associative unital superalgebra generated by
-$x^{a}_{i},\,\eta^{a}_{j},\,y^{a}_{r},\,\zeta^{a}_{s}$
-as well as their derivatives
+We call $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{V}$ the Bethe algebra for $V$ with respect to $\bf z$}.
+4.3. The Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$
+Let $\mathcal{F}:={\mathbb{C}}[{\bf x},{\bf y},{\bm{\eta}},{\bm{\zeta}}]$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$, and $\zeta^{a}_{s}$, for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$, and $a=1,\ldots,d$.
+Here $x^{a}_{i}$ and $y^{a}_{r}$ are even variables while $\eta^{a}_{j}$ and $\zeta^{a}_{s}$ are odd.
+With an appropriate inner product, $\mathcal{F}$ may be regarded as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators (cf.
+[CL, LZ]).
+Let $\mathcal{D}:={\mathbb{D}}[{\bf x},{\bf y},{\bm{\eta}},{\bm{\zeta}}]$ be the corresponding Weyl superalgebra, i.e., $\mathcal{D}$ is the associative unital superalgebra generated by $x^{a}_{i},\,\eta^{a}_{j},\,y^{a}_{r},\,\zeta^{a}_{s}$ as well as their derivatives
 $\partial_{x^{a}_{i}}:=\frac{\partial}{\partial x_{i}^{a}},\quad\partial_{\eta^{a}_{j}}:=\frac{\partial}{\partial\eta_{j}^{a}},\quad\partial_{y^{a}_{r}}:=\frac{\partial}{\partial y_{r}^{a}},\quad\partial_{\zeta^{a}_{s}}:=\frac{\partial}{\partial\zeta_{r}^{a}}$
-for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$ and $a=1,\ldots,d$.
-Here $x^{a}_{i}$, $y^{a}_{r}$, $\partial_{x^{a}_{i}}$ and $\partial_{y^{a}_{r}}$ are even while $\eta^{a}_{j}$, $\zeta^{a}_{s}$, $\partial_{\eta^{a}_{j}}$, and $\partial_{\zeta^{a}_{s}}$ are odd.
-The superalgebra $\mathcal{D}$ is called the Weyl superalgebra.
-There is an anti-involution $\omega:\mathcal{D}\longrightarrow\mathcal{D}$ defined by
-${\begin{array}[]{lll}x^{a}_{i}\mapsto\partial_{x^{a}_{i}},&\quad\partial_{x^{a}_{i}}\mapsto x^{a}_{i},\quad\eta^{a}_{j}\mapsto\partial_{\eta^{a}_{j}},&\quad\partial_{\eta^{a}_{j}}\mapsto\eta^{a}_{j},\\ y^{a}_{r}\mapsto\partial_{y^{a}_{r}},&\quad\partial_{y^{a}_{r}}\mapsto y^{a}_{r},\quad\zeta^{a}_{s}\mapsto\partial_{\zeta^{a}_{s}},&\quad\partial_{\zeta^{a}_{s}}\mapsto\zeta^{a}_{s},\end{array}}$ (4.1)
-for $1\leq i\leq m$, $1\leq j\leq n$, $1\leq r\leq p$, and $1\leq s\leq q$. In light of Proposition 3.6, the map $\omega$ extends to an anti-involution
-$\omega:\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!).$
+for $1\leq i\leq m$, $1\leq j\leq n$, $1\leq r\leq p$, $1\leq s\leq q$, and $1\leq a\leq d$.
+The superalgebra $\mathcal{D}$ acts naturally on $\mathcal{F}$, and so it is a subalgebra of ${\rm End}(\mathcal{F})$.
 According to [CLZ, pp. 789–790], there is a superalgebra homomorphism $\phi:U({\mathfrak{gl}}_{d})\longrightarrow\mathcal{D}$, given by
 $e_{ab}\mapsto\sum_{i=1}^{m}x_{i}^{a}\partial_{x^{b}_{i}}+\sum_{j=1}^{n}\eta_{j}^{a}\partial_{\eta^{b}_{j}}-\sum_{r=1}^{p}y_{r}^{b}\partial_{y^{a}_{r}}-\sum_{s=1}^{q}\zeta_{s}^{b}\partial_{\zeta^{a}_{s}},$ (4.9)
-for $1\leq a,b\leq d$, and a superalgebra homomorphism $\overline{\phi}:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$, given by
-${\begin{array}[]{cc}E^{r}_{r^{\prime}}\mapsto-\sum_{a=1}^{d}\partial_{y^{a}_{r}}y_{r^{\prime}}^{a},&E^{r}_{p+s}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\zeta_{s}^{a},\\ E^{p+s}_{r}\mapsto-\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}y_{r}^{a},&E^{p+s}_{p+s^{\prime}}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\zeta_{s^{\prime}}^{a},\\ E^{r}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{x^{a}_{i}},&E^{r}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{\eta^{a}_{j}},\\ E^{p+s}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{x^{a}_{i}},&E^{p+s}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{\eta^{a}_{j}},\\ E^{p+q+i}_{r}\mapsto-\sum_{a=1}^{d}x^{a}_{i}y_{r}^{a},&E^{p+q+i}_{p+s}\mapsto\sum_{a=1}^{d}x^{a}_{i}\zeta_{s}^{a},\\ E^{p+q+m+j}_{r}\mapsto-\sum_{a=1}^{d}\eta^{a}_{j}y_{r}^{a},&E^{p+q+m+j}_{p+s}\mapsto\sum_{a=1}^{d}\eta^{a}_{j}\zeta_{s}^{a},\\ E^{p+q+i}_{p+q+i^{\prime}}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{x^{a}_{i^{\prime}}},&E^{p+q+i}_{p+q+m+j}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{\eta^{a}_{j}},\\ E^{p+q+m+j}_{p+q+i}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{x^{a}_{i}},&E^{p+q+m+j}_{p+q+m+j^{\prime}}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{\eta^{a}_{j^{\prime}}}.\end{array}}$ (4.3)
-where $1\leq i,i^{\prime}\leq m$, $1\leq j,j^{\prime}\leq n$, $1\leq r,r^{\prime}\leq p$, and $1\leq s,s^{\prime}\leq q$.
-Let $\mathcal{F}$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$ and $\zeta^{a}_{s}$, for $1\leq i\leq m$, $1\leq j\leq n$, $1\leq r\leq p$, and $1\leq s\leq q$.
-With an appropriate inner product, $\mathcal{F}$ may be regarded as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators.
-The Weyl superalgebra $\mathcal{D}$ acts naturally on $\mathcal{F}$, and so it is a subalgebra of ${\rm End}(\mathcal{F})$.
-We simplify notations for the rest of this section. Let
+for $1\leq a,b\leq d$, and a superalgebra homomorphism $\varphi:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$, given by
+$\begin{array}[]{cc}E^{r}_{r^{\prime}}\mapsto-\sum_{a=1}^{d}\partial_{y^{a}_{r}}y_{r^{\prime}}^{a},\quad&E^{r}_{p+s}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\zeta_{s}^{a},\\ E^{p+s}_{r}\mapsto-\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}y_{r}^{a},\quad&E^{p+s}_{p+s^{\prime}}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\zeta_{s^{\prime}}^{a},\\ E^{r}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{x^{a}_{i}},\quad&E^{r}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{\eta^{a}_{j}},\\ E^{p+s}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{x^{a}_{i}},\quad&E^{p+s}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{\eta^{a}_{j}},\\ E^{p+q+i}_{r}\mapsto-\sum_{a=1}^{d}x^{a}_{i}y_{r}^{a},\quad&E^{p+q+i}_{p+s}\mapsto\sum_{a=1}^{d}x^{a}_{i}\zeta_{s}^{a},\\ E^{p+q+m+j}_{r}\mapsto-\sum_{a=1}^{d}\eta^{a}_{j}y_{r}^{a},\quad&E^{p+q+m+j}_{p+s}\mapsto\sum_{a=1}^{d}\eta^{a}_{j}\zeta_{s}^{a},\quad\\ E^{p+q+i}_{p+q+i^{\prime}}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{x^{a}_{i^{\prime}}},\quad&E^{p+q+i}_{p+q+m+j}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{\eta^{a}_{j}},\\ E^{p+q+m+j}_{p+q+i}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{x^{a}_{i}},\quad&E^{p+q+m+j}_{p+q+m+j^{\prime}}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{\eta^{a}_{j^{\prime}}}.\end{array}$} (4.10)
+where $1\leq r,r^{\prime}\leq p$, $1\leq s,s^{\prime}\leq q$, $1\leq i,i^{\prime}\leq m$, and $1\leq j,j^{\prime}\leq n$.
+These maps induce a ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-action on $\mathcal{F}$.
+Let ${\bm{\pi}}:{\mathfrak{gl}}_{d}\longrightarrow{\mathfrak{gl}}_{d}$ be the *-structure on ${\mathfrak{gl}}_{d}$ defined by
+${\bm{\pi}}(e_{ab})=e_{ba}\qquad\mbox{for $1\leq a,b\leq d$.}$
+Note that any unitarizable highest weight ${\mathfrak{gl}}_{d}$-module with respect to $\bm{\pi}$} is finite-dimensional.
+Recall the *-structure ${\bm{\sigma}}_{p}$ on ${\mathfrak{gl}}_{p+m|q+n}$ defined by (3.4).
+The maps $\bm{\pi}$} and ${\bm{\sigma}}_{p}$ give rise to a *-structure $({\bm{\pi}},{\bm{\sigma}}_{p})$ on ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$.
+It is well-known that $\mathcal{F}$ decomposes into a direct sum of unitarizable ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-modules (with respect to $({\bm{\pi}},{\bm{\sigma}}_{p})$).
+This is called the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ ([CLZ]).
+Such unitarizable modules are, in general, infinite-dimensional.
+Taking $p=q=0$, the maps $\phi$ and $\varphi$ specialize to the superalgebra homomorphisms $U({\mathfrak{gl}}_{d})\longrightarrow{\mathbb{D}}[{\bf x},{\bm{\eta}}]$ and $U({\mathfrak{gl}}_{m|n})\longrightarrow{\mathbb{D}}[{\bf x},{\bm{\eta}}]$, respectively, and the duality reduces to the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$, which yields a decomposition of ${\mathbb{C}}[{\bf x},{\bm{\eta}}]$ into a direct sum of finite-dimensional ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{m|n}$-modules ([CW, Se1, Se2]).
+Assuming further that $n=0$, it gives the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ ([H]).
+We simplify notations for the rest of this subsection. Let
 $x^{a}_{m+j}=\eta^{a}_{j},\quad\partial_{x^{a}_{m+j}}=\partial_{\eta^{a}_{j}},\quad y^{a}_{p+s}=\xi^{a}_{s},\quad\partial_{y^{a}_{p+s}}=\partial_{\zeta^{a}_{s}},$
@@
 The map $\phi$ induces a superalgebra homomorphism $\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])\longrightarrow\mathcal{D}$ defined by
-$e_{ab}[k]\mapsto\sum_{i=1}^{m+n}z_{p+q+i}^{k}x^{a}_{i}\partial_{x^{b}_{i}}-\sum_{r=1}^{p+q}z_{r}^{k}y^{b}_{r}\partial_{y^{a}_{r}},$ (4.4)
-for $1\leq a,b\leq d$ and $k\in{\mathbb{Z}}_{+}$, and $\overline{\phi}$ induces a superalgebra homomorphism $\overline{\phi}_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])\longrightarrow\mathcal{D}$ defined by
+$e_{ab}[k]\mapsto-\sum_{r=1}^{p+q}z_{r}^{k}y^{b}_{r}\partial_{y^{a}_{r}}+\sum_{i=1}^{m+n}z_{p+q+i}^{k}x^{a}_{i}\partial_{x^{b}_{i}},$ (4.11)
+for $1\leq a,b\leq d$ and $k\in{\mathbb{Z}}_{+}$, and $\varphi$ induces a superalgebra homomorphism $\varphi_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])\longrightarrow\mathcal{D}$ defined by
 $\begin{array}[]{cll}E^{r}_{s}[k]&\mapsto\sum_{a=1}^{d}(-1)^{|s|+1}w_{a}^{k}\partial_{y^{a}_{r}}y^{a}_{s},\quad E^{r}_{p+q+j}[k]&\mapsto\sum_{a=1}^{d}w_{a}^{k}\partial_{y^{a}_{r}}\partial_{x^{a}_{j}},\\ E^{p+q+i}_{s}[k]&\mapsto\sum_{a=1}^{d}(-1)^{|s|+1}w_{a}^{k}x^{a}_{i}y^{a}_{s},\quad E^{p+q+i}_{p+q+j}[k]&\mapsto\sum_{a=1}^{d}w_{a}^{k}x^{a}_{i}\partial_{x^{a}_{j}},\end{array}$} (4.12)
-for $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$.
-The Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ are respectively commutative subalgebras of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$.
-Thus, the maps $\phi_{\bf z}$ and $\overline{\phi}_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ and an action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$, respectively, on the polynomial superalgebra $\mathcal{F}$.
-The Cartan anti-involution $\boldsymbol{\theta}$ on ${\mathfrak{gl}}_{d}$, defined in (3.7), gives rise to an anti-involution $\boldsymbol{\theta}:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$, defined by
-$\hskip 42.67912pt\boldsymbol{\theta}(A[r])=\boldsymbol{\theta}(A)[r]\qquad\mbox{for $A\in{\mathfrak{gl}}_{d}$ and $r\in{\mathbb{Z}}_{+}$}.$
-Proposition 4.1 ([MTV1, Proposition 8.4]).
-The anti-involution $\boldsymbol{\theta}:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$ fixes each element of $\mathcal{B}_{d}^{\bf w}$.
-Recall the anti-involution $\omega$ on $\mathcal{D}$ defined in (4.1). It is straightforward to verify that the diagram
-[…]
-commutes.
-Proposition 4.2.
-For any ${\boldsymbol{b}}\in\mathcal{B}_{d}^{\bf w}$, $\omega\big{(}\phi_{\bf z}({\boldsymbol{b}})\big{)}=\phi_{\bf z}({\boldsymbol{b}})$.
-Proof.
-This follows from (4.6) and Proposition 4.1. ∎
-4.2. From Berezinians to column determinants
-The map $\phi_{\bf z}$ and $\overline{\phi}_{\bf w}$, defined in (4.4) and (4.5), extend naturally to the superalgebra homomorphisms
+for $1\leq r,s\leq p+q$, $1\leq i,j\leq m+n$, and $k\in{\mathbb{Z}}_{+}$.
+Since the Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and $\mathcal{B}_{p+m|q+n}^{\bf z}$} are subalgebras of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$, respectively, the maps $\phi_{\bf z}$ and $\varphi_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ on $\mathcal{F}$ and an action of $\mathcal{B}_{p+m|q+n}^{\bf z}$} on $\mathcal{F}$, respectively.
+In view of the Howe duality, it is natural to ask if any duality exists in the context of Bethe algebras as well.
+When $p=q=n=0$, a duality phenomenon is noticed in [TL] (see also [TV]).
+In [MTV3], the precise statement is made and proved.
+So far, various generalizations have emerged (see [HM, TU, VY]).
+We will show that the desired duality exists between $\mathcal{B}_{d}^{\bf w}$ and $\mathcal{B}_{p+m|q+n}^{\bf z}$} in a moment.
+The maps $\phi_{\bf z}$ and $\varphi_{\bf w}$ extend naturally to the superalgebra homomorphisms
 $\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$
 and
-$\overline{\phi}_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!),$
-respectively.
-Let
-$\mathcal{C}=\left[\delta_{a,b}(z-w_{a})-\left(\sum_{i=1}^{m+n}\frac{x^{b}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}-\sum_{r=1}^{p+q}\frac{y^{a}_{r}\partial_{y^{b}_{r}}}{\partial_{z}-z_{r}}\right)\right]_{a,b=1,\ldots,d}.$
-Lemma 4.3.
-We have
-$\omega\big{(}\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\big{)}={\rm cdet}\,(\mathcal{C}).$
-Proof.
-For $1\leq a,b\leq d$,
-$\phi_{\bf z}\big{(}e_{ab}(z)\big{)}=\sum_{i=1}^{m+n}\frac{x^{a}_{i}\partial_{x^{b}_{i}}}{z-z_{p+q+i}}-\sum_{r=1}^{p+q}\frac{y^{b}_{r}\partial_{y^{a}_{r}}}{z-z_{r}}.$ (4.7)
-We define
-$F^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(\partial_{z}-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{b}_{i}\partial_{y^{a}_{i}}}{z-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{a}_{i-p-q}\partial_{x^{b}_{i-p-q}}}{z-z_{i}}}&\mbox{if }\,p+q+1\leq i\leq p+q+m+n,\end{cases}$
-and
-$G^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(z-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{a}_{i}\partial_{y^{b}_{i}}}{\partial_{z}-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{b}_{i-p-q}\partial_{x^{a}_{i-p-q}}}{\partial_{z}-z_{i}}}&\mbox{if }\,p+q+1\leq i\leq p+q+m+n.\end{cases}$
-We have
-$\omega(F^{i}_{a,b})=G^{i}_{a,b}$ (4.8)
-for all $0\leq i\leq{p+q+m+n}$ and $1\leq a,b\leq d$.
-Note that ${\mathcal{L}}_{d}^{\bf w}$ is an amply invertible Manin matrix of type $(0^{d})$.
-Let $\pi\in\mathfrak{S}_{d}$ be the permutation given by $\pi(i)=d+1-i$ for $i=1,\ldots,d$.
-Clearly, $\pi^{2}=1$.
-By (4.7), Proposition 3.5 and Proposition 3.12, we obtain
-$\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})$ $=$ $\phi_{\bf z}{\rm cdet}\,\big{(}({\mathcal{L}}_{d}^{\bf w})^{\pi}\big{)}$
-$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\pi(\sigma(1)),\pi(1)}\ldots F^{i_{1}}_{\pi(\sigma(d)),\pi(d)}$
-$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\pi\sigma(1),d}\ldots F^{i_{1}}_{\pi\sigma(d),1}$
-$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\sigma(d),d}\ldots F^{i_{1}}_{\sigma(1),1}.$
-Here the last equality holds since $\pi^{-1}\mathfrak{S}_{d}\pi=\mathfrak{S}_{d}$ and $l(\pi^{-1}\sigma\pi)=l(\sigma)$ for all $\sigma\in\mathfrak{S}_{d}$. Thus,
-$\omega\big{(}\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\big{)}=\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}\omega\!\left(F^{i_{1}}_{\sigma(1),1}\right)\ldots\omega\!\left(F^{i_{d}}_{\sigma(d),d}\right).$
-In view of (4.8), we deduce that $\omega\big{(}{\rm cdet}\,\phi_{\bf z}\,{\mathcal{L}}_{d}^{\bf w}\big{)}={\rm cdet}\,(\mathcal{C})$. ∎
-Fix ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. For $u=z$ or $\partial_{z}$, let
+$\varphi_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!),$
+respectively. Furthermore, there is an anti-involution $\omega:\mathcal{D}\longrightarrow\mathcal{D}$ defined by
+$x^{a}_{i}\mapsto\partial_{x^{a}_{i}},\quad\partial_{x^{a}_{i}}\mapsto x^{a}_{i},\quad y^{a}_{r}\mapsto\partial_{y^{a}_{r}},\quad\partial_{y^{a}_{r}}\mapsto y^{a}_{r},$ (4.13)
+for $1\leq i\leq m+n$, $1\leq r\leq p+q$, and $1\leq a\leq d$. In light of Proposition 2.1, the map $\omega$ extends naturally to an anti-involution
+$\omega:\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!).$
+For $u=z$ or $\partial_{z}$, let
 $R_{p|q}(u)=\frac{(u-z_{1})\ldots(u-z_{p})}{(u-z_{p+1})\ldots(u-z_{p+q})}$
@@
 $R^{\prime}_{m|n}(u)=\frac{(u-z_{p+q+1})\ldots(u-z_{p+q+m})}{(u-z_{p+q+m+1})\ldots(u-z_{p+q+m+n})}.$
+We have the following proposition. We defer its proof until Appendix A to avoid disrupting the flow of the paper.
 Proposition 4.7.
-$(z-w_{1})\ldots(z-w_{d})\,\overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)=\omega\left(\!R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\cdot R_{p|q}(z)\!\right).$
+Let ${\bf s}=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. Then
+$(z-w_{1})\ldots(z-w_{d})\,\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)=\omega\left(\!R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\cdot R_{p|q}(z)\!\right).$
+The Cartan anti-involution $\varpi$ on ${\mathfrak{gl}}_{d}$, defined by (3.14), gives rise to an anti-involution $\varpi:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$, defined by
+$\hskip 42.67912pt\varpi(A[k])=\varpi(A)[k]\qquad\mbox{for $A\in{\mathfrak{gl}}_{d}$ and $k\in{\mathbb{Z}}_{+}$}.$
+Proposition 4.8 ([MTV1, Proposition 8.4]).
+The anti-involution $\varpi:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$ fixes all elements of $\mathcal{B}_{d}^{\bf w}$.
+Proposition 4.9.
+For any ${\bm{b}}\in\mathcal{B}_{d}^{\bf w}$, $\omega\big(\phi_{\bf z}({\bm{b}})\big)=\phi_{\bf z}({\bm{b}})$.
 Proof.
-For $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$, we have
-${\begin{array}[]{cll}\overline{\phi}_{\bf w}\big{(}E^{r}_{s}(z)\big{)}&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}\partial_{y^{a}_{r}}y^{a}_{s}}{z-w_{a}},\quad\overline{\phi}_{\bf w}\big{(}E^{r}_{p+q+j}(z)\big{)}&=\sum_{a=1}^{d}\frac{\partial_{y^{a}_{r}}\partial_{x^{a}_{j}}}{z-w_{a}},\\ \overline{\phi}_{\bf w}\big{(}E^{p+q+i}_{s}(z)\big{)}&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}x^{a}_{i}y^{a}_{s}}{z-w_{a}},\quad\overline{\phi}_{\bf w}\big{(}E^{p+q+i}_{p+q+j}(z)\big{)}&=\sum_{a=1}^{d}\frac{x^{a}_{i}\partial_{x^{a}_{j}}}{z-w_{a}}.\end{array}}$ (4.9)
-Let
-$D=\Big{[}\delta_{a,b}(z-w_{a})\Big{]}_{a,b=1,\ldots,d}\quad\text{and}\quad D^{\prime}=\Big{[}\delta_{i,j}(\partial_{z}-z_{i})\Big{]}_{i,j\in\mathbb{I}}.$
-Consider the $(m+n)\times d$ matrices
-$X:=\big{[}(-1)^{|i|}x^{a}_{i}\big{]}^{a=1,\ldots,d}_{i=1,\ldots,m+n}\quad\text{and}\quad P_{X}:=\big{[}\partial_{x^{a}_{i}}\big{]}^{a=1,\ldots,d}_{i=1,\ldots,m+n}$
-and the $(p+q)\times d$ matrices
-$Y:=\big{[}(-1)^{|r|+1}y^{a}_{r}\big{]}^{a=1,\ldots,d}_{r=1,\ldots,p+q}\qquad\text{and}\qquad P_{Y}:=\big{[}(-1)^{|r|}\partial_{y^{a}_{r}}\big{]}^{a=1,\ldots,d}_{r=1,\ldots,p+q}.$
-Write
-$T=\begin{bmatrix}Y^{t}\hskip 5.69054ptP_{X}^{t}\end{bmatrix}\quad\text{and}\quad T^{\prime}=\begin{bmatrix}P_{Y}\vspace{3mm}\\ X\end{bmatrix}.$
-By (4.9), we get
-$\overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)={\rm Ber}^{\bf s}\,(D^{\prime}-T^{\prime}D^{-1}T).$
-Let
-$\mathfrak{L}=\begin{bmatrix}D&T\vspace{2mm}\\ T^{\prime}&D^{\prime}\end{bmatrix}.$
-By Proposition 3.1, we have
-${\rm Ber}\,^{\!\hat{{\bf s}}}\,(\mathfrak{L})=(z-w_{1})\ldots(z-w_{d})\,\overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right),$
-where $\hat{{\bf s}}:=(0^{d},{\bf s})\in\mathcal{S}_{d+p+m|q+n}$.
-As $D-T(D^{\prime})^{-1}T^{\prime}$ is a sufficiently invertible Manin matrix of type $(0^{d})$, Proposition 3.4 and Proposition 3.5 imply that
-${\rm Ber}\,^{\!\hat{{\bf s}}}\,(\mathfrak{L})$ $=$ ${\rm Ber}^{\bf s}\,(D^{\prime})\cdot{\rm cdet}\,(D-T(D^{\prime})^{-1}T^{\prime})$
-$=$ $R_{p|q}(\partial_{z})\cdot R^{\prime}_{m|n}(\partial_{z})\cdot{\rm cdet}\,(\mathcal{C}^{\prime}),$
-where
-$\mathcal{C}^{\prime}:=\left[\delta_{a,b}(z-w_{a})-\left(\sum_{i=1}^{m+n}\frac{(-1)^{|i|}\partial_{x^{a}_{i}}x^{b}_{i}}{\partial_{z}-z_{p+q+i}}-\sum_{r=1}^{p+q}\frac{y^{a}_{r}\partial_{y^{b}_{r}}}{\partial_{z}-z_{r}}\right)\right]_{a,b=1,\ldots,d}.$
-We claim that
-$R^{\prime}_{m|n}(\partial_{z})\cdot{\rm cdet}\,(\mathcal{C}^{\prime})={\rm cdet}\,(\mathcal{C})\cdot R^{\prime}_{m|n}(\partial_{z}).$
-This yields
-${\rm Ber}\,^{\!\hat{{\bf s}}}\,(\mathfrak{L})=R_{p|q}(\partial_{z})\cdot{\rm cdet}\,(\mathcal{C})\cdot R^{\prime}_{m|n}(\partial_{z}),$
-which proves the proposition by Lemma 4.3.
-To prove the claim, it suffices to show that
-$R^{\prime}_{m|n}(\partial_{z})\cdot\mathcal{C}^{\prime}_{a,b}=\mathcal{C}_{a,b}\cdot R^{\prime}_{m|n}(\partial_{z}),\qquad\mbox{for $1\leq a,b\leq d$.}$ (4.10)
-Here, for instance, $\mathcal{C}_{a,b}$ denotes the $(a,b)$ entry of $\mathcal{C}$. Since
-$\partial_{x^{a}_{i}}x^{b}_{i}=(-1)^{|i|}x^{b}_{i}\partial_{x^{a}_{i}}+\delta_{a,b},\qquad\mbox{for $1\leq a,b\leq d$},$ (4.11)
-the off-diagonal entries of $\mathcal{C}^{\prime}$ coincide with those of $\mathcal{C}$.
-For $1\leq i\leq m+n$, the term $\partial_{z}-z_{p+q+i}$ commutes with $\partial_{x^{a}_{j}}x^{b}_{j}$ and $y^{a}_{r}\partial_{y^{b}_{r}}$ for all $1\leq j\leq m+n$, $1\leq r\leq p+q$ and $1\leq a,b\leq d$.
-In particular, $R^{\prime}_{m|n}(\partial_{z})$ commutes with any off-diagonal entry of $\mathcal{C}^{\prime}$.
-This proves (4.10) for $a\not=b$.
-Moreover, for $1\leq i\leq m+n$ and $1\leq a\leq d$,
-$(\partial_{z}-z_{p+q+i})(z-w_{a})=(z-w_{a})(\partial_{z}-z_{p+q+i})+1.$ (4.12)
-If $1\leq i\leq m$, we see from (4.11) and (4.12) that
-${(\partial_{z}-z_{p+q+i})}\left(z-w_{a}-\frac{\partial_{x^{a}_{i}}x^{a}_{i}}{\partial_{z}-z_{p+q+i}}\right)=\left(z-w_{a}-\frac{x^{a}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}\right)(\partial_{z}-z_{p+q+i}).$
-Similarly, if $m+1\leq i\leq m+n$, we have
-${(\partial_{z}-z_{p+q+i})^{-1}}\left(z-w_{a}+\frac{\partial_{x^{a}_{i}}x^{a}_{i}}{\partial_{z}-z_{p+q+i}}\right)=\left(z-w_{a}-\frac{x^{a}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}\right)(\partial_{z}-z_{p+q+i})^{-1}.$
-The above discussion proves (4.10) for $a=b$. Thus, the proof of (4.10) is now complete. ∎
-We relate ${\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)$ to ${\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})$ as follows.
-Proposition 4.5.
-We may write
+It is straightforward to verify that the diagram
+[…]
+commutes. This, together with Proposition 4.8, proves the proposition. ∎
+The following theorem is called the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$.
+It not only relates the coefficients of $\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ to those of $\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$, but also gives an equality of algebras.
+Theorem 4.10.
+Let ${\bf s}=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. We may write
 $R_{p|q}(z)\cdot R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})=\sum_{\ell=0}^{d}\sum_{k=-\infty}^{p+m-q-n}b_{k,\ell}^{{\bf w},{\bf z}}z^{k}\partial_{z}^{\ell},$
 where $b_{k,\ell}^{{\bf w},{\bf z}}:=\phi_{\bf z}(b_{k,\ell}^{\bf w})$ for some $b_{k,\ell}^{\bf w}\in\mathcal{B}_{d}^{\bf w}$. Then
-$(z-w_{1})\ldots(z-w_{d})\ \overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)$
+$(z-w_{1})\ldots(z-w_{d})\ \varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$
 $\hskip 85.35826pt=R_{p|q}(\partial_{z})\left(\sum_{k=-\infty}^{p+m-q-n}\sum_{\ell=0}^{d}b_{k,\ell}^{{\bf w},{\bf z}}z^{\ell}\partial_{z}^{k}\right)R_{p|q}(\partial_{z})^{-1}.$
+Moreover, $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\varphi_{\bf w}\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)$.
 Proof.
 By Proposition 4.7,
-$(z-w_{1})\ldots(z-w_{d})\,\overline{\phi}_{\bf w}{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)$
+$(z-w_{1})\ldots(z-w_{d})\,\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$
 $\hskip 42.67912pt=R_{p|q}(\partial_{z})\cdot\omega\left(R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\right)\!$
-$\hskip 42.67912pt=R_{p|q}(\partial_{z})\cdot\omega\left(R_{p|q}(z)\cdot R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\right)\cdot R_{p|q}(\partial_{z})^{-1}.$
-Therefore,
-$(z-w_{1})\ldots(z-w_{d})\ \overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)$
-$\hskip 85.35826pt=R_{p|q}(\partial_{z})\left(\sum_{k=-\infty}^{p+m-q-n}\sum_{\ell=0}^{d}\omega\big{(}b_{k,\ell}^{{\bf w},{\bf z}}\big{)}z^{\ell}\partial_{z}^{k}\right)R_{p|q}(\partial_{z})^{-1}.$
-By Proposition 4.2,
-$\omega\big{(}b_{k,\ell}^{{\bf w},{\bf z}}\big{)}=b_{k,\ell}^{{\bf w},{\bf z}}$
-for all $k\leq p+m-q-n$ and $0\leq\ell\leq d$. This completes the proof. ∎
-4.3. Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$
-Lemma 4.6.
-Let $c\in\{z_{1},\ldots,z_{p+q}\}$.
-Given $\alpha_{r,s}^{{\bf w},{\bf z}}\in\mathcal{D}$ for $r,s\in{\mathbb{Z}}$ with $0\leq r\leq d$ and $s\leq p+m-q-n$.
-Let $\mathcal{D}_{1}$ be the subalgebra of $\mathcal{D}$ generated by all $\alpha_{r,s}^{{\bf w},{\bf z}}$’s, and let $\mathcal{D}_{2}$ be the subalgebra of $\mathcal{D}$ generated by the coefficients of the series
-$(\partial_{z}-c)\left(\sum_{s=-\infty}^{p+m-q-n}\sum_{r=0}^{d}\alpha_{r,s}^{{\bf w},{\bf z}}z^{r}\partial_{z}^{s}\right)(\partial_{z}-c)^{-1}.$ (4.13)
-Then $\mathcal{D}_{1}=\mathcal{D}_{2}$.
-Proof.
-For any $r\in{\mathbb{Z}}_{+}$,
-$(\partial_{z}-c)z^{r}(\partial_{z}-c)^{-1}=z^{r}+rz^{r-1}(\partial_{z}-c)^{-1}.$
-Let $f$ denote the series (4.13). We find that
-$f=\sum_{s=-\infty}^{p+m-q-n}\sum_{r=0}^{d}\left(\alpha_{r,s}^{{\bf w},{\bf z}}z^{r}\partial_{z}^{s}+r\alpha_{r,s}^{{\bf w},{\bf z}}z^{r-1}\partial_{z}^{s}(\partial_{z}-c)^{-1}\right)\!,$
-and hence
-$f=\sum_{\ell=-\infty}^{p+m-q-n}\sum_{k=0}^{d}\beta_{k,\ell}^{{\bf w},{\bf z}}z^{k}\partial_{z}^{\ell},$
-where
-$\beta_{d,\ell}^{{\bf w},{\bf z}}=\alpha_{d,\ell}^{{\bf w},{\bf z}},$ (4.14)
-$\beta_{k,\ell}^{{\bf w},{\bf z}}=\alpha_{k,\ell}^{{\bf w},{\bf z}}+\sum_{i=0}^{\infty}(k+1)\alpha_{k+1,\ell+i+1}^{{\bf w},{\bf z}}c^{i},$ (4.15)
-for $10\leq k\leq d-$ 1 and $\ell\leq p+m-q-n$. Note that the right-hand side of (4.15) has only finitely many nonzero terms.
-The algebra $\mathcal{D}_{2}$ is, by definition, generated by $\beta_{k,\ell}^{{\bf w},{\bf z}}$’s.
-So, it is a subalgebra of $\mathcal{D}_{1}$ by (4.14) and (4.15).
-Also, (4.14) means that $\alpha_{d,\ell}^{{\bf w},{\bf z}}\in\mathcal{D}_{2}$ for all $\ell\leq p+m-q-n$.
-Using (4.15), we see, by induction, that $\alpha_{r,\ell}^{{\bf w},{\bf z}}\in\mathcal{D}_{2}$ for all $0\leq r\leq d$ and $\ell\leq p+m-q-n$.
-It follows that $\mathcal{D}_{1}\subseteq\mathcal{D}_{2}$, and thus $\mathcal{D}_{1}=\mathcal{D}_{2}$.
+$\hskip 42.67912pt=R_{p|q}(\partial_{z})\cdot\omega\left(R_{p|q}(z)\cdot R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\right)\cdot R_{p|q}(\partial_{z})^{-1}$
+$\hskip 42.67912pt=R_{p|q}(\partial_{z})\left(\sum_{k=-\infty}^{p+m-q-n}\sum_{\ell=0}^{d}\omega\big(b_{k,\ell}^{{\bf w},{\bf z}}\big)z^{\ell}\partial_{z}^{k}\right)R_{p|q}(\partial_{z})^{-1}.$
+By Proposition 4.9, $\omega\big(b_{k,\ell}^{{\bf w},{\bf z}}\big)=b_{k,\ell}^{{\bf w},{\bf z}}$ for all $k\leq p+m-q-n$ and $0\leq\ell\leq d$.
+This proves the first assertion.
+The last assertion clearly follows from the first one.
 ∎
-We have the following duality between the Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$.
-We call it the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$.
-Theorem 4.7.
-$\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}$.
-Proof.
-This follows from Proposition 4.5 and repeated use of Lemma 4.6. ∎
 Remark 4.11.
-We will demonstrate another duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for Gaudin models with irregular singularities, which specializes to a variant of Theorem 4.7, in [ChL3].
-The classical version of the duality will also be investigated there.
-If some of $p,q,m,n$ are 0, we will suppress (most of) them in the symbol ${\mathcal{B}_{p+m|q+n}^{\bf z}}$. For instance,
-$\mathcal{B}^{\bf z}_{p+m|n}:=\mathcal{B}^{\bf z}_{p+m|0+n},\quad\mathcal{B}^{\bf z}_{m|n}:=\mathcal{B}^{\bf z}_{0+m|0+n},\quad\mathcal{B}^{\bf z}_{0|n}:=\mathcal{B}^{\bf z}_{0+0|0+n},\quad\text{etc.}$.
-(4.16)
-Theorem 4.7 gives the Bethe dualities of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|n})$ and $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m})$.
-Corollary 4.9.
-  1. (i)
-    Suppose $q=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{p+m|n}\big{)}$.
-  2. (ii)
-    Suppose $q=n=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{p+m}\big{)}$.
-Theorem 4.7 recovers the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$ due to Huang and Mukhin as well as the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ due to Mukhin, Tarasov, and Varchenko.
-Corollary 4.10 ([HM, Theorem 5.2]).
-Suppose $p=q=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{m|n}\big{)}$.
-Corollary 4.11 ([MTV3, Theorem 6]).
-Suppose $p=q=n=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{m}\big{)}$.
-We also obtain a variant of the duality of Bethe algebras established by Tarasov and Uvarov (cf. [TU, Theorem 5.2]).
-Corollary 4.12.
-Suppose $p=q=m=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{0|n}\big{)}$.
-5. An application of Bethe duality
-In this section, we recall *-structures and unitarizable modules.
-We apply the Bethe duality to the Bethe algebra of ${\mathfrak{gl}}_{p+m|q+n}$ acting on a tensor product of infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules.
-5.1. *-structures and unitarizable modules
-A *-superalgebra is an associative superalgebra $A$ with an even anti-linear anti-involution ${\boldsymbol{\sigma}}:A\longrightarrow A$.
-A homomorphism $f:(A,{\boldsymbol{\sigma}})\longrightarrow(A^{\prime},{\boldsymbol{\sigma}}^{\prime})$ of *-superalgebras is a homomorphism of superalgebras satisfying ${\boldsymbol{\sigma}}^{\prime}\circ f=f\circ{\boldsymbol{\sigma}}$.
-[…]
-$\langle av|v^{\prime}\rangle=\langle v|{\boldsymbol{\sigma}}(a)v^{\prime}\rangle,$
-for all $a\in A$ and $v,v^{\prime}\in V$. An $A$-module equipped with a positive definite contravariant Hermitian form is called a unitarizable $A$-module (with respect to ${\boldsymbol{\sigma}}$).
-Let ${\mathfrak{g}}$ be a Lie superalgebra.
-An anti-linear anti-involution ${\boldsymbol{\sigma}}:{\mathfrak{g}}\longrightarrow{\mathfrak{g}}$ is called a *-structure on ${\mathfrak{g}}$.
-A homomorphism $f:({\mathfrak{g}},{\boldsymbol{\sigma}})\longrightarrow({\mathfrak{g}}^{\prime},{\boldsymbol{\sigma}}^{\prime})$ of Lie superalgebras with *-structures is a homomorphism of Lie superalgebras satisfying ${\boldsymbol{\sigma}}^{\prime}\circ f=f\circ{\boldsymbol{\sigma}}$.
-Moreover, ${\boldsymbol{\sigma}}$ is a *-structure on ${\mathfrak{g}}$ if and only if the natural extension of ${\boldsymbol{\sigma}}$ to $U({\mathfrak{g}})$ is an anti-linear anti-involution.
-[…]
-The Lie superalgebra ${\mathfrak{gl}}_{p+m|q+n}$ admits a *-structure ${\boldsymbol{\sigma}}:{\mathfrak{gl}}_{p+m|q+n}\longrightarrow{\mathfrak{gl}}_{p+m|q+n}$ (see [CLZ, Section 3.2]), defined by
-$\hskip 42.67912pt\sum_{i,j\in\mathbb{I}}a_{ij}E^{i}_{j}\mapsto\sum_{i,j\in\mathbb{I}}(-1)^{[i]+[j]}\bar{a}_{ij}E^{j}_{i}\qquad\mbox{for $1\leq i,j\leq p+q+m+n$},$
-where
-$[i]:=\begin{cases}\,0&\,\,\mbox{if}\ \ 1\leq i\leq p,\\ \,1&\,\,\mbox{otherwise},\end{cases}$
-and $\bar{a}_{ij}$ denotes the complex conjugate of $a_{ij}\in{\mathbb{C}}$. Let $\pi:{\mathfrak{gl}}_{d}\longrightarrow{\mathfrak{gl}}_{d}$ be the *-structure on ${\mathfrak{gl}}_{d}$ defined by
-$\pi(e_{ab})=e_{ba}\quad\mbox{for $1\leq a,b\leq d$.}$
-The maps $\pi$ and ${\boldsymbol{\sigma}}$ give rise to a *-structure $(\pi,{\boldsymbol{\sigma}})$ on ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$.
-5.2. Bethe algebras of unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules
-The Weyl superalgebra $\mathcal{D}$ is a subalgebra of the endomorphism algebra ${\rm End}(\mathcal{F})$ of the polynomial superalgebra $\mathcal{F}$.
-For ${\bf w}\in{\mathbb{C}}^{d}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$, Theorem 4.7 can be rephrased as saying that the image of $\mathcal{B}_{d}^{\bf w}$ in ${\rm End}(\mathcal{F})$ under $\phi_{\bf z}$ coincides with the image of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ in ${\rm End}(\mathcal{F})$ under $\overline{\phi}_{\bf w}$.
+We have learned from an anonymous expert that the last assertion of Theorem 4.10 can be obtained directly from the results of [HM].
+An outline of his approach is given as follows.
+Let $\pi$ be an automorphism on $\mathcal{D}$ defined by
+$x^{a}_{i}\mapsto x^{a}_{i},\quad\partial_{x^{a}_{i}}\mapsto\partial_{x^{a}_{i}},\quad y^{a}_{r}\mapsto(-1)^{|r|+1}\partial_{y^{a}_{r}},\quad\partial_{y^{a}_{r}}\mapsto y^{a}_{r},$
+for $1\leq i\leq m+n$, $1\leq r\leq p+q$, and $1\leq a\leq d$, and let $\vartheta$ be an automorphism on $U({\mathfrak{gl}}_{d}[t])$ defined by
+$e_{ab}[k]\mapsto e_{ab}[k]-\delta_{a,b}\sum_{r=1}^{p+q}(-1)^{|r|}z_{r}^{k}$
+for $1\leq a,b\leq d$ and $k\in{\mathbb{Z}}_{+}$. Note that $\vartheta$ restricts to an automorphism on $\mathcal{B}_{d}^{\bf w}$. As a consequence of [HM, Theorem 5.2],
+$(\pi\circ\phi_{\bf z}\circ\vartheta)(\mathcal{B}_{d}^{\bf w})=(\pi\circ\varphi_{\bf w})({\mathcal{B}_{p+m|q+n}^{\bf z}}),$
+which is equivalent to the last assertion of Theorem 4.10.
+4.4. An application of the Bethe duality
+The Weyl superalgebra $\mathcal{D}$ is a subalgebra of ${\rm End}(\mathcal{F})$.
+For ${\bf w}\in{\mathbb{C}}^{d}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$, Theorem 4.10 implies that the image of $\mathcal{B}_{d}^{\bf w}$ in ${\rm End}(\mathcal{F})$ under $\phi_{\bf z}$ coincides with the image of $\mathcal{B}_{p+m|q+n}^{\bf z}$} in ${\rm End}(\mathcal{F})$ under $\varphi_{\bf w}$.
 This makes it possible to study the action of $\mathcal{B}_{p+m|q+n}^{\bf z}$} on $\mathcal{F}$ via the action of $\mathcal{B}_{d}^{\bf w}$ on $\mathcal{F}$, and vice versa.
-[…]
-[…]
-for $1\leq a\leq d$. As ${\mathfrak{gl}}_{d}$-modules,
-$\mathcal{F}\cong\left(\!\bigotimes_{r=1}^{p}{\mathbb{C}}[{\bf y}_{r}]\!\right)\otimes\left(\!\bigotimes_{s=1}^{q}{\mathbb{C}}[{\boldsymbol{\zeta}}_{s}]\!\right)\otimes\left(\!\bigotimes_{i=1}^{m}{\mathbb{C}}[{\bf x}_{i}]\!\right)\otimes\left(\!\bigotimes_{j=1}^{n}{\mathbb{C}}[{\boldsymbol{\eta}}_{j}]\!\right)\!,$ (5.1)
-and as ${\mathfrak{gl}}_{p+m|q+n}$-modules,
-$\mathcal{F}\cong{\mathbb{C}}[{\boldsymbol{\Sigma}}^{1}]\otimes\cdots\otimes{\mathbb{C}}[{\boldsymbol{\Sigma}}^{d}].$ (5.2)
-The Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ asserts that $\mathcal{F}$ decomposes into a direct sum of unitarizable ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-modules (with respect to $(\pi,{\boldsymbol{\sigma}})$).
-We refer the reader to [CLZ, Theorem 3.3] for further details.
+Let
+${\bf d}_{\mathcal{F}}=\sum_{a=1}^{d}\left(\!\sum_{i=1}^{m}x_{i}^{a}\partial_{x^{a}_{i}}+\sum_{j=1}^{n}\eta_{j}^{a}\partial_{\eta^{a}_{j}}-\sum_{r=1}^{p}y_{r}^{a}\partial_{y^{a}_{r}}-\sum_{s=1}^{q}\zeta_{s}^{a}\partial_{\zeta^{a}_{s}}\!\right)\!.$
+For any monomial $f\in\mathcal{F}$, we have ${\bf d}_{\mathcal{F}}(f)=kf$ for some $k\in{\mathbb{Z}}$.
+We define the degree of $f$ to be $k$.
+In particular, the degrees of $x^{a}_{i}$} and $\eta^{a}_{j}$} are $1$ while the degree of $y^{a}_{r}$ and $\zeta^{a}_{s}$ are 1-1.
+We call $\bf d$$_{\mathcal{F}}$ the degree operator on $\mathcal{F}$.
+For $1\leq r\leq p$, $1\leq s\leq q$, $1\leq i\leq m$, and $1\leq j\leq n$, let $\mathcal{F}_{(r)}={\mathbb{C}}\big[y^{1}_{r},\ldots,y^{d}_{r}\big]$, $\mathcal{F}_{(p+s)}={\mathbb{C}}\big[\zeta^{1}_{s},\ldots,\zeta^{d}_{s}\big]$, $\mathcal{F}_{(p+q+i)}={\mathbb{C}}\big[x^{1}_{i},\ldots,x^{d}_{i}\big]$, and $\mathcal{F}_{(p+q+m+j)}={\mathbb{C}}\big[\eta^{1}_{j},\ldots,\eta^{d}_{j}\big]$.
+It is evident that as ${\mathfrak{gl}}_{d}$-modules,
+$\mathcal{F}=\mathcal{F}_{(1)}\otimes\cdots\otimes\mathcal{F}_{(p+q+m+n)}.$ (4.14)
+For $1\leq a\leq d$, let $\mathcal{F}^{(a)}={\mathbb{C}}\big[x^{a}_{i},\eta^{a}_{j},y^{a}_{r},\zeta^{a}_{s}\,\big|\,1\leq i\leq m,1\leq j\leq n,1\leq r\leq p,1\leq s\leq q\big]$.
+As ${\mathfrak{gl}}_{p+m|q+n}$-modules,
+$\mathcal{F}=\mathcal{F}^{(1)}\otimes\cdots\otimes\mathcal{F}^{(d)}.$ (4.15)
+For any weight $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ of a ${\mathfrak{gl}}_{p+m|q+n}$-module $M$, we denote by $M_{\mu}$ the $\mu$-weight space of $M$.
+The Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ of ${\mathfrak{gl}}_{p+m|q+n}$ acts on $\mathcal{F}$ as follows (see (4.10)): $E^{i}_{i}$ acts on $\mathcal{F}$ by $\sum_{a=1}^{d}y^{a}_{i}\partial_{y^{a}_{i}}-d,$ for $1\leq i\leq p$, and by $\sum_{a=1}^{d}\zeta^{a}_{i-p}\partial_{\zeta^{a}_{i-p}}+d,$ for $p+1\leq i\leq q$.
+Also, $E^{p+q+i}_{p+q+i}$ acts on $\mathcal{F}$ by $\sum_{a=1}^{d}x^{a}_{i}\partial_{x^{a}_{i}},$ for $1\leq i\leq m$, and by $\sum_{a=1}^{d}\eta^{a}_{i-m}\partial_{\eta^{a}_{i-m}},$ for $m+1\leq i\leq m+n$.
+Clearly, there is a weight space decomposition
+$\mathcal{F}=\bigoplus_{\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}}\mathcal{F}_{\mu}.$ (4.16)
 For $\xi\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$, we denote by $L_{p+m|q+n}(\xi)$ the irreducible highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module with highest weight $\xi$ with respect to the Borel subalgebra $\mathfrak{b}_{{p+m|q+n}}$.
@@
 Suppose $p\not=0\not=m$. For $r\in{\mathbb{Z}}$, let
-$V_{r}=\begin{cases}\displaystyle{L_{{p+m|q+n}}\big{(}r\epsilon_{p+q+1}-\mathbbm{1}_{p|q}\big{)}}&\quad\mbox{if $r\geq 0$,}\\ \displaystyle{L_{{p+m|q+n}}\left(\!-\sum_{i=q+r+1}^{q}\epsilon_{p+i}-\mathbbm{1}_{p|q}\!\right)}&\quad\mbox{if $-q\leq r<0$ and $q\not=0$,}\\ \displaystyle{L_{{p+m|q+n}}\left(\!(r+q)\epsilon_{p}-\sum_{i=1}^{q}\epsilon_{p+i}-\mathbbm{1}_{p|q}\!\right)}&\quad\mbox{if $r<-q$ and $q\not=0$,}\\ \displaystyle{L_{{p+m|q+n}}\big{(}r\epsilon_{p}-\mathbbm{1}_{p}\big{)}}&\quad\mbox{if $r<0$ and $q=0$}.\end{cases}$
-Note that $V_{r}$ is an infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module (with respect to ${\boldsymbol{\sigma}}$), and that the highest weight of $V_{r}$ corresponds to a generalized partition of depth 1; see [CCL2, Section 2.5] and also [CLZ].
+[…]
+The space $V_{r}$ is an infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module (with respect to ${\bm{\sigma}}_{p}$), and the highest weight of $V_{r}$, appearing in (4.17), corresponds to a generalized partition of depth 1; see [CCL1, Section 2.5] and also [CLZ].
 (Such a partition is called a generalized partition of length 1 in [CLZ].)
-For any weight $\mu$ of a ${\mathfrak{gl}}_{p+m|q+n}$-module $M$, we denote by $M_{\mu}$ the $\mu$-weight space of $M$.
+For $a=1,\ldots,d$, let $\gamma_{a}\in{\mathbb{Z}}$.
+If $\gamma_{a}\geq 0$, then we may identify $V_{\gamma_{a}}$ with the ${\mathfrak{gl}}_{p+m|q+n}$-submodule of the polynomial superalgebra $\mathcal{F}^{(a)}$ generated by the highest weight vector $(x^{a}_{1})^{\gamma_{a}}$.
+If $\gamma_{a}<0$, then we may identify $V_{\gamma_{a}}$ with the ${\mathfrak{gl}}_{p+m|q+n}$-submodule of $\mathcal{F}^{(a)}$ generated by the highest weight vector
+$\begin{cases}\zeta^{a}_{q+\gamma_{a}+1}\ldots\zeta^{a}_{q}&\quad\text{if }\mbox{$\gamma_{a}\geq-q$ and $q\not=0$,}\\ (y^{a}_{p})^{-\gamma_{a}-q}\zeta^{a}_{1}\ldots\zeta^{a}_{q}&\quad\text{if }\mbox{$\gamma_{a}<-q$ and $q\not=0$,}\\ (y^{a}_{p})^{-\gamma_{a}}&\quad\text{if }q=0.\end{cases}$
+In any case, $V_{\gamma_{a}}$ is the subspace of $\mathcal{F}^{(a)}$ spanned by all monomials of degree $\gamma_{a}$.
+Via (4.15), the tensor product $V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}}$ is a ${\mathfrak{gl}}_{p+m|q+n}$-submodule of $\mathcal{F}$.
+We thus obtain a direct sum decomposition of ${\mathfrak{gl}}_{p+m|q+n}$-modules:
+$\mathcal{F}=\bigoplus_{\gamma_{1},\ldots,\gamma_{d}\in{\mathbb{Z}}}V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}}.$ (4.18)
+Let ${\mathfrak{h}}_{d}$ denote the standard Cartan subalgebra of ${\mathfrak{gl}}_{d}$.
+The action of ${\mathfrak{h}}_{d}$ on $\mathcal{F}$ is determined by the action of $e_{aa}$ on $\mathcal{F}$, which is given by
+$\sum_{i=1}^{m}x_{i}^{a}\partial_{x^{a}_{i}}+\sum_{j=1}^{n}\eta_{j}^{a}\partial_{\eta^{a}_{j}}-\sum_{r=1}^{p}y_{r}^{a}\partial_{y^{a}_{r}}-\sum_{s=1}^{q}\zeta_{s}^{a}\partial_{\zeta^{a}_{s}},$
+for $1\leq a\leq d$ (see (4.9)). There is a weight space decomposition
+$\mathcal{F}=\bigoplus_{\gamma\in{\mathfrak{h}}_{d}^{*}}\mathcal{F}_{\gamma}.$ (4.19)
+For $1\leq i\leq p+q$ and $k\in-{\mathbb{Z}}_{+}$, let $\mathcal{F}_{(i)}^{k}$ be the subspace of $\mathcal{F}_{(i)}$ spanned by all monomials of degree $k$.
+For $1\leq j\leq m+n$ and $l\in{\mathbb{Z}}_{+}$, let $\mathcal{F}_{(p+q+j)}^{l}$ be the subspace of $\mathcal{F}_{(p+q+j)}$ spanned by all monomials of degree $l$.
+Evidently, $\mathcal{F}_{(i)}^{k}$ and $\mathcal{F}_{(p+q+j)}^{l}$ are finite-dimensional ${\mathfrak{gl}}_{d}$-modules.
+We have the following direct sum decompositions of ${\mathfrak{gl}}_{d}$-modules:
+$\mathcal{F}_{(i)}=\bigoplus_{k\in-{\mathbb{Z}}_{+}}\mathcal{F}_{(i)}^{k}\qquad\text{and}\qquad\mathcal{F}_{(p+q+j)}=\bigoplus_{l\in{\mathbb{Z}}_{+}}\mathcal{F}_{(p+q+j)}^{l}.$
+It is well known that $\mathcal{F}_{(i)}^{k}$ and $\mathcal{F}_{(p+q+j)}^{l}$ are irreducible ${\mathfrak{gl}}_{d}$-modules for $k\in-{\mathbb{Z}}_{+}$ and $l\in{\mathbb{Z}}_{+}$ (see, for instance, [CLZ, Theorem 3.3] by setting exactly one of $p,q,m,n$ (given there) to 1 and the others to 0).
+By (4.14), $\mathcal{F}$ decomposes into a direct sum of ${\mathfrak{gl}}_{d}$-modules
+$\mathcal{F}=\bigoplus_{\begin{subarray}{c}k_{1},\ldots,k_{p+q}\in-{\mathbb{Z}}_{+},\\ k_{p+q+1},\ldots,k_{p+q+m+n}\in{\mathbb{Z}}_{+}\end{subarray}}\mathcal{F}_{(1)}^{k_{1}}\otimes\cdots\otimes\mathcal{F}_{({p+q+m+n})}^{k_{p+q+m+n}}.$ (4.20)
+Proposition 4.12.
+For $\gamma_{1},\ldots,\gamma_{d}\in{\mathbb{Z}}$, let $\gamma=\sum_{a=1}^{d}\gamma_{a}\epsilon_{a}\in{\mathfrak{h}}_{d}^{*}$.
+Let $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ be any weight of $V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}}$.
+Then
+$(V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}})_{\mu}=\big(\mathcal{F}_{(1)}^{\overline{\mu}_{1}}\otimes\cdots\otimes\mathcal{F}_{({p+q+m+n})}^{\overline{\mu}_{p+q+m+n}}\big)_{\gamma},$
+where
+$\overline{\mu}_{i}=\begin{cases}\,\mu(E^{i}_{i})+(-1)^{|i|}d&\mbox{if}\ \ 1\leq i\leq p+q;\\ \,\mu(E^{i}_{i})\ \ &\mbox{if}\ \ p+q+1\leq i\leq p+q+m+n.\end{cases}$ (4.21)
+Proof.
+By (4.18) and (4.19), we have
+$\mathcal{F}_{\gamma}=V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}}.$
+Also, (4.16) and (4.20) yield
+$\mathcal{F}_{\mu}=\mathcal{F}_{(1)}^{\overline{\mu}_{1}}\otimes\cdots\otimes\mathcal{F}_{({p+q+m+n})}^{\overline{\mu}_{p+q+m+n}}.$
+Consequently,
+$(V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}})_{\mu}=\mathcal{F}_{\gamma}\cap\mathcal{F}_{\mu}=\big(\mathcal{F}_{(1)}^{\overline{\mu}_{1}}\otimes\cdots\otimes\mathcal{F}_{\overline{\mu}_{p+q+m+n}}^{(p+q+m+n)}\big)_{\gamma},$
+as asserted. ∎
 For ${\mathfrak{gl}}_{p+m|q+n}$-modules $M_{1},\ldots,M_{d}$ and ${\bf w}=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$, we write
 $\underline{M\mkern-4.0mu}\mkern 4.0mu=M_{1}\otimes\cdots\otimes M_{d}\qquad\text{and}\qquad\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})=M_{1}(w_{1})\otimes\cdots\otimes M_{d}(w_{d}).$
-Analogous to (4.16), we write
-$L_{p+m}(\xi)=L_{p+m|0+0}(\xi)\qquad\text{and}\qquad L_{m}(\xi)=L_{0+m|0+0}(\xi).$
 Theorem 4.13.
 Suppose $p\not=0\not=m$.
-For $a=1,\ldots,d$, let $M_{a}=V_{\gamma_{a}}$ for some $\gamma_{a}\in{\mathbb{Z}}$.
-For each weight $\mu$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!p+q+m+n}$, we have:
+For $a=1,\ldots,d$, let $L_{a}=V_{\gamma_{a}}$ for some $\gamma_{a}\in{\mathbb{Z}}$.
+For any weight $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ of $\underline{L}$, ${\bf z}\in{\bm{X}}_{\!p+q+m+n}$, and ${\bf w}\in{\bm{X}}_{\!d}$, we have:
   1. (i)
-    $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}$ is a cyclic ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module.
+    $\underline{L}({\bf w})_{\mu}$ is a cyclic $\mathcal{B}_{p+m|q+n}^{\bf z}$}-module.
   2. (ii)
-[…]
+    $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{\underline{L}({\bf w})_{\mu}}$ is a Frobenius algebra.
   3. (iii)
-[…]
+    $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{\underline{L}({\bf w})_{\mu}}$ is a maximal commutative subalgebra of ${\rm End}(\underline{L}({\bf w})_{\mu})$ of dimension $\dim\!\left(\underline{L}({\bf w})_{\mu}\right)$.
   4. (iv)
-[…]
-[…]
+    Every eigenspace of $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{\underline{L}({\bf w})_{\mu}}$ is one-dimensional.
+Moreover, $\mathcal{B}_{p+m|q+n}^{\bf z}$} is diagonalizable with a simple spectrum on $\underline{L}({\bf w})_{\mu}$ for generic $\bf z$} and generic $\bf w$}.
 Proof.
-Fix $1\leq a\leq d$.
-If $\gamma_{a}\geq 0$, then we may identify $M_{a}$ with the ${\mathfrak{gl}}_{p+m|q+n}$-submodule of the polynomial superalgebra ${\mathbb{C}}[{\boldsymbol{\Sigma}}^{a}]$ generated by the highest weight vector $(x^{a}_{1})^{\gamma_{a}}$.
-If $\gamma_{a}<0$, then we may identify $M_{a}$ with the ${\mathfrak{gl}}_{p+m|q+n}$-submodule of ${\mathbb{C}}[{\boldsymbol{\Sigma}}^{a}]$ generated by the highest weight vector
-$\begin{cases}\zeta^{a}_{q+\gamma_{a}+1}\ldots\zeta^{a}_{q}&\quad\text{if }\mbox{$\gamma_{a}\geq-q$ and $q\not=0$,}\\ (y^{a}_{p})^{-\gamma_{a}-q}\zeta^{a}_{1}\ldots\zeta^{a}_{q}&\quad\text{if }\mbox{$\gamma_{a}<-q$ and $q\not=0$,}\\ (y^{a}_{p})^{-\gamma_{a}}&\quad\text{if }q=0.\end{cases}$
-Under these identifications, the tensor product $\underline{M\mkern-4.0mu}\mkern 4.0mu$ is a ${\mathfrak{gl}}_{p+m|q+n}$-submodule of ${\mathbb{C}}[{\boldsymbol{\Sigma}}^{1}]\otimes\cdots\otimes{\mathbb{C}}[{\boldsymbol{\Sigma}}^{d}]$ and hence a ${\mathfrak{gl}}_{p+m|q+n}$-submodule of $\mathcal{F}$ via (5.2).
-We will see in a moment that for each weight $\mu:=\sum_{i=1}^{p+q+m+n}\mu_{i}\epsilon_{i}$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, the $\mu$-weight space $\underline{M\mkern-4.0mu}\mkern 4.0mu_{\!\mu}$ is also a weight space of a tensor product of finite-dimensional irreducible ${\mathfrak{gl}}_{d}$-modules.
-For $1\leq i\leq p$ (resp., $p+1\leq i\leq p+q$, $p+q+1\leq i\leq p+q+m$ and $p+q+m+1\leq i\leq p+q+m+n$), let $N_{i}$ denote the ${\mathfrak{gl}}_{d}$-submodule of ${\mathbb{C}}[y^{1}_{i},\cdots,y^{d}_{i}]$ (resp., ${\mathbb{C}}[\zeta^{1}_{i-p},\cdots,\zeta^{d}_{i-p}]$, ${\mathbb{C}}[x^{1}_{i-p-q},\cdots,x^{d}_{i-p-q}]$ and ${\mathbb{C}}[\eta^{1}_{i-p-q-m},\cdots,\eta^{d}_{i-p-q-m}]$) generated by the highest weight vector $(y^{d}_{i})^{-\mu_{i}-d}$ (resp., $\zeta^{\mu_{i}+1}_{i-p}\cdots\zeta^{d}_{i-p}$, $(x^{1}_{i-p-q})^{\mu_{i}}$ and $\eta^{1}_{i-p-q-m}\cdots\eta^{\mu_{i}}_{i-p-q-m}$).
-Then
-$N_{i}\cong\begin{cases}L_{d}\big{(}(\mu_{i}+d)\epsilon_{d}\big{)}&\quad\text{if }1\leq i\leq p,\\ \displaystyle{L_{d}\left(\!-\sum_{a=\mu_{i}+1}^{d}\epsilon_{a}\!\right)}&\quad\text{if }p+1\leq i\leq p+q,\\ L_{d}(\mu_{i}\epsilon_{1})&\quad\text{if }p+q+1\leq i\leq p+q+m,\\ \displaystyle{L_{d}\left(\sum_{a=1}^{\mu_{i}}\epsilon_{a}\right)}&\quad\text{if }p+q+m+1\leq i\leq p+q+m+n,\end{cases}$
-and so $N_{i}$ is an irreducible highest weight ${\mathfrak{gl}}_{d}$-module with a dominant integral highest weight. Write ${\bf z}=(z_{1},\ldots,z_{p+q+m+n})$. Let
-$\underline{N\mkern-4.0mu}\mkern 4.0mu=N_{1}\otimes\cdots\otimes N_{p+q+m+n}\quad\text{and}\quad\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})=N_{1}(z_{1})\otimes\cdots\otimes N_{p+q+m+n}(z_{p+q+m+n}).$
-The tensor product $\underline{N\mkern-4.0mu}\mkern 4.0mu$ is a ${\mathfrak{gl}}_{d}$-submodule of $\mathcal{F}$ via (5.1). By [CLZ, Theorem 3.3], we have
-$\underline{M\mkern-4.0mu}\mkern 4.0mu_{\mu}\subset\underline{N\mkern-4.0mu}\mkern 4.0mu\qquad\hbox{and}\qquad\underline{N\mkern-4.0mu}\mkern 4.0mu_{\gamma}\subset\underline{M\mkern-4.0mu}\mkern 4.0mu,$
-and no nonzero element in the $\nu$-weight space $\underline{N\mkern-4.0mu}\mkern 4.0mu_{\nu}$ is contained in $\underline{M\mkern-4.0mu}\mkern 4.0mu$ for all $\nu\not=\gamma$, where $\gamma:=\sum_{a=1}^{d}\gamma_{a}\epsilon_{a}\in{\mathfrak{h}}_{d}^{*}$.
-We deduce that $\underline{M\mkern-4.0mu}\mkern 4.0mu_{\mu}\subseteq\underline{N\mkern-4.0mu}\mkern 4.0mu_{\gamma}$.
-Similarly, $\underline{N\mkern-4.0mu}\mkern 4.0mu_{\gamma}\subseteq\underline{M\mkern-4.0mu}\mkern 4.0mu_{\mu}$.
-Thus we have
-$\underline{M\mkern-4.0mu}\mkern 4.0mu_{\mu}=\underline{N\mkern-4.0mu}\mkern 4.0mu_{\gamma}.$
-This gives
-[…]
-in view of Theorem 4.7.
+Let $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ be a weight of $\underline{L}$.
+For $i\in\mathbb{I}$, let $N_{i}=\mathcal{F}_{(i)}^{\overline{\mu}_{i}}$, where $\overline{\mu}_{i}$ is defined as in (4.21).
+Write ${\bf z}=(z_{1},\ldots,z_{p+q+m+n})$.
+We set $\underline{N\mkern-4.0mu}\mkern 4.0mu=N_{1}\otimes\cdots\otimes N_{p+q+m+n}$ and $\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})=N_{1}(z_{1})\otimes\cdots\otimes N_{p+q+m+n}(z_{p+q+m+n})$.
+In view of Theorem 4.10 and Proposition 4.12,
+$\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{\underline{L}({\bf w})_{\mu}}=(\mathcal{B}_{d}^{\bf w})_{\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}}.$ (4.22)
 By Theorem 3.3, $\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}$ is a cyclic $\mathcal{B}_{d}^{\bf w}$-module as $\mathcal{B}_{d}^{\bf w}$ preserves weight spaces, and by Theorem 3.14, $(\mathcal{B}_{d}^{\bf w})_{\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}}$ is a Frobenius algebra.
@@
 A similar statement corresponding to $p=0$ or $m=0$ can be formulated. The details will be left to the reader.
+Theorem 4.13 suggests that the super analog of Conjecture 3.13 should be true for the general linear Lie superalgebra.
+Conjecture 4.15.
+Suppose $p$, $q$, $m$, and $n$ are not all zero.
+For $a=1,\ldots,d$, let $L_{a}$ be a unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module (with respect to ${\bm{\sigma}}_{p}$).
+For any weight $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ of $\underline{L}$, ${\bf z}\in{\bm{X}}_{\!p+q+m+n}$, and ${\bf w}\in{\bm{X}}_{\!d}$, parts (i)–(iv) of Theorem 4.13 hold for the Bethe algebra $({\mathcal{B}_{p+m|q+n}^{\bf z}})_{\underline{L}({\bf w})_{\mu}}$.
+Moreover, $\mathcal{B}_{p+m|q+n}^{\bf z}$} is diagonalizable with a simple spectrum on $\underline{L}({\bf w})_{\mu}$ for generic $\bf z$} and generic $\bf w$}.
+Remark 4.16.
+An analog of Conjecture 4.15 when the entries of $\bf z$} are all zero is studied in [CCL2].
 Suppose $p\not=0\not=m$. For $r\in{\mathbb{Z}}$, let
-$V_{r}=\begin{cases}L_{p+m}(r\epsilon_{p+1}-\mathbbm{1}_{p})&\quad\mbox{if $r\geq 0$,}\\ L_{p+m}(r\epsilon_{p}-\mathbbm{1}_{p})&\quad\mbox{if $r<0$}.\end{cases}$
-We have the following result.
+$W_{r}=\begin{cases}L_{p+m|0}(r\epsilon_{p+1}-\mathbbm{1}_{p})&\quad\mbox{if $r\geq 0$,}\\ L_{p+m|0}(r\epsilon_{p}-\mathbbm{1}_{p})&\quad\mbox{if $r<0$}.\end{cases}$
+The following provides evidence for Conjecture 3.13.
 Corollary 4.17.
-For $a=1,\ldots,d$, let $M_{a}=V_{\gamma_{a}}$ for some $\gamma_{a}\in{\mathbb{Z}}$.
-For each weight $\mu$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!p+m}$, we have:
-  1. (i)
-    $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}$ is a cyclic $\mathcal{B}_{p+m}^{\bf z}$-module.
-  2. (ii)
-    $(\mathcal{B}_{p+m}^{\bf z})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is a Frobenius algebra.
-  3. (iii)
-    $(\mathcal{B}_{p+m}^{\bf z})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is a maximal commutative subalgebra of ${\rm End}(\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu})$ of dimension $\dim\!\left(\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}\right)$.
-  4. (iv)
-    Every eigenspace of $(\mathcal{B}_{p+m}^{\bf z})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is one-dimensional.
-Moreover, $(\mathcal{B}_{p+m}^{\bf z})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is diagonalizable with a simple spectrum for generic ${\bf w}$ and generic ${\bf z}$.
+For $a=1,\ldots,d$, let $L_{a}=W_{\gamma_{a}}$ for some $\gamma_{a}\in{\mathbb{Z}}$.
+For any weight $\mu\in{\mathfrak{h}}_{p+m}^{*}$ of $\underline{L}$, ${\bf z}\in{\bm{X}}_{\!p+m}$, and ${\bf w}\in{\bm{X}}_{\!d}$, parts (i)–(iv) of Theorem 4.13 hold for the Bethe algebra $(\mathcal{B}_{p+m}^{\bf z})_{\underline{L}({\bf w})_{\mu}}$.
+Moreover, $\mathcal{B}_{p+m}^{\bf z}$ is diagonalizable with a simple spectrum on $\underline{L}({\bf w})_{\mu}$ for generic $\bf z$} and generic $\bf w$}.
 Proof.
 This is an immediate consequence of Theorem 4.13 by taking $q=n=0$. ∎
-The structures of the Bethe algebras of finite-dimensional modules over a general linear Lie algebra have been well studied (see [FFRy, MTV2]).
-However, very little is known about the Bethe algebras of infinite-dimensional modules.
-Corollary 5.3, which concerns the action of the Bethe algebra of ${\mathfrak{gl}}_{p+m}$ on a tensor product of infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m}$-modules, makes good progress toward the infinite-dimensional cases.
-We expect that Theorem 2.1 and Theorem 2.2 should have unitarizable analogs. More generally, we have the following conjecture.
-Conjecture 5.4.
-Let $d\in{\mathbb{N}}$ and $p,q,m,n\in{\mathbb{Z}}_{+}$ that are not all zero.
-For $a=1,\ldots,d$, let $M_{a}$ be a unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module.
-For each weight $\mu$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!p+q+m+n}$, parts (i)–(iv) of Theorem 5.1 hold for the Bethe algebra $({\mathcal{B}_{p+m|q+n}^{\bf z}})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$.
-Moreover, $({\mathcal{B}_{p+m|q+n}^{\bf z}})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is diagonalizable with a simple spectrum for generic ${\bf w}$ and generic ${\bf z}$.
-Remark 5.5.
-An analog of the conjecture when the entries of ${\bf z}$ are all zero will be studied in [CCL2].
-Acknowledgments. The first author was partially supported by NSTC grant 113-2115-M-006-010 of Taiwan. The second author was partially supported by NSTC grant 112-2115-M-006-015-MY2 of Taiwan.
+Acknowledgments.
+The first author was partially supported by NSTC grants 113-2115-M-006-010 and 114-2115-M-006-005.
+The second author was partially supported by NSTC grant 112-2115-M-006-015-MY2.
+The authors are grateful to an anonymous expert for valuable comments on Theorem 4.10 (see Remark 4.11).
+Appendix A The proof of Proposition 4.7
+The appendix is devoted to proving Proposition 4.7. We keep the notations defined in Section 4. Let
+$\mathcal{C}=\left[\delta_{a,b}(z-w_{a})+\sum_{r=1}^{p+q}\frac{y^{a}_{r}\partial_{y^{b}_{r}}}{\partial_{z}-z_{r}}-\sum_{i=1}^{m+n}\frac{x^{b}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}\right]_{a,b=1,\ldots,d}.$
+Lemma A.1.
+$\omega\big(\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\big)={\rm cdet}(\mathcal{C})$.
+Proof.
+For $1\leq a,b\leq d$,
+$\phi_{\bf z}\big(e_{ab}(z)\big)=-\sum_{r=1}^{p+q}\frac{y^{b}_{r}\partial_{y^{a}_{r}}}{z-z_{r}}+\sum_{i=1}^{m+n}\frac{x^{a}_{i}\partial_{x^{b}_{i}}}{z-z_{p+q+i}}.$ (A.1)
+For $i\in\mathbb{I}\cup\{0\}$ and $1\leq a,b\leq d$, we define
+$F^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(\partial_{z}-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{b}_{i}\partial_{y^{a}_{i}}}{z-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{a}_{i-p-q}\partial_{x^{b}_{i-p-q}}}{z-z_{i}}}&\mbox{otherwise},\end{cases}$
+and
+$G^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(z-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{a}_{i}\partial_{y^{b}_{i}}}{\partial_{z}-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{b}_{i-p-q}\partial_{x^{a}_{i-p-q}}}{\partial_{z}-z_{i}}}&\mbox{otherwise}.\end{cases}$
+We see that
+$\omega(F^{i}_{a,b})=G^{i}_{a,b}.$ (A.2)
+Note that ${\mathcal{L}}_{d}^{\bf w}$ is an amply invertible Manin matrix of type $(0^{d})$.
+Let $\pi\in\mathfrak{S}_{d}$ be the permutation given by $\pi(i)=d+1-i$ for $i=1,\ldots,d$.
+Clearly, $\pi^{2}=1$.
+By (A.1), Proposition 4.5 and Proposition 4.6, we obtain
+$\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ $=$ $\phi_{\bf z}{\rm cdet}\big(({\mathcal{L}}_{d}^{\bf w})^{\pi}\big)$
+$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\pi(\sigma(1)),\pi(1)}\ldots F^{i_{1}}_{\pi(\sigma(d)),\pi(d)}$
+$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\pi\sigma(1),d}\ldots F^{i_{1}}_{\pi\sigma(d),1}$
+$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\sigma(d),d}\ldots F^{i_{1}}_{\sigma(1),1}.$
+Here the last equality holds since $\pi^{-1}\mathfrak{S}_{d}\pi=\mathfrak{S}_{d}$ and $l(\pi^{-1}\sigma\pi)=l(\sigma)$ for all $\sigma\in\mathfrak{S}_{d}$. Thus,
+$\omega\big(\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\big)=\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}\omega\!\left(F^{i_{1}}_{\sigma(1),1}\right)\ldots\omega\!\left(F^{i_{d}}_{\sigma(d),d}\right).$
+In view of (A.2), we deduce that $\omega\big(\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\big)={\rm cdet}(\mathcal{C})$. ∎
+Now we are ready to prove Proposition 4.7. For $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$, we have
+$\begin{array}[]{cll}\varphi_{\bf w}\big(E^{r}_{s}(z)\big)\hskip-28.45274pt&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}\partial_{y^{a}_{r}}y^{a}_{s}}{z-w_{a}},\quad\varphi_{\bf w}\big(E^{r}_{p+q+j}(z)\big)\hskip-8.5359pt&=\sum_{a=1}^{d}\frac{\partial_{y^{a}_{r}}\partial_{x^{a}_{j}}}{z-w_{a}},\\ \varphi_{\bf w}\big(E^{p+q+i}_{s}(z)\big)\hskip-8.5359pt&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}x^{a}_{i}y^{a}_{s}}{z-w_{a}},\quad\varphi_{\bf w}\big(E^{p+q+i}_{p+q+j}(z)\big)\hskip-8.5359pt&=\sum_{a=1}^{d}\frac{x^{a}_{i}\partial_{x^{a}_{j}}}{z-w_{a}}.\end{array}$} (A.3)
+Let
+$D=\Big[\delta_{a,b}(z-w_{a})\Big]_{a,b=1,\ldots,d}\qquad\text{and}\qquad D^{\prime}=\Big[\delta_{i,j}(\partial_{z}-z_{i})\Big]_{i,j\in\mathbb{I}}.$
+Consider the $(m+n)\times d$ matrices
+$X:=\big[(-1)^{|p+q+i|}x^{a}_{i}\big]^{a=1,\ldots,d}_{i=1,\ldots,m+n}\qquad\text{and}\qquad P_{X}:=\big[\partial_{x^{a}_{i}}\big]^{a=1,\ldots,d}_{i=1,\ldots,m+n}$
+and the $(p+q)\times d$ matrices
+$Y:=\big[(-1)^{|r|+1}y^{a}_{r}\big]^{a=1,\ldots,d}_{r=1,\ldots,p+q}\qquad\text{and}\qquad P_{Y}:=\big[(-1)^{|r|}\partial_{y^{a}_{r}}\big]^{a=1,\ldots,d}_{r=1,\ldots,p+q}.$
+Write
+$T=\begin{bmatrix}Y^{t}\hskip 5.69054ptP_{X}^{t}\end{bmatrix}\qquad\text{and}\qquad T^{\prime}=\begin{bmatrix}P_{Y}\vskip 8.53581pt\\ X\end{bmatrix}$.
+By (A.3), we get
+$\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)={\rm Ber}^{\bf s}(D^{\prime}-T^{\prime}D^{-1}T).$
+Let
+$\mathfrak{L}=\begin{bmatrix}D&T\vskip 5.69054pt\\ T^{\prime}&D^{\prime}\end{bmatrix}$.
+It is straightforward to check that $\mathfrak{L}$ is an amply invertible Manin matrix of type $\hat{{\bf s}}:=(0^{d},{\bf s})\in\mathcal{S}_{d+p+m|q+n}$. By Proposition 4.1, we have
+${\rm Ber}^{\hat{{\bf s}}}(\mathfrak{L})=(z-w_{1})\ldots(z-w_{d})\,\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big).$
+As $D-T(D^{\prime})^{-1}T^{\prime}$ is a sufficiently invertible Manin matrix of type $(0^{d})$, Proposition 4.4 and Proposition 4.5 imply that
+${\rm Ber}^{\hat{{\bf s}}}(\mathfrak{L})$ $=$ ${\rm Ber}^{\bf s}(D^{\prime})\cdot{\rm cdet}(D-T(D^{\prime})^{-1}T^{\prime})$
+$=$ $R_{p|q}(\partial_{z})\cdot R^{\prime}_{m|n}(\partial_{z})\cdot{\rm cdet}(\mathcal{C}^{\prime}),$
+where
+$\mathcal{C}^{\prime}:=\left[\delta_{a,b}(z-w_{a})+\sum_{r=1}^{p+q}\frac{y^{a}_{r}\partial_{y^{b}_{r}}}{\partial_{z}-z_{r}}-\sum_{i=1}^{m+n}\frac{(-1)^{|p+q+i|}\partial_{x^{a}_{i}}x^{b}_{i}}{\partial_{z}-z_{p+q+i}}\right]_{a,b=1,\ldots,d}.$
+We claim that
+$R^{\prime}_{m|n}(\partial_{z})\cdot{\rm cdet}(\mathcal{C}^{\prime})={\rm cdet}(\mathcal{C})\cdot R^{\prime}_{m|n}(\partial_{z}).$
+This yields
+${\rm Ber}^{\hat{{\bf s}}}(\mathfrak{L})=R_{p|q}(\partial_{z})\cdot{\rm cdet}(\mathcal{C})\cdot R^{\prime}_{m|n}(\partial_{z}),$
+which proves Proposition 4.7 by Lemma A.1.
+It remains to prove the claim. We only need to show that
+$R^{\prime}_{m|n}(\partial_{z})\cdot\mathcal{C}^{\prime}_{a,b}=\mathcal{C}_{a,b}\cdot R^{\prime}_{m|n}(\partial_{z}),\qquad\mbox{for $1\leq a,b\leq d$.}$ (A.4)
+Here, for instance, $\mathcal{C}_{a,b}$ denotes the $(a,b)$ entry of $\mathcal{C}$. Since
+$\partial_{x^{a}_{i}}x^{b}_{i}=(-1)^{|p+q+i|}x^{b}_{i}\partial_{x^{a}_{i}}+\delta_{a,b},\qquad\mbox{for $1\leq a,b\leq d$},$ (A.5)
+the off-diagonal entries of $\mathcal{C}^{\prime}$ coincide with those of $\mathcal{C}$.
+For $1\leq i\leq m+n$, the term $\partial_{z}-z_{p+q+i}$ commutes with $\partial_{x^{a}_{j}}x^{b}_{j}$ and $y^{a}_{r}\partial_{y^{b}_{r}}$ for all $1\leq j\leq m+n$, $1\leq r\leq p+q$, and $1\leq a,b\leq d$.
+In particular, $R^{\prime}_{m|n}(\partial_{z})$ commutes with any off-diagonal entry of $\mathcal{C}^{\prime}$.
+This proves (A.4) for $a\not=b$.
+Moreover, for $1\leq i\leq m+n$ and $1\leq a\leq d$,
+$(\partial_{z}-z_{p+q+i})(z-w_{a})=(z-w_{a})(\partial_{z}-z_{p+q+i})+1.$ (A.6)
+If $1\leq i\leq m$, we see from (A.5) and (A.6) that
+${(\partial_{z}-z_{p+q+i})}\left(z-w_{a}-\frac{\partial_{x^{a}_{i}}x^{a}_{i}}{\partial_{z}-z_{p+q+i}}\right)=\left(z-w_{a}-\frac{x^{a}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}\right)(\partial_{z}-z_{p+q+i}).$
+Similarly, if $m+1\leq i\leq m+n$, we have
+${(\partial_{z}-z_{p+q+i})^{-1}}\left(z-w_{a}+\frac{\partial_{x^{a}_{i}}x^{a}_{i}}{\partial_{z}-z_{p+q+i}}\right)=\left(z-w_{a}-\frac{x^{a}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}\right)(\partial_{z}-z_{p+q+i})^{-1}.$
+The above discussion proves (A.4) for $a=b$. Thus, the proof of (A.4) is now complete.